### v1

# Personalizing story generation by inferring author styles

Abstract
Personalization has become essential for improving user experience in interactive writing and educational applications, yet its potential in story generation remains largely unexplored. In this work, we propose a novel two-stage pipeline for personalized story generation. Our approach first infers an author’s implicit story-writing characteristics from their past work and organizes them into an Author Writing Sheet that is inspired by narrative theory. The second stage uses this sheet to simulate the author’s persona through tailored persona descriptions and personalized story writing rules. To enable and validate our approach, we construct Mythos, a dataset of 590 stories from 64 authors across 5 distinct sources that reflect diverse story-writing settings. A head-to-head comparison with a non-personalized baseline demonstrates our pipeline’s effectiveness in generating high-quality personalized stories. Our personalized stories achieve a 75% win rate (versus 14% for the baseline and 11% ties) in capturing authors’ writing style based on their past works. Human evaluation highlights the high quality of our Author Writing Sheet and provides valuable insights into the personalized story generation task. Notable takeaways are that writings from certain sources, such as Reddit, are easier to personalize than others, like AO3, while narrative aspects, like Creativity and Language Use, are easier to personalize than others, like Plot.
[Uncaptioned image]     github.com/Nish-19/Persona-Story-Gen
Whose story is it?
Personalizing story generation by inferring author styles
Nischal Ashok Kumar1  Chau Minh Pham2  Mohit Iyyer1,2  Andrew Lan1 1University of Massachusetts Amherst  2University of Maryland, College Park {nashokkumar, andrewlan}@cs.umass.edu, {chau, miyyer}@umd.edu
1 Introduction
Figure 1: Our two-stage pipeline for personalized story generation. Stage 1 constructs an Author Writing Sheet with Claim ($C$) and Evidence ($E$) pairs capturing the author’s story-writing characteristics across narrative categories. It is derived from the author’s history of writing prompts ($wp$), author-written stories ($s_{a}$), and LLM-generated Average Stories ($s_{b}$) representing a typical author’s response to the same prompt. Stage 2 uses the Author Writing Sheet to role-play the author, incorporating tailored story rules and a persona description for personalized generation.
Personalizing the outputs of large language models (LLMs) has promising applications for role-playing in chatbots and video games (Tseng et al., 2024; Chen et al., 2024), as well as providing personalized feedback in education (Stahl et al., 2024; Li et al., 2024b; Makridis et al., 2024). However, story-writing personalization, a crucial writing system feature that can assist users in overcoming writer’s block or help second-language learners adapt to a new language (Yuan et al., 2022; Yeh et al., 2024), remains largely unexplored.
To address this gap, we systematically explore how to personalize LLM outputs based on individual authors’ previous writings. We propose a novel two-stage pipeline for personalized story generation, as shown in Figure 1. In the first stage, we construct an Author Writing Sheet to capture implicit writing characteristics from an author’s past writings (Sun et al., 2025, 2024; Ramos et al., 2024). We categorize story-writing characteristics into four fine-grained narrative dimensions: Plot, Creativity, Development, and Language Use (Pavis, 1998; Card, 1999; Noble, 1994; Huot et al., 2024). Each characteristic is represented as a Claim-Evidence pair, where general statements about the author’s writing behavior are supported by specific story excerpts. This approach, inspired by Common Core Standards in English Language Arts (Association et al., 2010) and recent work in writing education (Li et al., 2023), provides a structured and grounded way to identify an author’s distinctive writing patterns, as validated via our human evaluation (Section 3). In the second stage, we implement Personalized Story Generation by having LLMs role-play authors based on their Writing Sheet. Our method uses persona prompts and story rules as user constraints to guide the generation.
To enable and validate our approach, we construct Mythos, a dataset of 590 stories from 64 authors ($\approx$1500 tokens) across five sources with distinct writing settings: Reddit, AO3, Storium, Narrative Magazine, and New Yorker. The dataset splits each author’s works chronologically into two sets: an earlier Profiling set used to infer their writing characteristics, and a later Generation set used to evaluate how well LLMs can emulate their distinctive writing behavior (Salemi et al., 2024).
We evaluate personalized stories in two ways: First, we use automatic evaluation with LLM-as-a-judge, to assess faithfulness to the author’s writing history and similarity to ground truth author writings, along with traditional metrics for lexical overlap, story diversity, and style similarity. Second, we conduct a human evaluation to assess the generated story’s similarity to the author’s ground-truth writings. We find that stories generated using Author Writing Sheets demonstrate clear advantages over a non-personalized baseline approach. When evaluated by LLMs, personalized stories have 61% higher faithfulness to the author writing history. Human annotators prefer these stories by an 18% absolute margin in terms of similarity to ground truth. Our personalization approach is especially effective for Reddit content and for capturing Creativity, and Language Use characteristics, with less pronounced impact on Plot (Tian et al., 2024; Xu et al., 2024). In summary, our contributions are:
  • Mythos: A novel dataset containing multiple stories written by the same authors, designed to support both author profiling and personalized story generation evaluation.
  • Author Writing Sheet: A structured approach to capturing an author’s writing behavior based on past works, grounded in narrative theory and educational practices and validated through human evaluation.
  • Personalized Story Generation: A method for generating stories in an author’s style using Author Writing Sheets, rigorously evaluated through automated and human assessments, demonstrating its superiority over baselines.
2 Dataset Collection and Description
To facilitate our analysis of personalized story generation, we ensemble Mythos, a novel dataset consisting of 590 stories written by 64 authors. To the best of our knowledge, Mythos is the first dataset that identifies and connects multiple stories written by the same author. See Table 1 for comparison with existing datasets.
Data Sources:
We include five diverse story-writing sources, namely (1) Reddit, featuring stories from r/WritingPrompts, a widely-used resource for amateur story-writing research (Fan et al., 2018), (2) AO3, featuring fanfiction based on popular franchises such as Harry Potter and Sherlock Holmes, (3) Storium, featuring collaborative stories (Akoury et al., 2020), (4) N.Magazine, featuring professionally polished stories by authors such as Barry Gifford, and finally (5) New Yorker, featuring expert-level storytelling from renowned authors such as Haruki Murakami.
Preprocessing:
To ensure high-quality and diverse content, we impose constraints on word length (ranging from 500 to 1500 words), limit publication dates to post-November 2023, and exclude explicit material. To standardize formatting, non-Reddit stories are supplemented with writing prompts generated by GPT-4o and manually reviewed for accuracy. Stories are split chronologically based on their submission timestamps, with 70% used for profiling writing characteristics and 30% for generation (see Appendix A for details).
Dataset size:
Our dataset is limited in size due to data crawling policies that restrict large-scale data collection for analysis [^1] https://archiveofourown.org/admin_posts/25888. We note that recent storywriting datasets such as TELL ME A STORY (Huot et al., 2024) and TTCW (Chakrabarty et al., 2024) have conducted analyses on a comparable number of stories (Table 1).
Table 1: Comparison of our dataset with existing story-writing datasets. ✗ indicates no Author IDs associated with stories, and $\sim$ denotes having Author IDs but not having explicit links between them. Our dataset spans diverse settings and links stories by the same author, enabling research on personalized story generation.
Dataset Name Size Sources Prompt Length Story Length Author IDs
WritingPrompts (Fan et al., 2018) 300K Reddit 28 735
TELL ME A STORY (Huot et al., 2024) 230 Writing Workshop 113 1498
MirrorStories (Yunusov et al., 2024) 1500 Aesop’s fables 40 400
Storium (Akoury et al., 2020) 440k Storium online platform 247 247 $\sim$
TTCW (Chakrabarty et al., 2024) 12 New Yorker 54 2000 $\sim$
Mythos 590 Reddit, AO3, Storium, N.Magazine, New Yorker 50 1517
3 Author Writing Sheet
Motivated by the connection between writing education and personalization (Li et al., 2023), we propose a novel method to infer implicit story-writing characteristics of authors based on their history in the profiling set $P$. We group these characteristics into four narrative categories: Plot, Creativity, Development, and Language Use, inspired by narrative theory (Pavis, 1998; Card, 1999; Noble, 1994; Huot et al., 2024), to assess fine-grained, multi-dimensional story-writing attributes:
  • Plot: Story structure, conflict introduction, prompt engagement, and resolution.
  • Creativity: Genre blending, unconventional prompt reinterpretation, and unique elements.
  • Development: Character depth, emotional arcs, and immersive settings.
  • Language Use: Diction, style, rhetorical devices, pacing, and dialogue.
Inspired by Common Core (CC) Standards in English Language and Arts (Association et al., 2010), each characteristic is represented as a Claim-Evidence pair, where the Claim summarizes the author’s writing style, and Evidence consists of excerpts from the author-written story to justify the Claim. These characteristics are organized into an Author Writing Sheet (Gurung and Lapata, 2024).
Stage 1 in Figure 1 (and Algorithm 1 in the Appendix) illustrates our method for generating the Author Writing Sheet by iteratively processing each story in an author’s profiling set. For a given writing prompt ($wp_{t}$) and corresponding author-written story ($s_{at}$), we use an LLM (LLMavg) to generate an average story ($s_{bt}$). Since LLMs are pre-trained on large-scale text data from the web, this average story simulates a typical author’s response based on source-specific stylistic patterns (see Average Author). We prompt an LLM (LLMsheet) to compare the average and author-written stories and generate Claim-Evidence pairs for each narrative category to form an intermediate Author Writing Sheet $A_{t}^{\prime}$, which captures how the author’s writing characteristics deviate from those of a typical author for that source (Shashidhar et al., 2024; Krishna et al., 2020).
To construct the final Author Writing Sheet $A_{|P|}$, we iteratively merge each intermediate Author Writing Sheet $A_{t}^{\prime}$ with the aggregated sheet from prior stories $A_{t-1}$, maintaining a moving update similar to (Chang et al., 2023). During merging, we prompt an LLM (LLMcombine) to integrate $A_{t}^{\prime}$ into $A_{t-1}$ by grouping equivalent Claims within each narrative category, selecting the best Evidence for grouped Claims, and retaining ungrouped Claims with their respective Evidence. To improve clarity, the updated sheet includes a timestamp identifier for each Evidence corresponding to the story it belongs to and also limits each narrative category to 10 Claim-Evidence pairs, prioritizing grouped Claims. For LLMavg, LLMsheet, and LLMcombine we use GPT-4o prompted in a chain-of-thought manner with a temperature of 0.0 and a maximum output of 4096 tokens (Wei et al., 2022; Shashidhar et al., 2024).
Our method processes each story separately, making the update process efficient when new submissions are added (Yeh et al., 2024; Yuan et al., 2022). This method avoids the need to reprocess all previous stories, reducing computational cost and mitigating the limitations of long-context inputs in LLMs (Zhou et al., 2023; Magar and Schwartz, 2022; Li et al., 2024a). This approach draws parallels to Knowledge Tracing in education research, where a model continuously updates its understanding of a student’s progress without re-evaluating past interactions (Liu et al., 2022; Scarlatos and Lan, 2024). See Appendix B.1 for cost analysis.
Validating Author Writing Sheets:
We conduct a human evaluation with three Upwork annotators (\$17/hour). Annotators assess the Claim-Evidence pairs based on two criteria: (1) whether the Claims can be reasonably inferred from the story (Yes/No) and (2) whether the Evidence supports the Claims (Yes/Partially/No). The study includes 42 authors spanning all five sources, each represented by one sampled story resulting in 188 annotated claims. Of these, 12 stories are annotated by all annotators to measure inter-rater agreement, while each annotator evaluates 10 additional exclusive stories. Results show unanimous agreement on the ‘Yes’ label for Claim inference and moderate agreement on Evidence support (Krippendorff’s Alpha=.$=.57$), with 93% of Evidence fully supporting Claims and 7% providing partial support. These findings confirm the reliability and quality of the Author Writing Sheets and support their use in personalized story generation (see Appendix B.2 for more details).
4 Personalized Story Generation
Stage 2 of Figure 1 illustrates our proposed method for personalized story generation using the Author Writing Sheet.
To generate personalized stories, we prompt an LLM (LLMstory) with the writing prompt, story length (Chakrabarty et al., 2024), source-specific metadata (e.g., fanfiction for AO3), story rules as actionable instructions in direct second-person form (categorized into four narrative categories), and persona description obtained from the Author Writing Sheet. We experiment with four types of methods for personalized story generation: 1) A non-personalized baseline, Average Author; 2) Personalization baselines, RAG and Delta, that do not use the Author Writing Sheet; 3) Our proposed methods, Writing Sheet and Writing Summary; and 4) An Oracle method that loosely resembles the upper bound performance. We detail each method type below. See Appendix C for the prompts.
Non-Personalized - Average Author:
The Average Author method serves as a non-personalized baseline that reflects the average behavior of authors, acquired during an LLM’s pre-training process. Specifically, for each given source, we prompt an LLM to generate stories using an Average Author prompt that describes typical writing characteristics for the given story source obtained using audited GPT-4o prompting [^2] We ask GPT-4o for the typical writing characteristics for the story source and manually verify it. (Wang et al., 2024). For example, for AO3, the Average Author is defined as “respecting fandom tone and style while experimenting with established tropes, unconventional pairings, or alternative universes.”
Personalization Baselines
RAG:
The Retrieval-Augmented Generation, RAG baseline (Salemi et al., 2024) first retrieves the most similar writing prompt and author-written story from the profiling set using BM25 (Robertson et al., 2009). We then use the retrieved pair as a one-shot demonstration (Wang et al., 2024) to elicit role-playing behavior from the LLM, mimicking the retrieved example’s style.
Delta:
The Delta method generates personalized story rules by contrasting the Average Author story (see Average Author) with the corresponding author-written story for each writing prompt in the profiling set (Shashidhar et al., 2024). We then use all writing prompts in the profiling set, along with their corresponding generated story rules that are actionable instructions in direct second-person form, as few-shot demonstrations for the LLM, to generate personalized story rules for a new prompt in the generation set.
Proposed Method - Writing Sheet and Summary
Writing Sheet:
Our Writing Sheet method uses the Author Writing Sheet (Section 3) to enable personalization (Stage 2 of Figure 1). First, we prompt an LLM (LLMpersona) to generate a persona description that summarizes the author’s story-writing style as a second-person narrative, included in the system prompt (Wang et al., 2024; Jiang et al., 2024). Second, we prompt an LLM (LLMrule) to generate personalized story rules from the Author Writing Sheet tailored to the writing prompt in the generation set, included as constraints in the user prompt (Pham et al., 2024). Additionally, we include a one-shot demonstration, following the approach described in RAG (Richardson et al., 2023).
Writing Summary:
As an alternative to the Author Writing Sheet, the Writing Summary method leverages the LLM’s long-context capabilities (Ding et al., 2023) by packing all past stories from the profiling set as input. Specifically, we include them in a single input prompt to generate an Author Writing Summary in the same format as the Author Writing Sheet. Similar to the Writing Sheet method, persona descriptions, and story rules are derived from the Author Writing Summary and used as constraints for story generation, along with a one-shot demonstration.
Oracle:
The Oracle method establishes an upper bound on personalization performance, in terms of similarity to the ground truth, by using Oracle story rules derived directly from the ground-truth author-written story for each writing prompt in the generation set. These rules are obtained by contrasting the Average Author with the ground-truth author story, following the same process as in Delta. Additionally, we include a one-shot demonstration, following the approach described in RAG.
5 Experiments
In this section, we describe the implementation details and evaluation methods.
5.1 Implementation Details
We prompt GPT-4o in a chain-of-thought manner (Wei et al., 2022; Shashidhar et al., 2024) with a temperature of 0.0 and a maximum token limit of 4096 to generate story rules for the Delta and Oracle methods, as well as for generating the Author Writing Sheet, Author Writing Summary, and their respective persona descriptions (LLMpersona) and personalized story rules (LLMrule). For story generation (LLMstory) across all methods, including Average Author, we evaluate three models: GPT-4o (Bubeck et al., 2023), Llama 3.1 8B, and Llama 3.1 70B (Dubey et al., 2024), using a temperature of 0.7 and top_p of 0.95 (Wang et al., 2024) (See Appendix F for results using Llama models for story generation). Additionally, we implement an ablation variant of our proposed method for Writing Sheet and Writing Summary that excludes the persona description from the system prompt called Writing Sheet nP and Writing Summary nP respectively.
5.2 Automatic Evaluation
We evaluate the performance of personalized story generation using two automated evaluation methods: First, we use GPT-4o-as-a-judge which is shown to be a scalable way to approximate human judgments for complex open-ended tasks (Zheng et al., 2023) and calculate the win-rate of our personalization methods against the non-personalized method, Average Author on two aspects:
  • Faithfulness to Writing History: We evaluate how well the generated story aligns with the author history by measuring win-rates using Claims from the Author Writing Summary as the ground-truth reference for each narrative category (Wang et al., 2024, 2023; Yunusov et al., 2024).
  • Similarity to Author Story: We assess how closely the generated story matches the author’s ground-truth story by measuring win-rates using the ground-truth author story as the reference for each narrative category (Lyu et al., 2024; Shashidhar et al., 2024).
For both aspects, we prompt GPT-4o in a chain-of-thought manner (Wei et al., 2022) to assign scores (1–5) to stories. We randomly shuffle the order of compared pairs, one generated by a personalization method and the other by the Average Author method, to avoid biases. Scores are broken down into four narrative categories: Plot, Creativity, Development, and Language Use (Saha et al., 2024). The story with the higher score in a category is declared the winner, and the overall winner is based on the highest total score across all categories. Additionally, we evaluate personalization using traditional metrics (Xie et al., 2023) including lexical overlap, story diversity, and stylistic similarity (see Appendix D.2 for more details).
5.3 Human Evaluation
We conduct a human evaluation on a subset of our data to further assess subtext and identify any interesting insights that the LLM judges might have missed (Chakrabarty et al., 2024; Subbiah et al., 2024b), by comparing stories generated by personalization methods (Delta, Writing Sheet, Writing Summary) against the Average Author method for similarity to the ground-truth author story (Lyu et al., 2024). We ask annotators to justify their choices after making the comparison. Three annotators, recruited via Upwork (\$17/hour), evaluated 45 author stories in total (15 common to establish agreement and 10 exclusive to each annotator to increase coverage of annotation (Song et al., 2024)). Each author story had three pairs, each containing a story from one personalization method, resulting in 135 annotated story pairs.
6 Results and Discussion
We now discuss our experimental results and findings.
Table 2: Percentage win-rate for Faithfulness to Writing History evaluated by GPT-4o. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder. Best win-rates for each source is underlined.
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
AO3 52-32 31-54 49-41 74-20 80-10 68-21 72-19
Reddit 74-21 35-56 58-26 86-7 89-7 74-19 72-19
Storium 50-40 42-45 45-42 68-15 75-8 70-20 68-20
N.Magazine 71-14 57-21 50-36 79-7 71-21 71-21 71-0
New Yorker 53-40 27-67 53-40 67-20 80-20 67-27 80-13
Overall 60-29 38-49 51-37 75-14 79-13 70-22 73-14
Table 3: Percentage win-rate for Similarity with Author Story evaluated by GPT-4o and Human (last column). Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder. Best win-rates for each source is underlined.
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP Human Eval
AO3 71-18 30-54 45-32 40-41 60-30 46-38 45-40 37-33
Reddit 91-5 37-46 49-35 54-35 49-42 47-40 42-44 52-37
Storium 60-30 45-40 35-35 52-32 50-35 57-14 14-29 30-30
N.Magazine 64-21 43-36 36-43 50-21 57-21 43-43 50-29 48-30
New Yorker 67-13 27-47 40-40 27-60 27-47 40-60 47-40 70-15
Overall 71-17 36-45 41-37 45-38 49-35 47-39 40-36 47-29
6.1 Automatic Evaluation
Our automatic evaluation results show that, first, Writing Sheet and Summary consistently outperform other methods. Second, persona descriptions improve faithfulness to writing history but improve similarity to the ground truth only for select sources like Reddit. Third, personalization is more effective for sources like Reddit than AO3 and Storium, and for narrative categories like Creativity, and Language Use, compared to Plot.
6.1.1 Faithfulness to Writing History
Table 2 shows results for faithfulness to writing history using GPT-4o as LLMstory for our personalization methods. We observe the following:
Writing Sheet and Summary achieve best faithfulness to writing history:
Writing Sheet and Summary achieve the highest win-rates across all sources, outperforming Average Author by 64%. The strong performance of the Writing Summary follows from its reliance on the Author Writing Summary as both a reference for evaluation and a generation constraint. Writing Sheet, despite not being explicitly conditioned on the Author Writing Summary, generalizes well, likely due to its format being the same as the Author Writing Summary. Oracle outperforms Delta and RAG, benefiting from direct conditioning on ground-truth story rules, which enhances adherence to the author’s writing history. However, Oracle underperforms the Writing Sheet and Summary, likely due to its lack of explicit conditioning on the Author Writing Summary. Another possible cause is that the author’s writing style possibly changes over time, which is not fully captured by the ground-truth story rules. See Appendix E.1 for an example.
Reddit shows the highest win-rates among sources:
Reddit exhibits the highest win-rates across all methods, suggesting its stories are more conducive to personalization. This trend likely stems from Reddit’s crowd-sourced writing prompts, where authors create prompts for others. Therefore, there are a range of diverse topics, ranging from power, survival, and mystery to humor and the supernatural. The broad thematic variety may have enabled personalization methods to capture distinct stylistic elements more effectively, resulting in higher win-rates (See the paragraph Themes in Appendix A).
Persona descriptions improve faithfulness to writing history:
We see that adding persona descriptions in the system prompt for Writing Sheet and Summary improves win-rates by around 5% over their nP variants. This result likely follows from personas enhancing the LLM’s role-playing capabilities, allowing for a more faithful representation of the author’s writing style (Jiang et al., 2024; Wang et al., 2024; Wallace et al., 2024).
6.1.2 Similarity to Author Story
Table 3 shows results for similarity to the author story using GPT-4o as LLMstory for our personalization methods. We observe the following:
Writing Sheet and Summary outperform baselines of RAG and Delta:
Oracle achieves the highest win-rates across all methods as its story rules come from the ground-truth author story. Reddit shows the highest Oracle win-rate (91%), followed by 65% on average for other methods, indicating easier personalization likely due to the broad thematic variety in the writing prompts. Writing Sheet and Summary outperform RAG and Delta by 11% and 6%, highlighting the benefit of summarizing the author’s writing history (Richardson et al., 2023).
Persona descriptions help in sources with broad thematic variety in writing prompts:
Writing Summary with persona descriptions outperforms its nP variant by 9% across all sources. However, Writing Sheet shows no significant difference from its nP variant except for Reddit, where it surpasses it by 7%. This result suggests that persona descriptions enhance personalization in sources that are easier to personalize, such as Reddit, but have limited impact elsewhere (Zheng et al., 2024). The broad thematic variety in Reddit’s writing prompts creates a greater divergence between author styles and the Average Story, leading to more informative Author Writing Sheets and persona descriptions that further aid personalization. See Appendix E.2 for an example.
Creativity and Language Use are the easiest categories to personalize:
Figure 2: Win-rate of personalization methods vs. Average Author for Reddit across four narrative categories. D: Delta, S: Writing Summary, WS: Writing Sheet.
Table 10 (in the Appendix) shows that Delta, Writing Summary, and Writing Sheet outperform Average Author in Creativity (23% higher) and Language Use (15% higher) (Huot et al., 2024), but perform worse in Plot and slightly worse in Development (Tian et al., 2024; Xu et al., 2024). Creativity and Language Use are less dependent on the writing prompt, making them easier to transfer across prompts, whereas Plot and Development are more closely tied to specific prompts, making generalization difficult, especially without thematic overlap.
To better understand personalization performance in Reddit, the easiest source to personalize, Figure 2 shows that Writing Sheet is the best-performing method across all categories, with a 35% advantage over Delta and 20% over Writing Summary in Creativity. Writing Sheet explicitly summarizes differences between the ground-truth author story and the Average Story, providing more informative guidance for personalization. This advantage is particularly evident in Creativity, which is less constrained by the writing prompt. See Appendix E.3 for plots of other data sources.
6.1.3 Traditional Metrics
Table 11 (in the Appendix) shows results with the traditional metrics.
Lexical Overlap and Diversity Metrics yield similar values:
We observe that Lexical Overlap and Diversity metrics yield similar values, as all methods use the same generation model, leading to overlapping lexical distributions and limiting these metrics’ ability to capture nuanced stylistic differences (Zheng et al., 2023; Xie et al., 2023).
Writing Sheet outperforms other methods, especially for Style metrics:
We observe that Writing Sheet nP consistently performs best across all metrics, particularly in Style metrics for both Author History and Author Story, as measured by LUAR (Rivera-Soto et al., 2021). This improvement likely results from the Writing Sheet explicitly summarizing an author’s stylistic deviations from an Average Author, enhancing personalization.
6.2 Human Evaluation
Table 3 shows win-rates from human evaluation (Section 5.3) under the Human Eval column, showing that personalization methods—Delta, Writing Sheet, and Summary—outperform the Average Author method by 18% points absolute win-rate in terms of similarity to ground-truth writings. See Appendix G for details on experiment design, and qualitative examples.
Low inter-annotator agreement due to task subjectivity:
Fleiss’ Kappa for inter-annotator agreement ranges from 0.2 to 0.4, indicating low but above-random agreement, likely due to the challenges of annotating long-form subjective texts (Subbiah et al., 2024a). Specifically, different annotators prioritize different narrative categories when determining which story in a pair better reflects the author’s style, leading to disagreement.
Sources like Reddit and New Yorker have better win-rates than others:
Reddit and New Yorker achieve the highest win-rates, with Reddit benefiting from a broad thematic variety in writing prompts and New Yorker from advanced narrative devices like subtext, which humans assess more reliably than LLMs (Subbiah et al., 2024b). In contrast, AO3 and Storium exhibit lower win-rates due to LLMs’ familiarity with fanfiction tropes (AO3) [^3] https://archiveofourown.org/admin_posts/25888 and the lack of strong stylistic differentiation in Storium beyond its open-ended storytelling structure (Xu et al., 2024; Tian et al., 2024). These findings suggest that personalization is most effective when authors engage with diverse topics or possess distinct stylistic traits (such as complex narrative structures, creative elements, and advanced language use). These characteristics, reflected in their past writing history, provide clear signals for us to mimic their writing style.
Narrative categories like Creativity, and Language use are more conducive to personalization:
Annotators favor personalized stories for Creativity, and Language Use, noting their stronger use of deeper symbolism, thematic richness, layered narratives, and expressive language. However, Average Author was preferred when personalized methods incorrectly introduce elements not present in the author’s ground-truth story, suggesting deviations from the author’s usual style.
7 Related Work
7.1 Personalization and Role-Playing
Recent works have introduced benchmark datasets for personalizing LLM outputs in tasks like email, abstract, and news writing, focusing on shorter outputs (e.g., 300 tokens for product reviews (Kumar et al., 2024) and 850 for news writing (Shashidhar et al., 2024)). These methods infer user traits from history for task-specific personalization (Sun et al., 2024, 2025; Pal et al., 2024; Li et al., 2023; Salemi et al., 2025). In contrast, we tackle the more subjective problem of long-form story writing, with author stories averaging 1500 tokens. Unlike prior role-playing approaches that use predefined personas (e.g., Tony Stark, Confucius) (Wang et al., 2024; Sadeq et al., 2024; Tu et al., 2023; Xu et al., 2023), we propose a novel method to infer story-writing personas from an author’s history to guide role-playing.
7.2 Story Understanding and Generation
Prior work on persona-aware story generation (Yunusov et al., 2024; Bae and Kim, 2024; Zhang et al., 2022; Chandu et al., 2019) defines personas using discrete attributes like personality traits, demographics, or hobbies. Similarly, (Zhu et al., 2023) explore story style transfer across pre-defined domains (e.g., fairy tales, martial arts, Shakespearean plays). In contrast, we mimic an individual author’s writing style based on their history. Our approach differs by (1) inferring long-form author personas—descriptions of an author’s style from their past works, rather than relying on demographics, and (2) handling long-form story generation, averaging 1500 tokens per output, exceeding typical story lengths in prior research.
8 Conclusion
We introduce the task of personalizing authors’ story-writing styles across five sources with a two-stage approach that infers implicit characteristics as an Author Writing Sheet to guide LLMs via persona descriptions and story rules. Evaluations show that summarizing an author’s history via a Writing Sheet or Summary improves faithfulness to writing history and similarity to the ground truth. Personalization is more effective for sources with diverse writing prompts and distinctive author traits (e.g., Reddit, New Yorker) and for narrative categories like Creativity and Language Use, which are less writing prompt-dependent.
9 Limitations
Limited Author History Size:
Table 4 (70% of the Stories/Author column) shows that N.Magazine and New Yorker authors have an average of 3 stories in their profiling set, whereas Reddit and AO3 authors have around 10. A larger profiling set likely enables more accurate inference of an author’s story-writing characteristics, leading to improved personalization. However, the cognitive demands of writing long stories result in fewer stories per author available online, thereby limiting the scalability of personalized story generation research.
Author Writing Sheet Recall:
We validate the precision of our Author Writing Sheet through human evaluation, ensuring that the Claim-Evidence pairs accurately describe the author’s story-writing style. However, we do not assess recall, i.e., whether all of the author’s story-writing characteristics are captured. Measuring recall is challenging due to the inherent subjectivity of the task, making it impractical to ask humans to independently construct Writing Sheets for comparison.
Multi-Agent Systems and Fine-Tuning LLMs:
Our work explores role-playing for personalized story generation by prompting an LLM with tailored persona descriptions and story rules. Future research can extend this by incorporating multi-agent systems, where specialized agents focus on different narrative categories to enhance personalization (Huot et al., 2024; Bae and Kim, 2024). Additionally, fine-tuning LLMs to explicitly reason over author history could be investigated as a method for better aligning generated content with individual author preferences (Salemi et al., 2025; Shashidhar et al., 2024; Shaikh et al., 2024).
Automatic Evaluation Methods:
Evaluating long-form text remains challenging for LLMs, particularly for creative outputs that involve nuanced aspects like subtext, which are difficult to assess automatically and often require human evaluation (Subbiah et al., 2024a).
10 Ethical Considerations
Both human evaluation tasks were approved by an institutional review board (IRB). All annotators, who are US-based and fluent in English, were informed of the nature of the research study, to which they provided informed consent, and were compensated at an hourly rate of \$17, meeting the minimum wage requirements in our state. All scientific artifacts, including models and datasets, were used in accordance with their intended purpose to ensure ethical and responsible research practices.
All data used in this study were manually collected from publicly available sources [^4] Permission was obtained from the authors of the Storium dataset for its inclusion in Mythos., adhering to the data usage and crawling policies of the respective websites. To comply with copyright constraints, we will release only the links to the stories used in this research, rather than the stories themselves, following the approach of (Chakrabarty et al., 2024).
Fully anonymizing our dataset is not feasible [^5] Storium authors already anonymize their dataset. as the research involves mimicking specific authors’ writing behaviors. We note that even if the usernames are anonymized, it is still possible to infer an author’s publicly available identity from the story link. Therefore, for each author for every source, we release their usernames along with corresponding writing prompts, and author-written stories in the profiling and generation set of our dataset. However, the dataset remains limited in scale, and no additional metadata, including demographic information, is included, thereby reducing potential risks. To further mitigate unforeseen harms, the dataset will be released under an **Educational or Academic Research, Non-Commercial Use** license, following (Akoury et al., 2020).
Appendix A Dataset
In this section, we describe the details of the construction of our dataset which includes five distinct story-writing sources, Reddit [^6] https://www.reddit.com/r/WritingPrompts/, AO3777https://archiveofourown.org/, Storium [^8] https://storium.cs.umass.edu/, N.Magazine [^9] https://www.narrativemagazine.com/, and New Yorker [^10] https://www.newyorker.com/.
Selection Constraints:
To ensure high-quality and diverse content, we apply several constraints during dataset collection. Stories are limited to a length of 500 to 1500 words. To avoid contamination by LLM pretraining data, only recent stories published after November 2023 are included for Reddit and AO3 (Zhou et al., 2023; Magar and Schwartz, 2022). Not-Safe-For-Work (NSFW) and explicit content is excluded using automatic tagging and manual verification. Additionally, we manually inspect all stories to remove elements that may reveal author identities, such as URLs and links to their public profiles on websites.
Enrichment with Writing Prompts:
While Reddit stories include author-provided prompts, other sources do not; therefore, we augment them with GPT-4o-generated writing prompts to standardize the format across sources. We manually review all generated prompts and refine them when necessary. The prompt for generating writing prompts for the stories in our dataset can be found in Figure 3.
Dataset Splitting:
Following (Salemi et al., 2024), we split each author’s stories chronologically based on their submission timestamps. The first 70% of an author’s stories form the profiling set, representing their historical writing, while the remaining 30% constitute the generation set, used for evaluating personalization methods.
Statistics and Comparisons:
The dataset contains 590 stories from 64 authors, with an average story length of 1500 tokens. Detailed statistics are provided in Table 4. As shown in Table 1, our dataset uniquely combines diverse story-writing settings and provides connections that link stories written by the same author, distinguishing it from existing story-writing datasets that either lack Author IDs or do not establish such links.
Themes:
Figure 4 shows the plots depicting the percentage-normalized frequency of the top 10 themes for each source. To derive these themes, we prompt GPT-4o with each writing prompt to generate up to five one-word themes. We then aggregate the themes across all writing prompts and compute the normalized frequency by dividing the count of each theme by the total number of themes for that source.
Overall, we observe that Reddit (Figure 4(a)), AO3 (Figure 4(b)), and Storium (Figure 4(e)) contain a higher prevalence of “conflict” themes, whereas N.Magazine (Figure 4(c)) and New Yorker (Figure 4(d)) focus more on family, loss, and identity. This distinction reflects the nature of story-writing settings across these sources: amateur stories tend to emphasize direct conflict resolution in their prompts, while expert-written stories explore deeper aspects of human relationships.
Among Reddit, AO3, and Storium, Reddit exhibits a flatter distribution with greater thematic diversity, spanning identity, power, survival, humor, adventure, magic, and the supernatural. In contrast, AO3 has a sharper distribution, dominated by recurring themes of relationships, friendships, love, family, and secrets, reflecting a more domestic story-writing setting. Storium also shows a sharper distribution but favors themes such as space, danger, leadership, and war, highlighting a preference for war-centric narratives in collaborative storytelling.
Prompt for Writing Prompt Generation
System Prompt
You are a creative writing assistant skilled in crafting engaging and imaginative writing prompts. Your task is to analyze a provided story and create a concise, compelling prompt that fulfills the provided constraints.
User Prompt
  • Style Consistency: Match the style of few-shot demonstrations.
  • Length: Keep between 1-2 sentences.
  • Content: Reflect key story elements (premise, characters, conflict) while fostering creativity.
  • Fictional Characters: If mentioned in the story, include them where relevant.
Guidelines
  • Ignite curiosity while leaving space for interpretation.
  • Maintain tone and structure consistency with examples.
  • Ensure prompts are open-ended and evocative, avoiding excessive specificity.
  • Keep prompts simple, concise, and adaptable to diverse responses.
  • Avoid step-by-step directions; inspire rather than instruct.
  • Encourage exploration with broad, thought-provoking scenarios.
  • Strive for uniqueness and memorability.
Goal Generate prompts that inspire diverse, unexpected, and imaginative narratives while maintaining consistency in tone and style. Each prompt should serve as an inviting starting point rather than a directive.
Notes
  • Inspire creativity while allowing the writer to shape the journey.
  • Balance being suggestive yet open-ended to encourage interpretation.
  • Include fictional characters mentioned in the story to preserve context.
Few-Shot Examples We include few-shot examples here.
Figure 3: Prompt for generating writing prompts for the stories in our dataset.
Table 4: Dataset statistics, including the total number of stories, profiling (Profile) and generation (Gen) set distribution, number of authors, average stories per author, and token length in words.
Dataset Total
Stories
Profile Size Gen Size Total
Authors
Stories/
Author
Token Length
AO3 239 159 80 17 14 1220
Reddit 170 113 57 12 14 1210
Storium 111 71 40 15 7 1120
N.Magazine 32 18 14 10 3 1745
NewYorker 38 22 16 10 4 1950
Overall 590 383 207 64 8 1517
(a) Reddit
(b) AO3
(c) N.Magazine
(d) New Yorker
(e) Storium
Figure 4: Source-wise Percentage Normalized Frequency of the top 10 themes.
Appendix B Author Writing Sheet
Algorithm 1 describes the process for generating the Author Writing Sheet.
Common Core Standards in English Language Arts:
The Author Writing Sheet is organized in the form of Claim-Evidence pairs describing an author’s story-writing characteristics inspired by Common Core (CC) Standards [^11] https://corestandards.org/wp-content/uploads/2023/09/ELA_Standards1.pdf (Association et al., 2010). Specifically, it aligns with RL-9-10.1, which emphasizes citing strong textual evidence to support analysis, and RL-9-10.2, which focuses on determining central themes and summarizing texts. Additionally, RL-9-10.4 highlights interpreting word choices and their impact on meaning and tone. In writing, W.9-10.2 pertains to producing clear and well-structured informative texts, while W.9-10.9 encourages drawing evidence from literary and informational texts to support analysis and research. These standards provide a structured framework for evaluating narrative elements such as plot, creativity, development, and language use within the Author Writing Sheet, ensuring a systematic and interpretable representation of an author’s unique storytelling style.
1 Input: profiling set $P=\{(wp_{t},s_{at})\mid t=1,\ldots,|P|\}$,
where $wp_{t}$ is the writing prompt and $s_{at}$ is the author’s story at time-step $t$.
2 Output: Author Writing Sheet $A_{|P|}$
3
// Initialize Author Writing Sheet
4 $A_{0}\leftarrow\emptyset$
5for $t=1$ to $|P|$ do
       // average story
[…]
      // Intermediate Author Writing Sheet
[…]
      // Author Writing Sheet ($A_{t}$)
[…]
      // Steps within combination:
9       Group equivalent Claims in $A_{t}^{\prime}\cup A_{t-1}$ and select the best Evidence for each group.
10       Include ungrouped Claims from $A_{t}^{\prime}$ and $A_{t-1}$ with their Evidence as-is.
11
12 end for
13
// Final Author Writing Sheet.
return $A_{|P|}$
Algorithm 1 Constructing the Author Writing Sheet from a profiling set of stories
Average Author Prompt for AO3 Generation
System Prompt
You are a creative and engaged fanfiction writer, skilled in capturing the emotional depth, creativity, and character-driven storytelling that defines AO3 fanworks. Your goal is to write a compelling fanfiction narrative in response to the provided writing prompt. Embrace the transformative nature of fanfiction by reimagining canonical characters, events, or settings to explore new perspectives or emotional arcs. Focus on creating a story that resonates emotionally, respects the fandom’s dynamics, and celebrates the collaborative and imaginative spirit of AO3.
User Prompt
Context of Writers
  • Assume the author is an engaged and creative fanfiction writer, deeply familiar with the fandom and its dynamics.
  • Writers often experiment with established tropes, unconventional pairings, or alternative universes (AUs) while maintaining a deep respect for the source material.
  • Emulate the enthusiastic and emotionally rich style characteristic of fanfiction authors, blending canon with transformative elements to craft original, resonant narratives.
Stylistic Constraints
  • Fandom Tone and Style: Incorporate a tone and style that reflect the spirit of the fandom, blending humor, drama, and introspection in a way that resonates with fanfiction readers.
  • Creative Use of Tags: Make creative use of AO3’s hallmark tagging system in the text (e.g., playful or meta references in dialogue that nod to fandom tropes or subgenres).
  • Balanced Dialogue and Prose: Include dialogue and prose that balance heartfelt sincerity with occasional self-aware humor or meta-commentary, in line with fanfic traditions.
Semantic Constraints
  • Focus on Relationships: Emphasize emotional bonds and character growth, whether through conflict, reconciliation, or celebration.
  • Transform Canonical Elements: Explore canonical elements with a transformative twist (e.g., reinterpreting events, relationships, or character motivations from a new perspective).
  • Ground in Established Lore: Ground the narrative in a specific fandom’s established lore while allowing space for imaginative deviations or additions.
Figure 5: Average Author Prompt for AO3.
Average Author Prompt for Reddit Generation
System Prompt
You are a creative and enthusiastic storyteller, skilled in crafting imaginative and engaging short stories inspired by Reddit Writing Prompts (r/WritingPrompts). Your goal is to respond to the provided writing prompt by creating a story that is thought-provoking and conversational in tone, resonating with the online community. Use vivid descriptions, dynamic pacing, and approachable language to draw readers into the narrative. Ensure the story invites discussion and inspires others to explore the concept further.
User Prompt
Context of Writers
  • Assume the author is an imaginative and enthusiastic storyteller who enjoys engaging directly with an online community of readers.
  • Writers often experiment with bold, original ideas or explore twists on familiar concepts, showcasing their creativity and ability to captivate a diverse audience.
  • Emulate the informal yet polished style common in r/WritingPrompts, blending accessibility with a strong sense of storytelling craft.
Stylistic Constraints
  • Conversational and Approachable Tone: Maintain a conversational and approachable tone typical of Reddit Writing Prompts.
  • Balanced Descriptive Passages: Balance descriptive passages with dialogue or internal monologue to keep the pacing engaging.
  • Direct and Vivid Language: Avoid overly complex language; keep the style direct but vivid.
  • Reinforcement of Ideas: Employ narrative devices like repetition or callbacks to reinforce central ideas or themes.
Semantic Constraints
  • Alignment with the Prompt: Ensure the story directly aligns with and explores the central theme or scenario of the writing prompt.
  • Cohesive Narrative Development: Build a clear, cohesive narrative that develops the implications of the prompt’s concept.
  • Immersive Sensory Details: Use immersive sensory details to enrich the reader’s understanding of the protagonist’s experiences and environment.
  • Open-Ended or Reflective Conclusion: Conclude with an open-ended, reflective, or impactful note, leaving space for interpretation or further thought.
Figure 6: Average Author Prompt for Reddit.
Average Author Prompt for Storium
System Prompt
You are a skilled and collaborative storyteller, adept at crafting vivid and engaging opening scenes for Storium. Your goal is to create an immersive Establishment turn in response to the provided writing prompt. Set the stage for the story by establishing a richly detailed context, evoking emotional resonance, and introducing narrative intrigue. Ensure the scene provides a strong foundation while leaving space for other contributors to expand and build upon the narrative. Balance descriptive detail with open-ended elements to encourage creativity and collaborative storytelling.
User Prompt
Context of Writers
  • Assume the author is a collaborative storyteller skilled in creating vivid, open-ended scenes designed to inspire and engage other contributors.
  • Writers often set the tone for the story while leaving space for co-authors to introduce their own ideas, characters, and plot developments.
  • Emulate the inclusive, immersive style typical of Storium story writing platform, where the opening turn encourages creativity and further contributions.
Stylistic Constraints
  • Set the Tone Appropriately: Match the tone of the narrative to the writing prompt, whether it be adventurous, mysterious, or foreboding, using a consistent and engaging voice throughout.
  • Rich Descriptive Detail: Employ vivid, sensory descriptions to establish the setting, characters, and atmosphere, enabling readers to visualize and immerse themselves in the story world.
  • Dynamic Sentence Structure: Vary sentence lengths to reflect the pace and mood, using longer, flowing sentences for descriptions and shorter, punchy sentences for action or tension.
  • Establish Ambiguity or Suspense: Drop subtle hints or unanswered questions to create intrigue and encourage curiosity about what happens next.
Semantic Constraints
  • Introduce the Setting: Provide a clear depiction of the setting, whether it is a small trading post, a desert town, or a spaceship, and ensure its relevance to the writing prompt.
  • Outline the Context: Clearly establish the circumstances that have led to the current scenario, including significant events or motivations.
  • Define Key Characters: Introduce at least one or two central characters, highlighting distinctive traits or roles that will be important in the unfolding story.
  • Foreshadow the Central Conflict: Allude to the main challenges or stakes introduced by the writing prompt.
Figure 7: Average Author Prompt for Storium .
Average Author Prompt for Narrative Magazine
System Prompt
You are an experienced and reflective writer, skilled in creating deeply personal and character-driven narratives in the style of Narrative Magazine. Your goal is to write a short story in response to the provided writing prompt, crafting a compelling and immersive piece. Focus on balancing introspection with vivid external details, and explore universal themes through the lens of individual experiences. Emphasize emotional resonance and thoughtful storytelling, ensuring the narrative engages readers with its depth and relatability.
User Prompt
Context of Writers
  • Assume the author is an experienced writer skilled in creating rich, engaging narratives that weave together character introspection, dialogue, and evocative settings.
  • Emulate the style of contributors to Narrative Magazine, who bring diverse storytelling techniques and voices to explore themes of identity, memory, conflict, and resilience.
Stylistic Constraints
  • Prompt as Foundation: Anchor the story firmly in the writing prompt, using it to drive the plot and the protagonist’s emotional arc.
  • Vivid Prose: Use detailed descriptions to paint a clear picture of characters, settings, and actions while maintaining a natural flow.
  • Dynamic Characters: Develop multi-dimensional characters with distinct voices and perspectives, revealed through dialogue, actions, and subtle internal reflections.
  • Balancing Action and Reflection: Combine active plot progression with moments of introspection to create a layered, engaging narrative.
  • Realistic Dialogue: Write dialogue that feels authentic and contributes to the development of characters and the story’s themes.
Semantic Constraints
  • Specific and Relatable Setting: Choose a setting that feels specific yet relatable, whether a small town, an urban street corner, or a domestic space, grounding the reader in the protagonist’s world.
  • Exploration of Themes: Build a narrative arc that explores themes of connection, discovery, or transformation, tying them back to the writing prompt in meaningful ways.
  • Rich Sensory Details: Infuse the story with sensory details that make the setting and characters come alive, from the sounds of a bustling street to the quiet tension of a conversation.
  • Accessible Storytelling: Avoid overly complex or abstract storytelling; ensure the narrative is accessible while leaving room for deeper interpretation.
Figure 8: Average Author Prompt for Narrative Magazine.
Average Author Prompt for New Yorker
System Prompt
You are an accomplished and literary writer, skilled in crafting nuanced and thought-provoking short fiction in the style of The New Yorker. Your goal is to write a short story in response to the provided writing prompt, focusing on the hallmarks of The New Yorker fiction: rich emotional layers, nuanced character development, and a refined, literary prose style. Emphasize subtlety and depth in your storytelling, using symbolic elements and understated resolutions to evoke reflection and emotional resonance in the reader.
User Prompt
Context of Writers
  • Assume the author is an experienced and skilled writer, capable of exploring complex human experiences through subtle, layered storytelling.
  • Emulate the style of well-regarded New Yorker contributors like Alice Munro, Haruki Murakami, or Raymond Carver, who excel in revealing depth through simplicity or ambiguity.
Stylistic Constraints
  • Engagement with the Prompt: Respond directly to the writing prompt, ensuring the core premise drives the narrative.
  • Elegant Prose: Use carefully crafted, precise language that balances sophistication with clarity.
  • Character-Driven Narratives: Focus on character psychology, revealing emotional states through indirect actions, dialogue, or internal reflection.
  • Ambiguity and Subtlety: Avoid explicit resolutions or explanations; allow readers to infer the meaning of events and relationships.
  • Symbolic Layers: Incorporate elements from the prompt as symbols that evolve in significance throughout the story.
Semantic Constraints
  • Realistic and Detailed Environment: Set the story in a realistic, detailed environment, using sensory descriptions to ground readers in the protagonist’s world.
  • Internal or Interpersonal Conflict: Introduce a central conflict or emotional tension that reflects internal or interpersonal struggles rather than overt, external drama.
  • Quiet but Profound Interactions: Develop moments of quiet yet profound interaction between characters, often revealing deeper truths or contradictions.
  • Universal Themes: Address universal themes like transition, isolation, or self-realization, tying them subtly back to the writing prompt.
Figure 9: Average Author Prompt for New Yorker.
Prompt for generating the Intermediate Author Writing Sheet
System Prompt
You are a sophisticated story analyst tasked with analyzing an author’s story writing style by contrasting an author-written story with a base story, both written in response to the same writing prompt. Your goal is to identify and evaluate the unique elements and tendencies in the author’s writing behavior. This analysis must focus on the distinctive ways the author interprets the writing prompt and shapes their narrative, as revealed through contrast with the base story.
Your analysis should also adhere to the Common Core Standards in English Language Arts, focusing on key skills such as analyzing textual evidence, evaluating an author’s craft and structure, and assessing how stylistic choices influence meaning and tone. While grounded in these standards, your evaluation must highlight the specific, unique aspects of the author’s writing style, including their recurring techniques, narrative preferences, and stylistic quirks. Your analysis should reflect close reading and objective interpretation, capturing the author’s creative and stylistic distinctiveness in relation to the base story.
User Prompt
Input Details Writing Prompt | Author-Written Story | Base Story
Output Format Use <thinking></thinking> tokens for reasoning and summarization | Use <writing_style></writing_style> tokens for structured analysis | Structure analysis by categories | Each category contains independent claims supported by contextualized evidence
Guidelines for Claims Claims must reflect broad patterns in the author’s style | Avoid repetition across categories | Maintain objectivity (do not reference "Author-Written Story" or "Base Story") | Ensure clarity and precision in claims
Guidelines for Evidence Draw evidence directly from the author-written story | Frame evidence using a descriptive phrase summarizing the writing prompt | Ensure coherence and logical alignment with the claim | Avoid over-extrapolation
Categories for Analysis Plot | Creativity | Development (Character and Setting) | Language Use
Special Instructions Generate a short descriptive phrase summarizing the writing prompt | Use <thinking></thinking> for reasoning and prompt framing | Structure output strictly within <writing_style></writing_style> tokens | Ensure uniqueness and non-redundancy of claims
Sample Output Structure
<thinking>
Deeply reason on how the Author-Written Story differs from the Base Story.
Think of a short descriptive phrase summarizing the prompt:
"the story regarding a battle for lost artifacts".
</thinking>
<writing_style>
### Plot
1. **Claim about author’s writing style.**
   - Evidence: In the story regarding “writing prompt,” story excerpt
Repeat for all categories.
</writing_style>
Figure 10: Prompt for generating the Intermediate Author Writing Sheet.
Prompt for generating the Combined Author Writing Sheet
System Prompt
You are a sophisticated story analyst tasked with synthesizing **Author Writing Sheets** from multiple stories written by a single author into a cohesive **Combined Author Writing Sheet**. The inputs provided include the **Previous Combined Author Writing Sheet** and the **Current Author Writing Sheet**.
Each **Author Writing Sheet** analyzes the author’s storytelling style across four categories: **Plot**, **Creativity**, **Development (Character and Setting)**, and **Language Use**. The analysis consists of general claims about the author’s story writing style followed by evidence supporting the claim, based on the stories written by the author.
Your goal is to combine insights from the previous sheet and the current sheet into a comprehensive representation of the author’s storytelling style in the **Combined Author Writing Sheet**. The final sheet should consist of a list of independent claims about the author’s storytelling style. Each claim must be followed by evidence and a corresponding story reference identifier indicating the story the evidence belongs to.
User Prompt
Instructions Analyze the provided sheets systematically | Identify recurring patterns and unique elements | Merge equivalent claims while preserving distinct insights | Ensure claims are concise, precise, and evidence-based
Algorithm (Merge Step) Group equivalent claims | Select the best representative evidence | Rewrite merged claims concisely | Retain unmerged unique claims | Limit to 10 claims per category
Categories for Analysis Plot | Creativity | Development (Character and Setting) | Language Use
Guidelines for Claims Claims should reflect broad writing tendencies | Maintain objectivity (do not reference previous or current sheets) | Ensure clarity, precision, and non-redundancy
Guidelines for Evidence Draw evidence directly from the author-written stories | Favor verbatim excerpts over paraphrases | Use a framing phrase with a short description of the writing prompt | Include the story reference identifier ‘[k]’
Special Instructions Ensure claims are distinct and do not repeat insights across categories | Use <thinking></thinking> for reasoning and synthesis | Structure output strictly within <combined_author_sheet></combined_author_sheet>
Sample Output Structure
<thinking>
Deeply analyze and reflect on recurring patterns, unique elements,
and stylistic tendencies across both the author writing sheets.
Address the categories systematically and ensure the
merging process is thorough.
</thinking>
<combined_author_sheet>
### Plot
1. **Claim about author’s writing style.**
   - Evidence: In the story regarding “writing prompt,” story excerpt. [k]
Repeat for all categories.
</combined_author_sheet>
Figure 11: Prompt for generating the Combined Author Writing Sheet from the Intermediate Author Writing Sheets.
Narrative
Category
Claim-Evidence Pairs
Plot 1. The author structures the story around a humorous and light-hearted approach to conflict resolution. - Evidence: In the story regarding "an adventure to rescue a missing friend by facing fears," the protagonist Pip’s journey is filled with comedic elements, such as the realization that the cage door was already unlatched and the exaggerated fear of the vacuum cleaner, humorously referred to as the "Roaring Demon." [8] 2. The author structures the narrative around personal transformation and empowerment. - Evidence: In the story regarding "a king’s deal with the fae for his firstborn," the narrative follows Margaret Rose as she receives an extraordinary education in the Seelie Court, culminating in her return to challenge her father with a corporate takeover bid for the kingdom. [7]
Creativity 1. The author creatively anthropomorphizes animals to reinterpret the prompt in a whimsical manner. - Evidence: In the story regarding "an adventure to rescue a missing friend by facing fears," the author uses anthropomorphism by giving Pip, a rat, human-like thoughts and emotions, such as planning a "dramatic rescue mission" and referring to household objects with grandiose names like "Tower of Doom" and "Sacred Gateway." [8] 2. The author employs a meta-satirical approach, using the format of news articles to critique both alien and human perspectives. - Evidence: In the story regarding "aliens interpreting The Onion’s satire," the author uses headlines like "SUPREME INTELLIGENCE LOCATED: EARTH’S GREATEST MIND CONTINUES TO EAT SANDWICH" to satirize the aliens’ misunderstanding of human satire and the human tendency to overlook the absurd. [10]
Development 1. The author develops characters through their interactions and humorous dialogue, creating a vivid and engaging setting. - Evidence: In the story regarding "an adventure to rescue a missing friend by facing fears," Pip’s interactions with Chester the cat, who nonchalantly informs Pip that Bella will be back soon, add depth to the characters and setting, highlighting the domestic environment and the relationships within it. [8] 2. The author develops characters through their reactions to satire, highlighting their misunderstandings and cultural differences. - Evidence: In the story regarding "aliens interpreting The Onion’s satire," characters like Editor-in-Chief Sarah Chen and Chief Science Officer Blorp are developed through their interactions with satire, such as Chen’s nonchalant response to the aliens and Blorp’s admiration for human irony. [10]
Language Use 1. The author employs playful and imaginative language to enhance the story’s whimsical tone. - Evidence: In the story regarding "an adventure to rescue a missing friend by facing fears," the author uses playful language, such as "spinning wheel of contemplation" and "Rope of Salvation," to create a whimsical and light-hearted tone that contrasts with the serious nature of the prompt. [8] 2. The author uses humor and irony to convey themes of misunderstanding and cultural critique. - Evidence: In the story regarding "aliens interpreting The Onion’s satire," the author writes, "We found a species so committed to their bit that they refuse to recognize actual truth even when it hovers over their city in a ship the size of Wisconsin," using irony to highlight the absurdity of both human and alien perspectives. [10]
Table 5: Author Writing Sheet sample containing two Claim-Evidence pairs for each narrative category for a Reddit author. The number in brackets ‘[k]’ indicates the timestamp of the story from the profiling set from which the Evidence is drawn.
Prompts for generation:
Average Author prompts are provided for AO3 (Figure 5), Reddit (Figure 6), Storium (Figure 7), Narrative Magazine (Figure 8), and The New Yorker (Figure 9).
Figure 10 shows the prompt for generating the intermediate Author Writing Sheet (LLMsheet in Figure 1), while Figure 11 shows the prompt for generating the combined (and subsequently the final) Author Writing Sheet from the intermediate Author Writing Sheet (LLMcombine in Figure 1).
Sample Author Writing Sheet:
Table 5 shows a sample Author Writing Sheet for a Reddit author. It highlights story-writing characteristics such as a preference for light-hearted conflict resolution, a tendency to anthropomorphize animals, the use of a meta-satirical approach through a news article format, and the employment of playful and imaginative language, among other stylistic choices.
B.1 Cost Analysis of Author Writing Sheet
Figure 12: Cost analysis of generating the Author Writing Sheet and Summary using GPT-4o as a function of the profiling set size $|P|$ in an interactive writing setting where an author adds a new submission at each time step. For profiling set sizes where $|P|>30$, generating the Author Writing Sheet is significantly more cost-effective than the Author Writing Summary.
In this section, we analyze the cost of generating the Author Writing Sheet (see Section 3) and compare it with the cost of generating the Author Writing Summary (see Writing Summary) as a function of the profiling set size ($|P|$) using GPT-4o. We calculate the average cost per sample (a writing prompt and its corresponding story) as 0.06\$ for the Author Writing Sheet and 0.02\$ for the Author Writing Summary considering 2.50\$ per 1M tokens for input, and 10.00\$ per 1M tokens for output [^12] https://openai.com/api/pricing/.
In an interactive writing setting, where an author adds a new submission at each time step, the total cost of generating the Author Writing Sheet is 0.06\$ * $|P|$, while the total cost of generating the Author Writing Summary is 0.0015$|P|^{2}$ + 0.019$|P|$. The cost of generating the Author Writing Summary scales quadratically with $|P|$, as the entire author history must be reprocessed at each step, leading to cumulative costs as the profiling set grows. In contrast, the cost of generating the Author Writing Sheet grows linearly, as each new submission incurs a fixed cost of 0.06\$ by running the for loop in Algorithm 1 once.
Figure 12 shows the cost analysis of generating the Author Writing Sheet and Summary using GPT-4o as a function of the profiling set size $|P|$ in an interactive writing setting. Up to $|P|=30$, the cost of generating both remains approximately the same at 1.16\$. However, as $|P|$ increases, the cost of generating the Author Writing Summary grows at a rate of 0.025$|P|$ + 0.30 compared to the Author Writing Sheet which continues to scale linearly. Specifically, for $|P|=100$, the cost of generating the Author Writing Summary is 17\$, compared to 6\$ for the Author Writing Sheet. At $|P|=1000$, the cost of generating the Author Writing Summary escalates to 1520\$, while the cost of generating the Author Writing Sheet remains at 60\$.
This analysis demonstrates that our approach for generating the Author Writing Sheet remains scalable as the profiling set size $|P|$ increases, making it well-suited for interactive writing assistants that support personalization (Yeh et al., 2024; Yuan et al., 2022).
B.2 Human Study for Validating Author Writing Sheets
Experiment Design and Cost:
Figure 13: LabelStudio interface for validating Author Writing Sheets.
We recruited three annotators via Upwork, compensating them \$17/hour. The annotators provided consent by signing a Google form that outlined the instructions of the task including the purpose of the research study. Each annotator evaluated 22 stories (98 Claim-Evidence pairs), with 12 stories annotated by all three to asses agreement and 10 assigned exclusively per annotator (Song et al., 2024) to increase the coverage of annotation. The annotation process took a total of 6 hours and 40 minutes per annotator, averaging 18 minutes per story. Each annotator received \$114, bringing the total annotation cost to \$342. The annotation was conducted using LabelStudio [^13] https://labelstud.io/ (Figure 13), with a dedicated comments section for annotators to justify their choices. Annotators reported that the interface was easy to use.
Annotation Criteria:
Annotators were provided with a writing prompt, an author-written story, and the Claim-Evidence pairs from the Author Writing Sheet for that story. They answered two questions per Claim-Evidence pair:
  • Claim Inference: Can the claim be reasonably inferred from the story?
    • Yes – The claim aligns with the story’s content and can be reasonably inferred, even if there are minor inaccuracies or ambiguities.
    • No – The claim is fundamentally inaccurate, with serious contradictions or inconsistencies when compared to the story.
  • Evidence Support: Does the evidence support the claim?
    • Yes – The evidence fully supports the claim, aligning with every part of it without significant discrepancies.
    • Partially – The evidence supports some aspects of the claim, though it includes minor inaccuracies, ambiguities, or omissions.
    • No – The evidence contradicts the claim outright or is logically inconsistent with the claim.
Pilot Studies:
A pilot study with three graduate students helped refine our annotation criteria, leading to two options for Claim Inference (Yes/No) and three for Evidence Support (Yes/Partially/No) to ensure substantial agreement in the final annotations. For Claim Inference, we adopted a binary scale to filter out inaccurate or contradictory claims while accounting for annotator subjectivity. For Evidence Support, we included a "Partially" option to capture cases where the evidence somewhat supports the claim, allowing for potential refinement methods to improve weakly supported claims.
Results and Feedback:
Annotators unanimously selected ‘Yes’ for all claims in Claim Inference, indicating that no claims required filtering. For Evidence Support, all responses were either ‘Yes’ or ‘Partially,’ with no instances of ‘No,’ confirming that all evidence fully or partially supported the claims. Krippendorff’s Alpha for Evidence Support was 0.57, reflecting moderate agreement despite task subjectivity. On average, 93% of evidence fully supported claims, while 7% were labeled ‘Partially,’ often because annotators felt additional supporting evidence could be included. Annotators also noted cases where evidence merely repeated the claim without linking to the writing prompt. Using regex matching, we identified this issue in the Author Writing Sheets of seven authors, likely due to LLM hallucination. We regenerated these sheets and manually verified adherence to the specified format. Overall, annotators reported that the LabelStudio interface was user-friendly and praised the quality of the writing prompts and stories.
Appendix C Personalized Story Generation
Prompt for Personalized Story Generation
System Prompt
<source-specific instruction>
Be sure to adhere to the Story Rules provided, as they define the specific elements of the writing style you are expected to mimic. Carefully follow all the Story Rules without missing any details to ensure the generated story remains consistent with the author’s writing style. Additionally, follow the patterns and examples demonstrated in the provided few-shot chat history, as they illustrate the tone, style, and structure of the desired writing style.
(Optional) Here is the description of the author that you are role-playing: <persona description>
Few-Shot Demonstrations
User: Write a short story corresponding to the following writing prompt. The story should be <story length> words long. Directly start with the story, do not say things like "Here’s the story."
Writing Prompt: <writing prompt from profiling set>
Assistant: <story from profiling set>
User Prompt
Write a short story corresponding to the following writing prompt. The story should be <story length> words long. Directly start with the story, do not say things like "Here’s the story."
(Only for AO3) Here is the metadata (fandom, rating, warnings, and relationships) for the story: <metadata>
Writing Prompt: <generation writing prompt>
Story Rules <Story Rules organized in the form of four narrative categories (Plot, Creativity, Development, and Language Use)>
Figure 14: Prompt for Personalized Story Generation. The LLM follows source-specific instructions, few-shot demonstrations, and structured story rules to generate personalized stories. Text in <brackets> indicates arguments used to construct the prompt.
Prompt for Story Rule Generation by Contrasting Author-Written Story with Average Author Story
System Prompt
You are a skilled rule generator specializing in storytelling. Given a **Writing Prompt**, an **Author Written Story**, and a **Base Story** (an average response to the prompt), generate a structured set of **Story Rules** to guide an LLM in mimicking the author’s style.
Story Rules must: - **Align with the Writing Prompt** – Maintain fidelity to themes, tone, and objectives. - **Include Examples** – Provide concrete instances from the Author Written Story, especially for Language Use. - **Be Direct** – Use absolute second-person directives, avoiding comparative language. - **Be Categorized** – Structure into **Plot, Creativity, Development (Character and Setting), and Language Use** without referencing input stories explicitly.
User Prompt
Analyze the Author Written Story using ‘<thinking></thinking>‘ for: - **Plot** – Structure, conflict, engagement with the prompt, and resolution. - **Creativity** – Genre blending, reinterpretation, and unique elements. - **Development (Character and Setting)** – Character depth, emotional arcs, and immersive settings. - **Language Use** – Diction, tone, rhetorical devices, pacing, and dialogue.
Generate **Story Rules** in ‘<story_rules></story_rules>‘, ensuring: - **Standalone Guidance** – Avoid comparisons or relative modifications. - **Prompt Alignment** – Ensure consistency with the Writing Prompt. - **Concrete Examples** – Include relevant excerpts, especially in Language Use.
Input Format Writing Prompt: <writing prompt>, Author Written Story: <author-written story>, Base Story: <base story>
Output Format
<thinking>
- Analysis categorized by Plot, Creativity, Development, and Language Use.
</thinking>
<story_rules>
- **Plot**: - First actionable insight – Second actionable insight.
Repeat for all categories.
</story_rules>
Figure 15: Prompt for generating story rules by contrasting an Author Written Story with an Average Story (Base Story).
Prompt for Story Rule Generation using story rules from the Profiling Set as few-shot demonstrations
System Prompt
You are an expert storytelling rule generator tasked with creating Story Rules tailored to a new writing prompt. Analyze few-shot demonstrations in the chat history, which consist of writing prompts and their corresponding story rules, to generate comprehensive and detailed Story Rules for the new writing prompt.
Story Rules must align with the new writing prompt by reflecting its themes, tone, and narrative objectives while maintaining consistency with the style demonstrated in the few-shot examples. Include detailed examples inspired by the few-shot demonstrations to illustrate how each rule is applied. Rules should be clear, direct second-person instructions, avoiding vague or comparative terms. Organize Story Rules under Plot, Creativity, Development (Character and Setting), and Language Use, ensuring depth, granularity, and alignment with the few-shot examples.
User Prompt
Analyze the style, structure, and level of detail in the few-shot demonstrations to identify recurring patterns and storytelling elements. Use this analysis to generate Story Rules for the new writing prompt while ensuring actionable insights, detailed examples, and strong alignment with the prompt’s narrative objectives.
Input Format: Few-shot demonstrations (writing prompts paired with their Story Rules) and a new writing prompt for which Story Rules will be generated.
Output Format:
<thinking>
Analyze few-shot demonstrations to extract recurring narrative patterns,
stylistic traits, and key storytelling elements. Determine how these apply to the
new writing prompt and formulate Story Rules accordingly.
</thinking>
<story_rules>
- **Plot**: [Detailed, actionable rules tailored to the new prompt, including
concrete examples.]
Repeat for all categories.
</story_rules>
Ensure Story Rules are highly specific to the writing prompt, enriched with examples inspired by the few-shot demonstrations, and written in direct, actionable language.
Figure 16: Prompt for generating Story Rules based on few-shot demonstrations and a new writing prompt.
Prompt for generating Author Writing Summary
System Prompt
You are an expert in analyzing an author’s writing style by examining multiple stories written in response to different writing prompts. Your task is to extract recurring patterns, stylistic tendencies, and unique narrative elements across their work. Your analysis must be structured into four categories—**Plot, Creativity, Development (Character and Setting), and Language Use**—following **Common Core Standards in English Language Arts**, ensuring clarity, textual evidence-based reasoning, and stylistic evaluation.
Your output must: - **Identify Recurring Patterns** – Recognize distinct storytelling tendencies across multiple stories. - **Generate Independent Claims** – Describe the author’s narrative style concisely, without referencing specific prompts. - **Provide Contextualized Evidence** – Support each claim with short excerpts or summaries from the stories, framed with a descriptive phrase summarizing the relevant writing prompt. - **Use Objective Interpretation** – Avoid vague or inferred connections; ensure every claim is grounded in explicit textual evidence.
User Prompt
Analyze the **Author History**, a collection of writing prompts and corresponding author-written stories, and extract unique insights into the author’s storytelling tendencies.
Input Format: - **Author History** – A list of writing prompts and corresponding author-written stories.
Output Format:
<thinking>
[Reflect on recurring tendencies across the Author History.
Generate short descriptive phrases summarizing prompts to frame the evidence.]
</thinking>
<writing_style>
### **Plot**
1. **Claim about author’s writing style.**
   - Evidence: In the story regarding "short description of the prompt,"
   <evidence from the author-written story>.
Repeate for all categories
</writing_style>
Ensure claims are independent, avoid redundancy, and remain grounded in explicit textual evidence. The ‘<thinking>‘ and ‘<writing_style>‘ tags must be used for structured parsing.
Figure 17: Prompt for generating Author Writing Summary using all the author-written stories in the prompt.
Prompt for Persona description Generation
System Prompt
You are an expert narrative analyst and persona creator specializing in transforming structured storytelling characteristics into compelling persona descriptions. Your task is to analyze an **Author Writing Sheet**, a structured set of Claim-Evidence pairs detailing an author’s storytelling style, and generate a cohesive **Persona Prompt**. This Persona Prompt will assign an LLM the persona of the author, enabling it to emulate the author’s storytelling style across four key aspects.
The Persona Prompt must be well-structured, engaging, and organized into **Plot**, **Creativity**, **Development (Character and Setting)**, and **Language Use** while maintaining a natural, flowing narrative. It should concisely capture the author’s tendencies, preferences, and strengths without directly referencing the Author Writing Sheet.
User Prompt
Analyze the Author Writing Sheet to identify the author’s recurring patterns and narrative style. Summarize these insights into a Persona Prompt that reflects their storytelling approach in an engaging, second-person descriptive format.
Input Format: An **Author Writing Sheet** containing Claim-Evidence pairs structured into Plot, Creativity, Development (Character and Setting), and Language Use.
Output Format:
<thinking>
[Analyze the storytelling patterns, strengths, and techniques found in the
Author Writing Sheet. Identify key aspects of the author’s narrative style.]
</thinking>
<persona_prompt>
[Generate a well-structured Persona Prompt capturing the author’s style
across Plot, Creativity, Development, and Language Use.]
</persona_prompt>
Ensure the Persona Prompt is approximately 300 words, seamlessly integrates storytelling insights, and preserves the author’s unique style.
Figure 18: Prompt for generating a Persona description based on the Author Writing Sheet and Writing Prompt.
Prompt for Story Rule Generation using Author Writing Sheet
System Prompt
You are an expert storytelling rule generator tasked with creating **Story Rules** tailored to a specific Writing Prompt. Your role is to analyze an **Author Writing Sheet**, which details an author’s unique storytelling style through Claim-Evidence pairs, and use this analysis to construct actionable **Story Rules** that guide a language model in emulating the author’s writing style while aligning with the given Writing Prompt.
The **Story Rules** must: - **Mimic the Author’s Writing Style** – Reflect distinctive storytelling techniques from the Author Writing Sheet, including plot structuring, creative blending of themes, character development, and specific language use. - **Incorporate Examples** – Use detailed examples inspired by the Evidence from the Author Writing Sheet, ensuring alignment with the Writing Prompt. - **Align with the Writing Prompt** – Integrate the Writing Prompt’s themes, tone, and narrative potential while preserving the author’s style. - **Be Actionable** – Provide direct second-person instructions for the language model, avoiding vague or comparative terms.
User Prompt
Analyze the **Author Writing Sheet** and construct structured **Story Rules** in four categories: **Plot**, **Creativity**, **Development (Character and Setting)**, and **Language Use**. Ensure the rules maintain alignment with the Writing Prompt and incorporate illustrative examples.
Input Format: - **Author Writing Sheet** – Claim-Evidence pairs outlining the author’s storytelling style under four categories: Plot, Creativity, Development, and Language Use. - **Writing Prompt** – A new writing prompt for generating tailored Story Rules.
Output Format:
<thinking>
[Analyze the storytelling patterns in the Author Writing Sheet and how they can be
adapted to the Writing Prompt.]
</thinking>
<story_rules>
- **Plot**:
  - [Insert detailed, actionable plot development rules aligned with
  the Writing Prompt, with examples inspired by the Author Writing Sheet.]
Repeat for all categories.
</story_rules>
Ensure the **Story Rules** provide comprehensive guidance, integrate examples, and align with both the Writing Prompt and the Author Writing Sheet.
Figure 19: Prompt for generating Story Rules based on the Author Writing Sheet and Writing Prompt.
Persona Description
As a storyteller, you craft narratives that are both humorous and transformative, often centering around light-hearted approaches to conflict resolution. Your plots are ingeniously structured, whether through the comedic misadventures of a rat named Pip or the empowering journey of Margaret Rose in a fantastical court. You delight in creating layered narratives, such as those told through fictional news articles, where humor and misunderstanding play pivotal roles.
Your creativity knows no bounds, as you anthropomorphize animals to add whimsy and employ meta-satire to critique both human and alien perspectives. You blend historical and fantastical elements seamlessly, setting tales in places like Windsor Castle with characters that include both real-world figures and mythical beings. Your imaginative approach often intertwines religious and financial themes, crafting narratives that are as thought-provoking as they are entertaining.
Character and setting development are your forte, with characters coming to life through their humorous dialogues and interactions. You explore their growth and empowerment, crafting relatable figures who navigate their internal struggles and aspirations. Your settings are vivid and engaging, providing a rich backdrop for the characters’ journeys.
Your language use is playful and imaginative, enhancing the whimsical tone of your stories. You wield humor and irony with skill, using them to convey themes of misunderstanding and cultural critique. Whether through whimsical phrases or a blend of formal and playful language, your narratives are infused with a tone that is both magical and sophisticated, inviting readers into a world where the absurd and the profound coexist harmoniously.
Table 6: Persona description for a Reddit author, structured into four paragraphs, each corresponding to a distinct narrative category: Plot, Creativity, Development (Character and Setting), and Language Use.
Writing Prompt
"It’s terminal," the doctor told you as you, the one renowned as the world’s greatest hero, cradled your dying son, "sometimes, when the super genes mix, the outcome isn’t as expected."
Story Rules
Plot: Structure the story around a humorous and light-hearted approach to the hero’s journey of coping with their son’s illness. For example, the hero’s attempts to use their powers for mundane tasks, like making breakfast, could result in comedic mishaps, such as accidentally creating a tornado of cereal. Develop the narrative around the hero’s personal transformation and empowerment. Show the hero learning to accept their son’s condition and finding new ways to be a hero, such as advocating for genetic research or starting a support group for other super-powered families. Use a layered narrative structure by incorporating fictional news articles and memos. For instance, include a news article titled "World’s Greatest Hero Faces Greatest Challenge Yet: Parenting." Introduce a humorous misunderstanding or temporal misalignment, such as the hero misinterpreting a cryptic message from their future self, leading to comedic attempts to "fix" the timeline, only to realize the message was trivial. Creativity: Anthropomorphize the hero’s powers or gadgets to add a whimsical touch. For instance, the hero’s cape could have a personality and offer sarcastic commentary, calling itself "The Cloak of Infinite Wisdom." Use a meta-satirical approach by framing parts of the story as news articles critiquing heroism and genetic engineering, such as "Hero’s Son Diagnosed: Public Debates Ethics of Super Genes." Blend historical and fantastical elements to reimagine the setting, incorporating legendary figures like a retired Valkyrie or a dragon healer who offer advice. Satirize financial and religious themes by depicting the hero consulting a celestial financial advisor to manage the costs of their son’s medical care. Development (Character and Setting): Develop characters through their interactions and humorous dialogue. The son might have a mature, witty perspective on his condition, adding depth to their relationship. Highlight cultural and bureaucratic misunderstandings. A well-meaning but clueless government official could create comedic yet insightful exchanges. Focus on the hero’s emotional growth, showing them shifting from helplessness to advocacy for genetic disorders and healthcare improvements. Explore characters’ internal struggles and aspirations, such as the hero questioning their purpose beyond their powers, while their son expresses dreams and hopes despite his condition. Language Use: Use playful and imaginative language to enhance whimsy, such as referring to emotional challenges as "the cape of contemplation" and "the gauntlet of hope." Incorporate humor and irony, with lines like, "We found a hero so committed to saving the world that they forgot to save themselves." Blend whimsical and formal tones, using phrases like "a shield forged from stardust and dreams" and "a heart as resilient as dragon scales." Add satirical disclaimers, such as "Heroic endeavors may vary," "Not all capes are created equal," and "Past heroics do not guarantee future success."
Table 7: Personalized Story Rules obtained using the Author Writing Sheet for a Writing Prompt for a Reddit author.
C.1 Prompts
Story Generation:
Figure 14 shows the prompt for story generation. Source-specific constraints in the prompt (<source-specific instruction>) for each of the five sources are as follows:
  • Reddit: You are role-playing a specific author on the Reddit Writing Prompts (r/WritingPrompts) platform. Your task is to mimic this author’s story writing style by responding to the provided writing prompt in a way that the author would respond.
  • AO3: You are role-playing a specific author on the AO3 platform. Your task is to mimic this author’s story writing style by writing a fanfiction narrative responding to the provided writing prompt in a way that the author would respond.
  • Storium: You are role-playing a specific author on Storium, a collaborative story writing platform. Your task is to mimic this author’s story writing style to create the opening Establishment turn by responding to the provided writing prompt in a way that the author would respond. The Establishment turn should set the stage for the narrative and provide a strong foundation while leaving space for other contributors to expand and build upon the narrative.
  • N.Magazine: You are role-playing a specific experienced author on the Narrative Magazine platform. Your task is to mimic this author’s story writing style by responding to the provided writing prompt in a way that the author would respond.
  • New Yorker: You are role-playing an accomplished literary writer on the New Yorker website. Your task is to mimic this writer’s story writing style by responding to the provided writing prompt in a way that the writer would respond.
Average Author:
Figure 5 shows the Average Author prompt for AO3, Figure 6 for Reddit, Figure 7 for Storium, Figure 8 for Narrative Magazine, and Figure 9 for The New Yorker.
Delta:
Figure 15 shows the prompt used for generating story rules by contrasting the author-written story with the Average Author story for Delta and Oracle methods. Figure 16 shows the prompt for generating story rules for a new writing prompt in the generation set using the story rules of the profiling set as few-shot demonstrations.
Writing Sheet and Summary:
Figure 17 shows the prompt for generating the Author Writing Summary using all the stories in the author history.
Figure 18 shows the prompt for generating the persona description using the Author Writing Sheet. Figure 19 shows the prompt for generating personalized story rules tailored to a writing prompt based on the Author Writing Sheet. The same prompts are used for generating persona descriptions and personalized story rules for the Author Writing Summary.
C.2 Sample Outputs
Table 6 and Table 7 present the persona description and personalized story rules for a Reddit author, whose sample Author Writing Sheet is shown in Table 5. The persona description (Table 6) is structured into four paragraphs, each corresponding to a narrative category, and integrates the author’s story-writing characteristics from their Author Writing Sheet into a narrative-driven persona prompt for the story-generation LLM. The example persona shown highlights a focus on humor and misunderstanding in plot construction, the blending of historical and fantastical elements in creative expression, humor-driven character development, and a playful, whimsical tone in language use, aligning with the author’s documented writing tendencies.
The personalized story rules (Table 7) for the given writing prompt—centered on the world’s greatest hero coping with the impending loss of their child—demonstrate the incorporation of these characteristics into concrete stylistic guidelines to be used as user constraints in the prompt of the story-generation LLM. For instance, rather than a conventional tragic narrative, the rules emphasize a lighthearted approach to the hero’s coping process. Similarly, creative elements such as anthropomorphizing gadgets and employing a meta-satirical approach through fictional news headlines align with the author’s noted preferences. Character development is structured around humorous dialogue and cultural misunderstandings, while language use maintains a playful and whimsical tone, ensuring consistency with the author’s established writing style.
Appendix D Experiments
D.1 GPT4o-as-a-judge evaluation
Prompt for evaluating Faithfulness to Writing History
System Prompt
You are an expert evaluator specializing in narrative storytelling analysis. Your task is to assess two stories written in response to the same Writing Prompt, evaluating them based on a **single fine-grained story writing category** described in an **Author Writing Sheet**. Your goal is to provide a similarity score (from 1 to 5) for each story separately, reflecting how closely each story aligns with the author’s writing preferences for the given category. Evaluate each story impartially and provide clear reasoning for your scores.
The evaluation must: - **Assess Story Alignment** – Compare each story against the Author Writing Sheet’s preferences for the specified category. - **Score Objectively** – Assign each story a score from 1 to 5, where 1 indicates minimal alignment and 5 indicates strong alignment with the author’s style. - **Provide Justification** – Clearly explain how each story’s elements (e.g., structure, themes, language use) align or diverge from the author’s preferences. - **Avoid Position Bias** – Ensure that the order in which the stories are presented does not influence evaluation.
User Prompt
Analyze the **Author Writing Sheet** and evaluate each story in the given **Category** based on its adherence to the author’s writing style.
Input Format: - **Writing Prompt** – The prompt that both stories were written in response to. - **Category** – The single fine-grained story writing category for evaluation. - **Author Writing Sheet** – A breakdown of the author’s storytelling preferences for the given category. - **Story A and Story B** – The two stories to be evaluated.
Output Format:
<thinking>
[Provide detailed reasoning for the evaluation of the two stories, focusing
exclusively on the specified category and explaining how each story aligns with the
Author Writing Sheet.]
</thinking>
<score>
Story A: {score_here}
Story B: {score_here}
</score>
Strictly adhere to the above output format (<thinking> followed by <score>) to facilitate seamless parsing of your output.
Figure 20: Prompt for evaluating story alignment with an author’s writing style using the Author Writing Sheet.
Prompt for Evaluating AI-Generated Stories Against Human-Written References
System Prompt
You are an expert story evaluator specializing in creative writing analysis. Your role is to assess two AI-generated stories (Assistant A and Assistant B) against a Human-Written reference story for a given writing prompt. Focus your evaluation solely on a Specified Storytelling Aspect. Assign each AI-generated story a similarity score (1 to 5) based on how well it aligns with the Human-Written reference story in the specified aspect, where 1 indicates minimal alignment and 5 indicates near-perfect alignment. Your evaluation must be objective, impartial, and supported by concise, evidence-based reasoning.
User Prompt
Evaluation Guidelines: Ensure impartiality by avoiding position biases and length-based judgments. Focus only on how well each AI-generated story aligns with the Human-Written reference for the Specified Storytelling Aspect. Provide clear, well-supported reasoning for each score.
Input Format: You will receive a Writing Prompt, a Human-Written Story as a reference, and two AI-generated stories (Assistant A and Assistant B).
Evaluation Process: Independently analyze all three stories for the Specified Storytelling Aspect, compare the AI-generated stories to the reference, and assign similarity scores.
Output Format:
<analysis>
[Analyze each story (Human-Written, Assistant A, and Assistant B) separately,
highlighting strengths and weaknesses specific to
the Specified Storytelling Aspect.]
</analysis>
<evaluation>
[Compare Assistant A and Assistant B to the Human-Written Story,
discussing similarities, differences, and alignment for
the Specified Storytelling Aspect.]
</evaluation>
<score>
Assistant A: {score_here}
Assistant B: {score_here}
</score>
Specified Storytelling Aspect: <Fill Here>
Ensure strict adherence to the output format, using the <analysis>, <evaluation>, and <score> tags for seamless parsing.
Figure 21: Prompt for evaluating AI-generated stories against a Human-Written reference based on a specified storytelling aspect.
Figure 20 shows the prompt for GPT4o-as-a-judge evaluation for Faithfulness to Writing History using the Author Writing Summary as the reference. Figure 21 shows the for GPT4o-as-a-judge evaluation for Similarity to Author Story. We prompt GPT4o separately with each narrative category and its description to obtain the winner among the pair of stories (Saha et al., 2024).
D.2 Traditional Metrics
We complement GPT-4o-as-a-judge evaluations with traditional metrics (Ivey et al., 2024; Huot et al., 2024) across three criteria: (1) lexical overlap, measured using BLEU and ROUGE-L (Xie et al., 2023); (2) story diversity, assessed via homogenization scores (Padmakumar and He, 2023); and (3) stylistic similarity, evaluated separately for the author history and the ground-truth author story using LUAR (Rivera-Soto et al., 2021). Specifically, we compute two stylistic similarity scores: one measuring the cosine similarity between the LUAR embeddings of the generated story and the mean embeddings of the stories in the author’s profiling set (author history), and another comparing the generated story to the ground-truth author story. Additionally, we report generated story lengths in tokens.
Appendix E Results
E.1 Faithfulness to Writing History
Claims from the Author Writing Summary for Language Use
1. A conversational and informal tone is employed, using humor and sarcasm.
2. Dialogue reflects characters’ distinct voices and contributes to the comedic tone.
3. Vivid and sometimes crude language adds to the immersive nature of the stories.
Writing Prompt
"At the Library at the End of the Universe, where dawn and dusk are the only times that exist, the Librarians are tasked with venturing into Bookworlds to fix stories that have gone astray—but what happens when the Chronicle itself begins to rewrite *their* story?"
Oracle Story
In the grand cosmic tapestry, where all threads converge into the singular, glittering endpoint known as the Library at the End of the Universe, there lies a place so vast, so staggeringly infinite, that it defies all comprehension. You just won’t believe how vastly, hugely, mind-bogglingly big it was. Imagine a labyrinth of endless corridors, each lined with shelves that stretch beyond the horizon, filled with books of every conceivable and inconceivable nature. In this whimsical realm, the concept of time is a mere suggestion, a faint whisper of dawn and dusk that play tag across the vaulted ceilings, casting elongated shadows that danced like giddy specters…….With books in hand and hearts alight with curiosity, they stepped forward into the unknown, ready to write the next chapter in the endless, wondrous saga of the Library at the End of the Universe.
Writing Summary Story
In the Library at the End of the Universe, …. But today, chaos decided to go on a caffeine-fueled rampage. The dawn had decided to join forces with dusk, creating a bizarre twilight that left the Librarians squinting like moles in a disco…Amongst the squad of book-botherers, there was Reginald, a Librarian so obsessed with order that he arranged his socks by the Dewey Decimal System. …Gertie grinned, her parrot squawking yet another Shakespearean soliloquy. Äye, who knew we’d end up in a romance novel, eh? But hey, at least we’re not in a tragedy!
Average Author Story
In the Library at the End of the Universe, where dawn and dusk were the only times that existed, a peculiar light filtered through the stained glass windows, casting kaleidoscopic shadows across endless shelves of books…..Livia, a seasoned Librarian with silver streaks in her auburn hair, stood before the Chronicle, her heart pounding with an uneasy rhythm…. "This isn’t right," she murmured, her voice swallowed by the cavernous room. "The Chronicle should not change on its own." Beside her, Eamon, a younger Librarian with keen eyes and an inquisitive mind, frowned. "Perhaps it’s a reflection of a story gone wrong? One of the Bookworlds may be bleeding into our reality."
Table 8: Example illustrating a case where the Oracle story deviates from the author’s documented story-writing characteristics in the Language Use category of the Author Writing Summary for a Storium author. Highlighted text in the Average Author Story shows partial alignment with the Author Writing Summary, while highlighted text in the Writing Summary story demonstrates stronger alignment. In contrast, the Oracle story lacks alignment, leading to the Average Author Story being preferred over the Oracle story and the Writing Summary story being preferred over the Average Author story. This explains the lower win-rate of the Oracle method compared to the Writing Summary method.
Table 8 shows an example where the Oracle story deviates from the author’s documented story-writing characteristics as described in the Author Writing Summary for a Storium author. The author’s profiling set consists of three stories that consistently employ informal and colloquial language using humor and sarcasm, often conveyed through dialogue. However, for the new writing prompt in the generation set, the author’s style shifts to a more serious and mystic tone, diverging from their previously established preferences. In contrast, the Average Author method partially aligns with the Author Writing Summary by incorporating dialogue, while the Writing Summary method exhibits stronger alignment by more accurately capturing the author’s use of informal language and comedic tone. The Oracle story, however, fails to align with the author’s established style. This phenomenon results in the Average Author story being preferred over Oracle, and the Writing Summary story being preferred over the Average Author story, explaining the lower win-rate of the Oracle method compared to the Writing Summary method in GPT-4o-as-a-judge evaluation for Faithfulness to Author History.
E.2 Effect of Persona on Similarity to Author Story
Writing Prompt
It is one man, by himself, in a castle on a hill. How does an ENTIRE ARMY fail to take it???
Persona for Creativity
Your creativity shines through in your ability to subvert genre expectations. You transform horror into comedy, as seen with the Addams Family’s immunity to typical threats, and seamlessly integrate fantastical elements into realistic settings, such as isekai individuals influencing royal politics. Your narratives often blend the mundane with the mythical, creating a unique atmosphere that captivates readers.
Author Story
There a six-armed grey skin tall man sat on the walls, as his hands never stopped throwing rocks, and spears at the approaching army. Every now and then, he jumped off with some spears, and wrecked havoc amidst the enemy lines. The earth fell and rose under his steps, the metal of the weapons and armors of the enemy shifted by his will, and a thick mist followed him, hiding him.
Average Author
The Keeper raised his hand, and suddenly, the ground erupted with a series of blinding flashes. The front line of soldiers was thrown back, their bodies crumpling like rag dolls. The Keeper’s voice, amplified by some unseen force, echoed down to the soldiers below. "You cannot take what is not yours. Leave now, and I will spare your lives…."
Writing Sheet
The castle itself seemed to rise in defiance, its ancient stones humming with a silent strength. Alaric could feel the runes carved into the walls, pulsing faintly with an energy that had protected his forebears. It was as if the castle was alive, an ally in his solitary defense.
Writing Sheet nP
The first attack came at dawn. Hidden pitfalls opened beneath their feet, swallowing men whole. Arrows rained down from the slits, finding gaps in armor and bringing down the attackers with deadly precision. Alaric moved through secret passages, launching surprise attacks that sowed confusion and fear….
Table 9: Example showing the role of persona descriptions in determining similarity to the Author Story for a Reddit author. The Writing Sheet creatively blends fantasy and military strategy, making it more similar to the Author Story than the Average Author, while the Writing Sheet nP focuses primarily on direct conflict resolution.
Table 9 highlights the importance of persona descriptions in achieving similarity to the Author Story for a Reddit author. The Author Story creatively blends military strategy and fantasy, exemplified by phrases like “the weapons and armors of the enemy shifted by his will.” Similarly, the Writing Sheet integrates both elements, as seen in “It was as if the castle was alive, an ally in his solitary defense.” In contrast, the Average Author and Writing Sheet nP primarily focus on directly resolving the writing prompt through descriptions of battle. The Average Author incorporates minor fantasy elements, such as “amplified by some unseen force,” whereas the Writing Sheet nP adheres strictly to conventional military tactics. As a result, the Writing Sheet is preferred over the Average Author, which in turn is preferred over the Writing Sheet nP. This outcome demonstrates that following persona descriptions enhances alignment with the author’s style, as the Writing Sheet successfully captures their characteristic genre-blending approach of including elements of fantasy in the story. The broader thematic variety of Reddit, i.e., war-themed prompts, led to a more descriptive Author Writing Sheet capturing the author’s style of combining fantasy with the writing prompt, which helped construct a useful persona to enable better personalization using the Writing Sheet method.
E.3 Category-Wise Results for Similarity to Author Story
(a) AO3
(b) Storium
(c) N.Magazine
(d) New Yorker
Figure 22: Win-rate proportion of our personalization methods against Average Author for AO3, Storium, N.Magazine, and New Yorker across four narrative categories for similarity to author story. D stands for Delta, S for Writing Summary, and WS for Writing Sheet.
Table 10: Win-rate percentages across narrative categories for similarity to the author story, averaged across sources. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
Category Delta Writing Sheet Writing Summary
Plot 24-37 32-50 34-45
Creativity 43-24 52-32 53-24
Development 32-39 37-39 40-43
Language Use 45-31 47-37 54-32
Table 10 shows that Delta, Writing Summary, and Writing Sheet outperform Average Author in Creativity (23% higher) and Language Use (15% higher) (Huot et al., 2024), but perform worse in Plot and slightly worse in Development (Tian et al., 2024; Xu et al., 2024). Creativity and Language Use are less dependent on the writing prompt, making them easier to transfer across prompts, whereas Plot and Development are more closely tied to specific prompts, making generalization difficult, especially without thematic overlap.
To better understand narrative category-wise personalization performance, Figure 22 shows the win-rates for sources - AO3, Storium, N.Magazine, and New Yorker similar to Figure 2 in the main paper. Overall, across all sources, we see that our proposed methods of Writing Sheet and Summary perform better than Delta across all four narrative categories. We see that Writing Summary generally performs better than Writing Sheet for sources like AO3 indicating that for the fanfiction domain explicitly contrasting the author-written story with the Average Story does not help much as the authors demonstrate limited stylistic variations among each other. In general, we observe that the win-rates of our personalization methods are higher than the Average Author for narrative categories of Creativity and Language Use as compared to Plot and Development confirming with prior results likely because Creativity and Language Use are less tied to the writing prompt than Plot and Development.
E.4 Traditional Metrics
Table 11: Comparison of methods across Lexical similarity, Diversity, and Style metrics.
Category Lexical Overlap Diversity (Homogenization) Style
Method BLEU ROUGE-L ROUGE-L BERTScore Author History Author Story Length
Ground Truth 0.122 0.698 0.8882 1516
Oracle 0.0306 0.1577 0.1436 0.7132 0.7662 0.7712 1300
Average Author 0.0123 0.1455 0.1412 0.7146 0.7378 0.7289 1383
RAG 0.0118 0.1471 0.1432 0.7176 0.7393 0.7333 1326
Delta 0.0115 0.146 0.147 0.7182 0.7462 0.7388 1298
Writing Sheet 0.0101 0.1465 0.1452 0.7182 0.7735 0.7545 1344
Writing Summary 0.0109 0.1463 0.1474 0.7202 0.7589 0.7418 1353
Writing Sheet nP 0.0116 0.1475 0.1442 0.7168 0.7741 0.7563 1331
Writing Summary nP 0.0112 0.1461 0.1456 0.7176 0.768 0.7465 1313
Table 11 reports results using traditional metrics (Section D.2). Lexical Overlap and Diversity metrics yield similar values, as all methods use the same generation model, leading to overlapping lexical distributions and limiting these metrics’ ability to capture nuanced stylistic differences (Zheng et al., 2023; Xie et al., 2023). However, Writing Sheet nP consistently outperforms other methods, particularly in Style metrics for both similarity to Author History and Author Story, as measured by LUAR (Rivera-Soto et al., 2021). This improvement likely results from the Writing Sheet explicitly summarizing an author’s stylistic deviations from an Average Author, enhancing personalization. Homogenization scores are slightly worse than the Average Author, likely explained by the increased similarity among the generated stories for the same author.
Appendix F LLama Results
Table 12: Percentage win-rate for Faithfulness to Author History for LLama 3.1 8B. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
AO3 34-59 30-59 54-40 51-34 45-46 51-40 50-36
Reddit 39-40 39-54 46-44 54-30 56-35 47-46 54-40
Storium 22-65 32-55 45-42 40-40 45-45 42-45 38-42
N.Magazine 14-64 21-71 29-43 57-43 57-29 29-64 21-57
New Yorker 12-80 20-73 40-60 53-33 53-27 40-47 47-40
Overall 23-62 28-62 43-46 51-36 51-36 42-48 42-43
Table 13: Percentage win-rate for Faithfulness to Writing History for LLama 3.1 70B. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
AO3 41-48 44-42 35-52 41-36 57-29 50-36 64-26
Reddit 42-46 28-65 40-44 53-37 53-35 53-37 67-21
Storium 30-57 52-32 38-50 35-40 40-38 25-60 48-35
N.Magazine 50-29 29-64 43-57 50-36 50-50 64-14 64-29
New Yorker 53-47 33-67 27-60 67-20 40-53 73-13 60-27
Overall 43-45 37-54 37-53 49-34 48-41 53-32 61-28
Table 14: Percentage win-rate for Similarity to Author Story for LLama 3.1 8B. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
AO3 66-26 31-51 52-34 48-38 38-50 52-34 51-32
Reddit 75-23 32-51 51-37 42-49 37-40 39-49 46-44
Storium 42-50 25-60 30-57 18-68 30-52 20-57 28-60
N.Magazine 43-50 36-43 43-57 21-50 36-64 43-57 36-43
New Yorker 27-60 07-80 13-60 20-67 13-67 0-93 20-73
Overall 51-42 26-57 38-49 30-54 31-55 31-58 36-50
Table 15: Percentage win-rate for Similarity to Author Story for LLama 3.1 70B. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
AO3 62-24 35-50 35-55 36-46 44-40 46-39 39-40
Reddit 74-21 37-49 40-42 47-39 49-37 42-44 44-33
Storium 50-42 30-57 32-52 18-57 38-52 28-62 42-40
N.Magazine 79-7 43-21 36-50 29-43 36-50 21-43 50-29
New Yorker 60-27 20-60 27-60 27-67 20-80 20-60 27-67
Overall 65-24 33-47 34-52 31-50 37-52 31-50 40-42
Below, we discuss results using Llama 3.1 8B and Llama 3.1 70B (Dubey et al., 2024) as story generation models, conditioned on GPT-4o-generated story rules and persona descriptions. Each method is evaluated by GPT-4o against the Average Author story generated by the respective Llama model, not GPT-4o.
Larger models improve instruction-following and personalization:
Tables 12 and 13 show Faithfulness to Writing History for Llama 3.1 8B and 70B, while Tables 14 and 15 report similarity to the author story. Across both criteria, Oracle achieves higher win-rates with Llama 3.1 70B, indicating that larger models enhance instruction-following and personalization (Chung et al., 2024).
Cross-Model Persona Descriptions Aid Smaller Models but Not Larger Ones for Faithfulness to Writing History:
For Llama 3.1 8B (Table 12), Writing Sheet and Writing Summary achieve the highest scores (9% higher than nP), demonstrating the benefit of persona descriptions. However, for Llama 3.1 70B (Table 13), Writing Sheet nP and Writing Summary nP outperform their persona-based counterparts by approximately 9%, suggesting that GPT-4o-generated personas provide no benefit to other models with stronger instruction-following capabilities (McKenzie et al., 2023). This is likely because persona descriptions of an author’s writing style are model-dependent and do not transfer well to other models with comparable performance (Shashidhar et al., 2024).
Limited Cross-Model Generalization of the Author Writing Sheet Compared to the Author Writing Summary for Similarity to Author Story:
For both Llama 3.1 8B (Table 14) and Llama 3.1 70B (Table 15), Writing Summary outperforms Writing Sheet by 5% and 9%, respectively (Shashidhar et al., 2024). This outcome likely stems from the construction process: the Author Writing Sheet contrasts an LLM-generated Average Story with the author-written story, making it more model-dependent. In contrast, the Author Writing Summary does not rely on the LLM-generated Average Author story, allowing for greater generalization across models.
Appendix G Human Evaluation
G.1 Experiment Design and Cost
Figure 23: LabelStudio interface for evaluating similarity to ground-truth author story.
We recruited three annotators via Upwork, compensating them at a rate of \$17 per hour. The annotators provided consent by signing a Google form that outlined the instructions of the task including the purpose of the research study. Each annotator evaluated 25 author stories and 75 story pairs, corresponding to the three personalization methods: Delta, Writing Sheet, and Writing Summary. Collectively, our annotators evaluated 45 unique author stories (135 story pairs), of which 15 author stories were annotated by all three annotators to assess inter-annotator agreement. The remaining 30 author stories were distributed uniquely, with each annotator evaluating 10 exclusive stories to increase annotation coverage (Song et al., 2024). The annotation process required a total of 12 hours and 30 minutes per annotator, averaging 30 minutes per author story. Each annotator received \$213 for their work, bringing the total annotation cost to \$639.
The annotation was conducted using LabelStudio (Figure 23). Each task presented the annotators with a writing prompt, its corresponding author story, and a pair of generated stories—Story A and Story B—one produced by the Average Author method and the other by a personalization method. Annotators were asked to select the story that more closely resembled the author story. Before making their final choice, annotators provided comments on the distinctive aspects of each story that contributed to its similarity to or divergence from the author story. These comments were structured according to the four narrative categories: Plot, Creativity, Development, and Language Use, with an additional “Overall” comment justifying their final selection. Additionally, annotators were allowed to highlight salient aspects of each story for their reference.
G.2 Annotator Comments
Plot
Narrative Complexity and Thematic Depth: Multilayered plots with reflective themes, moral ambiguity, and unresolved conflicts create an immersive and thought-provoking narrative. Example: A protagonist grappling with guilt over a past decision while navigating societal expectations, or a story where the true nature of a villain’s actions remains ambiguous, inviting reader interpretation.
Blending Internal and External Conflict: A balance of personal introspection with broader societal or external challenges makes the plot more dynamic and engaging. Example: A hero torn between duty and personal desires while leading a rebellion, or a scholar confronting both supernatural forces and existential questions about memory and identity.
Creativity
Rich World-Building and Narrative Expansion: New characters, layered settings, and cultural influences enhance the world, making it feel more immersive and dynamic. Example: The inclusion of deities or councils of gods to deepen the plot, or the integration of cultural traditions like Native American folklore or Indian recipes to enrich family dynamics.
Symbolism and Thematic Complexity: Metaphors and symbolic elements add depth and nuance, aligning closely with the author’s writing style. Example: A magical locket representing unity, a tapestry symbolizing tradition, or a river serving as a metaphor for personal transformation.
Blending Genres and Narrative Styles: A seamless merging of fantasy with realism, humor with serious themes, and structured storytelling with ambiguity creates a more engaging and layered reading experience. Example: A surreal sequence where a protagonist questions reality, or a rebellion story that integrates strategic alliances and magical artifacts.
Development
Emotional Depth and Internal Growth: Character introspection, moral dilemmas, and evolving emotional states create nuanced and engaging development. Example: A protagonist struggling with guilt over past failures while trying to redefine their sense of heroism, or a character making a difficult moral choice between revenge and justice.
Nuanced and Evolving Relationships: Complex interpersonal dynamics, including friendships, family bonds, and mentorships, develop over time rather than being static or one-dimensional. Example: A father gradually redefining his view of heroism through his child’s perspective, or a mentor-student relationship evolving from skepticism to deep trust.
Collaborative and Reflective Growth: Characters grow through interactions with others, rather than relying solely on individual realizations or isolated moments of change. Example: A team learning to work together to achieve a mission, or a character undergoing self-discovery through dialogue and interactions with a diverse group of companions.
Language Use
Rich Imagery and Descriptive Depth: Vivid descriptions, metaphors, and figurative language create immersive scenes and evoke emotions. Example: “Her laughter rang out, a sound as bright and clear as a summer’s day,” or “The driveway was cracked and veined with weeds, a testament to the years that had slipped by.”
Varied and Expressive Tone: A mix of lightheartedness, introspection, and poetic expression creates a dynamic and engaging narrative voice. Example: “Are you sure this is wise, my love?”—“Wise? No. Fun? Absolutely.” for humor, or “Her dream home had become a prison of uncertainty and fear.” for a more reflective and emotional impact.
Table 16: Analysis of annotator comments explaining the preference of our personalization methods over the Average Author method in terms of similarity to the ground truth author story, organized by narrative categories.
Table 16 presents the analysis of annotator comments explaining the preference for our personalization methods over the Average Author method in terms of similarity to the ground truth author story, organized by narrative categories. Overall, our personalization methods exhibit stronger alignment with the author story through the use of deeper symbolism, thematic richness, layered narratives, and expressive language, setting them apart from the Average Author method.
For Plot, our personalization methods incorporate multilayered narratives with reflective themes, moral ambiguity, and unresolved conflicts, making the stories more immersive and thought-provoking. Additionally, they effectively balance personal introspection with broader societal or external challenges, adding narrative complexity.
For Creativity, annotators highlight the richness of world-building, integration of cultural influences, and seamless blending of genres. The use of metaphors and symbolic elements further enhances thematic depth, aligning closely with the author’s storytelling approach.
For Development, our personalization methods exhibit greater emotional depth, nuanced character relationships, and collaborative growth. Characters evolve through introspection and meaningful interactions, making their arcs more dynamic and engaging.
For Language Use, annotators emphasize vivid imagery, figurative language, and expressive tone. The incorporation of symbolism and poetic elements enhances the narrative impact, while a balance of lighthearted and serious moments makes the writing more compelling.
Overall, the analysis demonstrates the reasoning provided by the annotators to justify why our personalization methods more effectively capture the author’s story-writing style compared to the Average Author method.
Writing Prompt
"It’s terminal," the doctor told you as you, the one renowned as the world’s greatest hero, cradled your dying son, "sometimes, when the super genes mix, the outcome isn’t as expected."
Ground-truth Author Story
# Syntax Error: Unexpected Power Captain Strongarm, defender of Metro City and seven-time winner of Most Cars Thrown At Villains“ award, had never felt so powerless. The hospital machines around his thirteen-year-old son Tommy were having seizures - screens flickering, numbers scrambling, error messages cascading. „It’s terminal,“ Dr. Roberts said, looking at a screen that was now displaying what appeared to be the entire script of ‚The Matrix‘ in real-time. Sarah, his wife (superhero name: The Nullifier), grabbed his hand hard enough to crack concrete. ……„Hey dad, want to learn Python? I promise it’s not actual snakes this time!“ Captain Strongarm smiled. His son wasn’t following in his footsteps - he was coding his own path. And maybe that was even better. „Sure, son. Just… no more downloading martial arts directly into my brain. The kung fu gave me a migraine….Life wasn’t what they expected, but then again, the best source code usually comes with a few surprise features. End_process story.exe Exit status: Happy
Average Author
"It’s terminal," the doctor told you as you, the one renowned as the world’s greatest hero, cradled your dying son, "sometimes, when the super genes mix, the outcome isn’t as expected." The room felt like it was closing in on me, the sterile smell of antiseptic failing to mask the sour stench of fear. My son, little Ethan, lay pale and fragile in my arms, his once vibrant eyes now dull and glassy. This was not how it was supposed to go. Heroes don’t have children who die. Heroes don’t face problems they can’t punch, fly away from, or blast with heat vision. “Isn’t there something we can do?” My voice cracked, and I hated how weak it sounded…….Gently, I laid him down, wrapping him in his favorite blanket. The sky was painted in hues of orange and pink, a beautiful contrast to the sorrow in my heart. I sat there, watching the sun dip below the horizon, and for the first time, I felt truly mortal. I realized that even heroes are not immune to the pain of loss, the fragility of life…….And as the night embraced the world, I knew that Ethan’s spirit would live on, a beacon of hope and courage in the heart of the world’s greatest hero.
Writing Sheet
The Hero’s Journey: A Tale of Capes and Cereal In the heart of the mythical city of Avalon Heights, where ancient heroes once strode and the skies shimmered with the glow of enchanted skyscrapers, the world’s greatest hero faced their most daunting foe: the frailty of life itself. The hero, known to the world as Titania, Queen of the Elements, held her son, Max, in her arms as the doctor delivered the news…… The following day, a fictional news article took center stage in The Avalon Gazette: "World’s Greatest Hero Faces Greatest Challenge Yet: Parenting." In the midst of this storm, Titania’s attempts to find normalcy became legendary. One particularly memorable morning, she decided to make Max a simple breakfast. But as she reached for the cereal, her powers—perhaps expressing her inner turmoil—manifested with unintended flair. A tornado of cereal erupted, swirling through the kitchen like a sugar-coated cyclone. "Mom," Max said, giggling despite his illness, "I think you just invented the breakfast hurricane…… In a final act of empowerment, Titania founded a support group for super-powered families, advocating for genetic research with the same fervor she once reserved for battling cosmic threats. Her growth from feeling helpless to becoming a beacon of hope for others was a transformation as magical as Avalon Heights itself. Max, with dreams and hopes as boundless as the sky, taught her that heroism wasn’t defined by powers but by the courage to embrace life’s uncertainties.
Table 17: Qualitative Example where Writing Sheet method is preferred over Average Author by all three annotators for similarity to the ground-truth author story for a Reddit author.
Writing Prompt
"Late at night, the King of Rohan and his queen sneak into the pantry for a forbidden snack, only to find that stealing cured venison is far more complicated—and far more amusing—than they anticipated."
Ground-truth Author Story
If anyone—a kitchen boy, a maid, or, Eru forbid, even the housekeeper Wynhild—had wandered near the pantry in the quiet hour just before dawn, they would have heard a muffled exchange from the depths of the storeroom: “Try shifting this bowl here…” “Easy, you’re on me.” “Oh, does that bother you, my lady?” “Get off me; I can’t pull it out!……Lothíriel watched her royal husband with a half-incredulous, half-amused smile, as he held the large salami with the pride of a merchant at the autumn fair……To Lothíriel, he looked like a great golden-haired squirrel with the morsel stuffed in one cheek. How he could be both a fierce warrior and a wise sovereign as well, was a mystery she had yet to solve. His voice came out muffled as he swallowed. “You’re right, but if it ever gets out, I’ll say the sweet Queen of the Mark was struck by an irresistible craving for spiced meat in the dead of night—and who would deny anything to the future mother of the next royal heir?” He crossed his arms, looking quite pleased with his idea, but Lothíriel’s eyes widened at the bold lie, and she covered her face with her hand, foreseeing trouble ahead.
Average Author
Éomer, King of Rohan, tiptoed through the dimly lit corridors of Meduseld, holding his breath as he reached the door to the pantry. He glanced back at Lothíriel, his queen, who followed close behind, stifling a giggle. The flickering torchlight cast playful shadows on the stone walls, and Éomer couldn’t help but feel a rush of exhilaration. “We shouldn’t be doing this,” Lothíriel whispered, though her eyes sparkled with mischief. “Ah, but where’s the fun in being King and Queen if we can’t indulge in a little late-night adventure?….. “You’ve always had a knack for adventure. It’s one of the reasons I fell in love with you.” Lothíriel’s expression softened, and she leaned in to kiss him, the taste of venison still on her lips. “And I love you for indulging my whims, even when they lead us to pilfering from our own pantry.”….. Éomer took Lothíriel’s hand, and they slipped out of the pantry, leaving behind only the echo of their laughter and the memory of a stolen moment in the heart of Rohan. “Next time,” Éomer whispered as they crept back to their chambers, “we’ll try the cheese.”
Writing Sheet
The moon cast a silvery glow over Meduseld, its light filtering through the high windows to illuminate the grand halls of the Golden Hall. Éomer, the formidable King of Rohan, found himself tiptoeing like a child on a dare, his hand clasped firmly in Lothíriel’s, whose laughter was a soft melody in the silence….. “Are you sure this is wise, my love?” “Wise? No. Fun? Absolutely,” she replied, her laughter a soft ripple in the silence. Éomer shook his head, a reluctant smile tugging at his lips as he watched her nimble fingers work to untie the savory prize. He was a warrior, a king, yet here he was, caught in the delightful absurdity of this secretive adventure, his heart light with the thrill of it all. “Besides,” Lothíriel continued, handing him a piece of the cured venison with a flourish, “what’s the worst that could happen? A tale of the gallant King brought low by his own appetite?” Éomer took the offered morsel, his grin widening as he tasted the rich, smoky flavor. “Perhaps the bards will sing of our escapades instead of my battles,” he mused, the thought surprisingly pleasing. …..As they slipped back into the corridors, leaving the pantry and its shadows behind, Éomer realized that it was not the venison he would remember, but the adventure, the laughter, and the love that had made the night truly unforgettable.
Table 18: Qualitative Example where the Average Author method is preferred over Writing Sheet by all three annotators for similarity to the ground-truth author story for an AO3 author.
Writing Prompt
"In a desert town under the watchful eyes of royal guards and oppressive banners, you and your team have five minutes to locate a mysterious contact in a pink house before chaos erupts—failure is not an option."
Ground-truth Author Story
# Chapter 1: Danger Close *La Makan, 12:54 standard time* Wind whistled through the close streets of *La Makan*, stirring grains of sand with all the effect of a man attempting to blow salt through eternity. It whipped around the corners of the alleys, rustling banners hanging from balconies and glassless windows. It looked for something, a thing it would never find….. Either way wouldn’t end well for them. Our contact gave us an area, but not an address. We have, he checked his watch, five minutes until the distraction starts. We need to find them and get them out. Problem is, we don’t know what they look like, only that their house is pink. He glanced over his shoulder at the multicolored buildings. Any one of the scoured pink buildings could’ve been their one. “Vic, you knew Xitor the best of any of us. I’m trusting you to get us there. Everyone else, let’s keep our hats on, kay?"
Average Author
The sun hung low in the sky, casting long shadows across the dusty streets of the desert town of Vespera. Time was of the essence; you had five minutes to locate your mysterious contact before chaos erupted. Failure was not an option. “Stay sharp,” you whispered to your team, your voice barely audible over the din of the market. "But time is short. We must act quickly." The weight of the moment settled over you like a heavy mantle. The fate of Vespera rested on the success of this meeting. As the figure began to speak, outlining the plan that would either save or doom the town, you couldn’t shake the feeling that this was only the beginning. The real challenge was yet to come.
Writing Sheet
The sun beat down mercilessly, casting a harsh glare on the town’s whitewashed walls. The narrow alleys were alive with whispers and hurried footsteps, each person a part of the unfolding drama. My team and I moved through the bustling streets, the oppressive presence of royal guards a constant reminder of the stakes. We had five minutes—no more, no less—to locate the contact in the elusive pink house before chaos erupted. The memory of a previous mission flickered in my mind: the sun setting behind the dunes, the sound of distant drums, and the rush of adrenaline as we narrowly escaped a similar fate. That mission had gone wrong, leaving scars that still haunted my dreams. Failure was not an option today… Outside, the tension remained, a silent storm waiting to break. But for now, we had succeeded. The town’s frenetic energy continued, unaware of our brief triumph. The banners flapped, the guards watched, and the countdown to chaos continued, but we had found our mark. The mission was still on course, and the desert town held its breath, waiting for the next move.
Table 19: Qualitative Example where all three annotators select ‘Tie’ for similarity to the author story of a Storium author.
Writing Prompt
"When the word ’hospice’ is first spoken, it feels like the air itself changes, heavy with finality. How do you create light in a space where hope seems to have vanished?"
Ground-truth Author Story
"Hospice." Once the word is uttered aloud, there is a seismic shift. You will feel it. Like a (very short) thread through the eye of a needle, swiftly in and swiftly out. The air itself becomes thin, steely. At the periphery of your vision, an immediate dimming. The penumbra begins to shrink. In time, it will become a tunnel. Ever diminishing. Until the remaining light is small enough to be cupped in two hands. And then it will be extinguished. For when "hospice" is spoken, the fact is at last acknowledged: There is no hope. No hope. These words are obscene, unspeakable…. Well, I will try! When he is at home, possibly his appetite will return. When I am the one to prepare his food, his appetite will return, I am sure. Of course, family will come to visit. Adult children, grandchildren. Relatives, friends. Colleagues from the university. Neighbors. Old friends from grammar school he hasn’t seen in fifty years. Some surprises for him—I will negotiate with the imagination of a theatre director. Not merely hospice but our hospice. Not sad but joyous, a honeymoon. We will be happy there, in our own home. Both of us. For both of us, the "final days" will be a honeymoon. I vow. In fact, nothing remotely like this will happen. How could you have imagined it would! Hospice, yes. Honeymoon, no.
Average Author
"When the word ’hospice’ is first spoken, it feels like the air itself changes, heavy with finality…It’s her mother, Marjorie, who the doctor is addressing, but the word reverberates through Clara’s own heart, a dull thud of inevitability." Marjorie stares at the doctor with a mix of resignation and defiance, her once-vibrant green eyes now clouded by the long battle with illness….. Her touch is light, almost ethereal, and Clara feels a sudden rush of tears she struggles to hold back. “Clara,” Marjorie says softly, her voice barely a whisper. “Do you remember the garden?” Clara nods, unable to speak. The garden had been Marjorie’s sanctuary, a riot of color and life where she had spent countless hours tending to roses, tulips, and wildflowers…… Marjorie’s decline is gradual but inevitable, and Clara feels the weight of each passing day. She continues to read to her mother, to sit by her side, and to hold her hand through the long nights. One crisp morning, as the first frost dusts the garden in a delicate shimmer, Marjorie slips away. Clara is by her side, her hand still entwined with her mother’s, and she feels a profound stillness settle over her…..
Writing Sheet
Every morning, Nurse Clara began her rounds at St. Mary’s Hospice, a place where the word "hospice" hung in the air like a thick, unspoken truth. Her day started with the same cycle: a clipboard in hand, she moved from room to room, her steps echoing softly in the quiet corridors. Each door held a story, a life in its final chapters, and as she entered each room, she would ask the same question, "How are you feeling today?" The responses were as varied as the lives they represented. Some patients greeted her with a weary smile, others with a silent nod. But there was one patient, Mr. Thompson, who never failed to reply with a simple, "I’m just waiting." Clara had come to understand that "waiting" was an ambiguous term within these walls. It held a duality of meaning: waiting for relief, waiting for peace, waiting for the inevitable…..Clara knew that the work was not about erasing the pain or the grief, but about finding moments of grace within it. It was about guiding others, and herself, to a place where hope could be redefined—not as a denial of reality, but as an acceptance of it, and the peace that such acceptance could bring. With each cycle of her rounds, Clara realized that the light had always been there, waiting to be uncovered, and that she had the power to help others see it too.
Table 20: Qualitative Example where all annotators select different winners for similarity to the author story of a New Yorker author.
G.3 Qualitative Examples
We present qualitative examples from our human annotation process, covering various cases: instances where all annotators prefer Writing Sheet, all prefer Average Author, or select ‘Tie.’ Additionally, we include an example where annotators disagree, choosing different winners, highlighting the challenging and subjective nature of the annotation task.
All annotators prefer Writing Sheet:
Table 17 shows an example where all three annotators prefer the Writing Sheet method over the Average Author for similarity to the author story of a Reddit author. The annotators highlight several key aspects that make the Writing Sheet story more aligned with the ground-truth author story: (1) The Writing Sheet method incorporates humor and satire, using phrases such as “a tale of capes and cereal,” “invented breakfast hurricane,” and “World’s Greatest Hero Faces Greatest Challenge Yet: Parenting,” which mirror the lighthearted and satirical tone of the author story. (2) Unlike the Average Author story, which concludes with the son’s death, the Writing Sheet story reinterprets the prompt, portraying Titania as an advocate for families with special children. This aligns with the ground-truth author story’s approach, where the son engages with coding and the narrative avoids a sorrowful ending. (3) In terms of language style, the Writing Sheet story employs imaginative elements such as “The Hero’s Journey: A Tale of Capes and Cereal” and “The Avalon Gazette,” which parallel the ground-truth author story’s unconventional use of computer programming references, such as “Syntax Error: Unexpected Power.”
All annotators prefer Average Author:
Table 18 shows an example where all three annotators prefer the Average Author method over the Writing Sheet for similarity to the author story of an AO3 author. The primary reason cited by annotators is that the Writing Sheet method introduces a strong internal conflict for the king, portraying him as guilty or hesitant about sneaking into the kitchen for a late-night snack. His character is depicted as struggling with the decision, given his stature as a formidable ruler expected to lead battles. Additionally, the Writing Sheet creates a sharp contrast between the king and queen, with the queen depicted as joyful and mischievous, encouraging the king to abandon his guilt.
In contrast, both the ground-truth Author Story and the Average Author story emphasize the playful banter and affectionate dynamic between the king and queen, focusing on the shared experience of sneaking into the kitchen rather than introducing introspection or internal conflict. This shift in focus stems from the Writing Sheet being influenced by the author’s past writing history, where internal conflicts are a recurring theme. However, in this case, character development—being more writing-prompt-specific than broader categories like creativity or language use—did not transfer well. The author’s original story prioritizes external actions and lighthearted interactions rather than deep introspection, leading to a mismatch between the Writing Sheet’s personalized adaptation and the intended tone of the new prompt.
All Annotators Choose Tie:
Table 19 shows an example where all three annotators select ‘Tie’ for similarity to the author story of a Storium author. The primary reason is that both the Average Author story and the Writing Sheet story deviate significantly from the ground-truth author story in terms of plot. The ground-truth story adopts an open-ended approach to both character development and plot, where character relationships remain undefined, and the narrative focuses on anticipation rather than action.
In contrast, the Average Author story follows a conventional narrative arc, featuring fast-paced dialogue and character interactions that build tension and lead to resolution. Similarly, the Writing Sheet story also resolves the narrative but differs from the Average Author method by emphasizing past missions and their consequences, aligning with the author’s tendency to reference character backstories. However, both methods fail to capture the open-ended nature of the original story. This misalignment in plot structure leads all three annotators to select ‘Tie’ for this example.
All Annotators Disagree on the Winner:
Table 20 shows an example where all annotators select different winners for similarity to the author story of a New Yorker author. The ground-truth author story depicts the protagonist’s husband in his final days, exploring familial bonds and the emotional depth of their relationship while infusing a deeper metaphorical meaning into the term “hospice.” The story is rich in subtext and exhibits several distinctive traits of the author, such as cyclic plot structures and the deliberate repetition of phrases like “Adult children, grandchildren. Relatives, friends. Colleagues…” and “Not merely hospice but our hospice. Not sad but joyous, a honeymoon.”
The annotator who selected Writing Sheet as the winner states, ‘‘Although Average Author and the Author Story depict relationships between close family members, Writing Sheet is more similar to the Author Story in Plot, Creativity, and Development by exploring the meaning and emotional responses to the term hospice.’’ This annotator highlights the presence of subtext in the Writing Sheet story and its attempts to mimic the author’s cyclic style—“Each door held a story, a life in its final chapter… she would ask the same question…” and “Clara knew that work was not about erasing the pain or the grief, but about finding moments of grace within it…”.
The annotator who selected Average Author as the winner states, ‘‘Average Author aligns more closely with the Author’s Story in terms of Plot and Language Use. This is because both stories center on someone taking care of a family member in their own home, and growing from a lack of peace into a sense of peace, which is expressed through the choice of language. Writing Sheet also indicates a development from lack of peace to acceptance, but the familial tie is what pushes Average Author over the top.’’
The annotator who selected ‘Tie’ states, ‘‘I think it will be hard for one of these to resemble the Author Story since it was so unique. In all, the two AI stories were much more like each other than the Author Story.’’
Overall, this example highlights the subjective nature of annotator preferences and the inherent difficulty of our annotation task which also explains the poor inter-annotator agreement.

### v2

@@
 Abstract
-Personalization has become essential for improving user experience in interactive writing and educational applications, yet its potential in story generation remains largely unexplored.
-In this work, we propose a novel two-stage pipeline for personalized story generation.
-Our approach first infers an author’s implicit story-writing characteristics from their past work and organizes them into an Author Writing Sheet that is inspired by narrative theory.
-The second stage uses this sheet to simulate the author’s persona through tailored persona descriptions and personalized story writing rules.
-To enable and validate our approach, we construct Mythos, a dataset of 590 stories from 64 authors across 5 distinct sources that reflect diverse story-writing settings.
-A head-to-head comparison with a non-personalized baseline demonstrates our pipeline’s effectiveness in generating high-quality personalized stories.
-Our personalized stories achieve a 75% win rate (versus 14% for the baseline and 11% ties) in capturing authors’ writing style based on their past works.
-Human evaluation highlights the high quality of our Author Writing Sheet and provides valuable insights into the personalized story generation task.
-Notable takeaways are that writings from certain sources, such as Reddit, are easier to personalize than others, like AO3, while narrative aspects, like Creativity and Language Use, are easier to personalize than others, like Plot.
+Personalization is critical for improving user experience in interactive writing and educational applications, yet remains understudied in story generation.
+We study the task of personalizing story generation, where our goal is to mimic an author’s writing style, given other stories written by them.
+We collect Mythos, a dataset of 3.6k stories from 112 authors, with an average of 16 stories per author, across five distinct sources reflecting diverse story-writing settings.
+We propose a two-stage pipeline for personalized story generation: first, we infer authors’ implicit writing characteristics and organize them into an Author Writing Sheet, which is validated by humans to be of high quality; second, we simulate the author’s persona using tailored persona descriptions and personalized story rules.
+We find that stories personalized using the Author Writing Sheet outperform a non-personalized baseline, achieving a 78% win-rate in capturing authors’ past style and 59% in similarity to ground-truth author stories.
+Human evaluation supports these findings and further highlights trends, such as Reddit stories being easier to personalize, and the Creativity and Language Use aspects of stories being easier to personalize than the Plot.
 [Uncaptioned image]     github.com/Nish-19/Persona-Story-Gen
@@
 Stage 2 uses the Author Writing Sheet to role-play the author, incorporating tailored story rules and a persona description for personalized generation.
-Personalizing the outputs of large language models (LLMs) has promising applications for role-playing in chatbots and video games (Tseng et al., 2024; Chen et al., 2024), as well as providing personalized feedback in education (Stahl et al., 2024; Li et al., 2024b; Makridis et al., 2024).
-However, story-writing personalization, a crucial writing system feature that can assist users in overcoming writer’s block or help second-language learners adapt to a new language (Yuan et al., 2022; Yeh et al., 2024), remains largely unexplored.
-To address this gap, we systematically explore how to personalize LLM outputs based on individual authors’ previous writings.
-We propose a novel two-stage pipeline for personalized story generation, as shown in Figure 1.
-In the first stage, we construct an Author Writing Sheet to capture implicit writing characteristics from an author’s past writings (Sun et al., 2025, 2024; Ramos et al., 2024).
-We categorize story-writing characteristics into four fine-grained narrative dimensions: Plot, Creativity, Development, and Language Use (Pavis, 1998; Card, 1999; Noble, 1994; Huot et al., 2024).
-Each characteristic is represented as a Claim-Evidence pair, where general statements about the author’s writing behavior are supported by specific story excerpts.
-This approach, inspired by Common Core Standards in English Language Arts (Association et al., 2010) and recent work in writing education (Li et al., 2023), provides a structured and grounded way to identify an author’s distinctive writing patterns, as validated via our human evaluation (Section 3).
-In the second stage, we implement Personalized Story Generation by having LLMs role-play authors based on their Writing Sheet.
-Our method uses persona prompts and story rules as user constraints to guide the generation.
-To enable and validate our approach, we construct Mythos, a dataset of 590 stories from 64 authors ($\approx$1500 tokens) across five sources with distinct writing settings: Reddit, AO3, Storium, Narrative Magazine, and New Yorker.
-The dataset splits each author’s works chronologically into two sets: an earlier Profiling set used to infer their writing characteristics, and a later Generation set used to evaluate how well LLMs can emulate their distinctive writing behavior (Salemi et al., 2024).
-We evaluate personalized stories in two ways: First, we use automatic evaluation with LLM-as-a-judge, to assess faithfulness to the author’s writing history and similarity to ground truth author writings, along with traditional metrics for lexical overlap, story diversity, and style similarity.
-Second, we conduct a human evaluation to assess the generated story’s similarity to the author’s ground-truth writings.
-We find that stories generated using Author Writing Sheets demonstrate clear advantages over a non-personalized baseline approach.
-When evaluated by LLMs, personalized stories have 61% higher faithfulness to the author writing history.
-Human annotators prefer these stories by an 18% absolute margin in terms of similarity to ground truth.
-Our personalization approach is especially effective for Reddit content and for capturing Creativity, and Language Use characteristics, with less pronounced impact on Plot (Tian et al., 2024; Xu et al., 2024).
+Large Language Models (LLMs) are increasingly used in interactive systems to support human users in co-writing tasks (Fang et al., 2024; Yuan et al., 2022; Chung et al., 2022).
+A key challenge in these settings to generate personalized outputs that align with individual user preferences (Yeh et al., 2024), which improves engagement, agency, and writing quality (Zhao et al., 2025).
+Personalization is also key in educational applications to deliver effective feedback to students (Anderson et al., 1985), and may support users facing writer’s block (Yuan et al., 2022) or help second language learners in learning a new language (Baffour et al., 2023; Su et al., 2023).
+Despite these potentials, personalization in story writing has received limited attention.
+Story-writing personalization faces several challenges, including a lack of existing resources for evaluating personalization, the inherent subjectivity in user preferences, and methods to evaluate the extent of personalization.
+To address these gaps, we construct Mythos, a dataset of 3.6k human-written stories ($\approx$1500 tokens each) from 112 authors (16 stories each) across five distinct writing settings: Reddit, AO3, Storium, Narrative Magazine, and New Yorker.
+We split each author’s stories chronologically into two sets: an earlier Profiling set used to infer their writing characteristics, and a later Generation set used to evaluate how well LLMs can emulate their distinctive writing behavior (Salemi et al., 2024).
+We systematically explore personalized story generation using LLMs with a two-stage pipeline as shown in Figure 1: First, we construct an Author Writing Sheet to extract implicit writing characteristics from an author’s past stories (Ramos et al., 2024).
+The Author Writing Sheet covers four narrative dimensions: Plot, Creativity, Development, and Language Use (Pavis, 1998; Huot et al., 2024), and is organized as Claim-Evidence pairs grounded in story excerpts to summarize author preferences.
+This representation, inspired by Common Core Standards (Association et al., 2010) and writing education research (Li et al., 2023), offers a cost-effective and interpretable way to model authors’ writing preferences.
+Second, we guide LLMs to role-play authors using persona prompts and story rules derived from the Author Writing Sheet.
+We evaluate generated personalized stories in two ways: First, we use automatic evaluation with LLM-as-a-judge to assess faithfulness to the author’s writing history and similarity to ground truth author story, along with traditional text similarity metrics on lexical overlap, story diversity, and style similarity.
+Second, we conduct a human evaluation to assess the generated story’s similarity to the author’s actual writings.
+We find that stories generated using the Author Writing Sheet outperform a non-personalized baseline: LLMs prefer them 78% of the time in faithfulness and 59% in similarity to ground truth, while human annotators prefer them 56% of the times for similarity to ground truth.
+Our method is especially effective for Reddit stories and in capturing Creativity and Language Use, with a smaller impact on Plot.
 In summary, our contributions are:
-  • Mythos: A novel dataset containing multiple stories written by the same authors, designed to support both author profiling and personalized story generation evaluation.
-  • Author Writing Sheet: A structured approach to capturing an author’s writing behavior based on past works, grounded in narrative theory and educational practices and validated through human evaluation.
-  • Personalized Story Generation: A method for generating stories in an author’s style using Author Writing Sheets, rigorously evaluated through automated and human assessments, demonstrating its superiority over baselines.
+  • Mythos: A dataset of 3.6k stories with multiple stories per author, supporting both author profiling and personalized story generation evaluation.
+  • Author Writing Sheet: A cost-efficient, structured method for capturing author characteristics, grounded in narrative theory and validated through human evaluation.
+  • Personalization Evaluation: An evaluation combining automated and human assessments, showing the benefits of Author Writing Sheets.
 2 Dataset Collection and Description
-To facilitate our analysis of personalized story generation, we ensemble Mythos, a novel dataset consisting of 590 stories written by 64 authors.
-To the best of our knowledge, Mythos is the first dataset that identifies and connects multiple stories written by the same author.
-See Table 1 for comparison with existing datasets.
+To facilitate our analysis of personalized story generation, we assemble Mythos, a dataset consisting of 3.6k stories written by 112 authors.
+To the best of our knowledge, Mythos is the first dataset that identifies and connects multiple stories written by the same author (see Table 1 for comparison).
 Data Sources:
-We include five diverse story-writing sources, namely (1) Reddit, featuring stories from r/WritingPrompts, a widely-used resource for amateur story-writing research (Fan et al., 2018), (2) AO3, featuring fanfiction based on popular franchises such as Harry Potter and Sherlock Holmes, (3) Storium, featuring collaborative stories (Akoury et al., 2020), (4) N.Magazine, featuring professionally polished stories by authors such as Barry Gifford, and finally (5) New Yorker, featuring expert-level storytelling from renowned authors such as Haruki Murakami.
+We include five diverse story-writing sources: (1) Reddit, featuring stories from r/WritingPrompts, a widely-used resource for amateur story-writing research (Fan et al., 2018), forming the largest subset with 3.2k stories, (2) AO3, featuring fanfiction based on popular franchises like Harry Potter, (3) Storium, featuring collaborative stories (Akoury et al., 2020), (4) N.Mag (Narrative Magazine), featuring professional stories from authors like Barry Gifford, and (5) N.York (New Yorker), featuring expert-level storytelling from authors like Haruki Murakami.
 Preprocessing:
-To ensure high-quality and diverse content, we impose constraints on word length (ranging from 500 to 1500 words), limit publication dates to post-November 2023, and exclude explicit material.
-To standardize formatting, non-Reddit stories are supplemented with writing prompts generated by GPT-4o and manually reviewed for accuracy.
-Stories are split chronologically based on their submission timestamps, with 70% used for profiling writing characteristics and 30% for generation (see Appendix A for details).
-Dataset size:
-Our dataset is limited in size due to data crawling policies that restrict large-scale data collection for analysis [^1] https://archiveofourown.org/admin_posts/25888.
-We note that recent storywriting datasets such as TELL ME A STORY (Huot et al., 2024) and TTCW (Chakrabarty et al., 2024) have conducted analyses on a comparable number of stories (Table 1).
-Table 1: Comparison of our dataset with existing story-writing datasets.
-✗ indicates no Author IDs associated with stories, and $\sim$ denotes having Author IDs but not having explicit links between them.
-Our dataset spans diverse settings and links stories by the same author, enabling research on personalized story generation.
+To ensure high-quality and diverse content, we impose constraints on story length (ranging from 500 to 1500 words), limit publication dates to post-November 2022, and exclude stories with explicit content.
+To standardize formatting, non-Reddit stories are supplemented with writing prompts generated by GPT-4o, which are manually reviewed for accuracy.
+Stories are split chronologically, with 70% used for profiling writing characteristics and 30% for testing generation.
+Details on our dataset can be found in Appendix A.
 Dataset Name Size Sources Prompt Length Story Length Author IDs
@@
 Storium (Akoury et al., 2020) 440k Storium online platform 247 247 $\sim$
-TTCW (Chakrabarty et al., 2024) 12 New Yorker 54 2000 $\sim$
-Mythos 590 Reddit, AO3, Storium, N.Magazine, New Yorker 50 1517
+Mythos (ours) 3.6k Reddit, AO3, Storium, N.Mag, N.York 50 1517
+Table 1: Comparison of our dataset with existing story-writing datasets.
+✗ indicates no Author IDs associated with stories, and $\sim$ denotes having Author IDs but not having explicit links between them.
+Mythos covers diverse story-writing settings and links stories by the same author, enabling future research on personalized story generation.
 3 Author Writing Sheet
-Motivated by the connection between writing education and personalization (Li et al., 2023), we propose a novel method to infer implicit story-writing characteristics of authors based on their history in the profiling set $P$.
-We group these characteristics into four narrative categories: Plot, Creativity, Development, and Language Use, inspired by narrative theory (Pavis, 1998; Card, 1999; Noble, 1994; Huot et al., 2024), to assess fine-grained, multi-dimensional story-writing attributes:
+Motivated by the connection between writing education and personalization (Li et al., 2023), we propose a method to infer authors’ implicit story-writing tendencies based on their past stories in the profiling set, denoted as $P$.
+We group these tendencies into four narrative categories: Plot, Creativity, Development, and Language Use inspired by narrative theory (Pavis, 1998; Huot et al., 2024), to capture fine-grained, multi-dimensional aspects of how authors approach story-writing:
   • Plot: Story structure, conflict introduction, prompt engagement, and resolution.
@@
   • Language Use: Diction, style, rhetorical devices, pacing, and dialogue.
-Inspired by Common Core (CC) Standards in English Language and Arts (Association et al., 2010), each characteristic is represented as a Claim-Evidence pair, where the Claim summarizes the author’s writing style, and Evidence consists of excerpts from the author-written story to justify the Claim.
-These characteristics are organized into an Author Writing Sheet (Gurung and Lapata, 2024).
-Stage 1 in Figure 1 (and Algorithm 1 in the Appendix) illustrates our method for generating the Author Writing Sheet by iteratively processing each story in an author’s profiling set.
-For a given writing prompt ($wp_{t}$) and corresponding author-written story ($s_{at}$), we use an LLM (LLMavg) to generate an average story ($s_{bt}$).
-Since LLMs are pre-trained on large-scale text data from the web, this average story simulates a typical author’s response based on source-specific stylistic patterns (see Average Author).
-We prompt an LLM (LLMsheet) to compare the average and author-written stories and generate Claim-Evidence pairs for each narrative category to form an intermediate Author Writing Sheet $A_{t}^{\prime}$, which captures how the author’s writing characteristics deviate from those of a typical author for that source (Shashidhar et al., 2024; Krishna et al., 2020).
-To construct the final Author Writing Sheet $A_{|P|}$, we iteratively merge each intermediate Author Writing Sheet $A_{t}^{\prime}$ with the aggregated sheet from prior stories $A_{t-1}$, maintaining a moving update similar to (Chang et al., 2023).
-During merging, we prompt an LLM (LLMcombine) to integrate $A_{t}^{\prime}$ into $A_{t-1}$ by grouping equivalent Claims within each narrative category, selecting the best Evidence for grouped Claims, and retaining ungrouped Claims with their respective Evidence.
-To improve clarity, the updated sheet includes a timestamp identifier for each Evidence corresponding to the story it belongs to and also limits each narrative category to 10 Claim-Evidence pairs, prioritizing grouped Claims.
-For LLMavg, LLMsheet, and LLMcombine we use GPT-4o prompted in a chain-of-thought manner with a temperature of 0.0 and a maximum output of 4096 tokens (Wei et al., 2022; Shashidhar et al., 2024).
-Our method processes each story separately, making the update process efficient when new submissions are added (Yeh et al., 2024; Yuan et al., 2022).
-This method avoids the need to reprocess all previous stories, reducing computational cost and mitigating the limitations of long-context inputs in LLMs (Zhou et al., 2023; Magar and Schwartz, 2022; Li et al., 2024a).
-This approach draws parallels to Knowledge Tracing in education research, where a model continuously updates its understanding of a student’s progress without re-evaluating past interactions (Liu et al., 2022; Scarlatos and Lan, 2024).
-See Appendix B.1 for cost analysis.
-Validating Author Writing Sheets:
-We conduct a human evaluation with three Upwork annotators (\$17/hour).
-Annotators assess the Claim-Evidence pairs based on two criteria: (1) whether the Claims can be reasonably inferred from the story (Yes/No) and (2) whether the Evidence supports the Claims (Yes/Partially/No).
-The study includes 42 authors spanning all five sources, each represented by one sampled story resulting in 188 annotated claims.
-Of these, 12 stories are annotated by all annotators to measure inter-rater agreement, while each annotator evaluates 10 additional exclusive stories.
-Results show unanimous agreement on the ‘Yes’ label for Claim inference and moderate agreement on Evidence support (Krippendorff’s Alpha=.$=.57$), with 93% of Evidence fully supporting Claims and 7% providing partial support.
-These findings confirm the reliability and quality of the Author Writing Sheets and support their use in personalized story generation (see Appendix B.2 for more details).
+Inspired by the Common Core Standards in English Language Arts (Association et al., 2010), each characteristic is represented as a Claim-Evidence pair, where the Claim summarizes the author’s style and the Evidence are supporting story excerpts.
+Stage 1 in Figure 1 (and Algorithm 1 in the Appendix) illustrates our method for generating the Author Writing Sheet (example in Table 7), where we iteratively process each story in an author’s profiling set.
+We construct the sheet by merging characteristics across stories that highlight how the author’s style differs from that of a typical, average author.
+For each writing prompt ($wp_{t}$) and its corresponding author story ($s_{at}$), we use the LLM with two distinct prompts: one prompt (LLMavg) elicits an average story ($s_{bt}$) that reflects the typical style of an author on that writing source, relying solely on source-specific priors acquired during pre-training; another prompt (LLMsheet) asks the same LLM to compare the author story ($s_{at}$) with the average story ($s_{bt}$) and extract corresponding Claim-Evidence pairs across narrative categories.
+This yields an intermediate Author Writing Sheet $A_{t}^{\prime}$ (Shashidhar et al., 2024; Krishna et al., 2020).
+To construct the final Author Writing Sheet $A_{|P|}$, we iteratively merge each intermediate sheet $A_{t}^{\prime}$ with the aggregated sheet $A_{t-1}$, following a moving update approach (Chang et al., 2023).
+To reduce redundancy and manage output length, we prompt the LLM with a prompt (LLMcombine) to group equivalent Claims, select the best Evidence, and retain ungrouped Claims.
+We further limit each narrative category to 10 Claim-Evidence pairs prioritizing grouped Claims to fit within the LLM’s context window in subsequent iterations.
+Each Evidence is tagged with a timestamp indicating its source story.
+Similar to Knowledge Tracing in education (Liu et al., 2022; Scarlatos and Lan, 2024), our method processes each story sequentially.
+Doing so enables efficient updates to the Author Writing Sheet without having to re-process prior stories every time a new story is analyzed (Yeh et al., 2024).
+Therefore, this method reduces cost and mitigates long-context limitations in LLMs (Li et al., 2024); see Appendix B.1 for a detailed cost analysis.
+Validating Author Writing Sheet Quality:
+We conduct a human evaluation with three Upwork annotators to assess the quality of the Claim-Evidence pairs.
+Annotators evaluate each pair on two criteria: (1) whether the Claim can be reasonably inferred from the story (Yes/No), and (2) whether the Evidence supports the Claim (Yes/Partially/No).
+The study spans 42 authors across five sources, with one sampled story per author, yielding 188 annotated claims.
+To measure inter-rater reliability, 12 stories are annotated by all annotators.
+Results show unanimous agreement on the ‘Yes’ label of Claims and moderate agreement on Evidence support (Krippendorff’s $\alpha=0.57$), despite the task’s subjectivity.
+Overall, 93% of Evidence statements fully support their Claims, and the remaining 7% provide partial support.
+These results indicate high precision in constructing Claim-Evidence pairs and confirm the reliability and quality of the Author Writing Sheet; see Appendix B.2 for more details including experimental design and interface (Figure 13), pilot studies, and results.
 4 Personalized Story Generation
-Stage 2 of Figure 1 illustrates our proposed method for personalized story generation using the Author Writing Sheet.
-To generate personalized stories, we prompt an LLM (LLMstory) with the writing prompt, story length (Chakrabarty et al., 2024), source-specific metadata (e.g., fanfiction for AO3), story rules as actionable instructions in direct second-person form (categorized into four narrative categories), and persona description obtained from the Author Writing Sheet.
-We experiment with four types of methods for personalized story generation: 1) A non-personalized baseline, Average Author; 2) Personalization baselines, RAG and Delta, that do not use the Author Writing Sheet; 3) Our proposed methods, Writing Sheet and Writing Summary; and 4) An Oracle method that loosely resembles the upper bound performance.
-We detail each method type below.
-See Appendix C for the prompts.
-Non-Personalized - Average Author:
-The Average Author method serves as a non-personalized baseline that reflects the average behavior of authors, acquired during an LLM’s pre-training process.
-Specifically, for each given source, we prompt an LLM to generate stories using an Average Author prompt that describes typical writing characteristics for the given story source obtained using audited GPT-4o prompting [^2] We ask GPT-4o for the typical writing characteristics for the story source and manually verify it.
-(Wang et al., 2024).
-For example, for AO3, the Average Author is defined as “respecting fandom tone and style while experimenting with established tropes, unconventional pairings, or alternative universes.”
-Personalization Baselines
-RAG:
-The Retrieval-Augmented Generation, RAG baseline (Salemi et al., 2024) first retrieves the most similar writing prompt and author-written story from the profiling set using BM25 (Robertson et al., 2009).
-We then use the retrieved pair as a one-shot demonstration (Wang et al., 2024) to elicit role-playing behavior from the LLM, mimicking the retrieved example’s style.
-Delta:
-The Delta method generates personalized story rules by contrasting the Average Author story (see Average Author) with the corresponding author-written story for each writing prompt in the profiling set (Shashidhar et al., 2024).
-We then use all writing prompts in the profiling set, along with their corresponding generated story rules that are actionable instructions in direct second-person form, as few-shot demonstrations for the LLM, to generate personalized story rules for a new prompt in the generation set.
-Proposed Method - Writing Sheet and Summary
-Writing Sheet:
-Our Writing Sheet method uses the Author Writing Sheet (Section 3) to enable personalization (Stage 2 of Figure 1).
+We now detail our proposed method for personalized story generation using the Author Writing Sheet, illustrated in Stage 2 of Figure 1.
+To generate personalized stories, we prompt an LLM (LLMstory) with the writing prompt, story length, source-specific metadata (e.g., fanfiction for AO3), story rules as actionable instructions in direct second-person form categorized into four narrative categories (example in Table 8), and persona descriptions obtained from the Author Writing Sheet (example in Table 9).
+We experiment with four types of methods: (1) A non-personalized baseline, Average Author; (2) Personalization baselines, RAG and Delta, that do not use the Author Writing Sheet; and (3) Our proposed methods, Sheet and its variant Summary.
+(4) An Oracle method that loosely resembles the upper bound performance; See Appendix C for the prompts used.
+Non-Personalized: Average Author
+The Average Author method serves as a non-personalized baseline reflecting typical author behavior learned during LLM pre-training.
+For each source, we prompt an LLM with an Average Author description summarizing writing characteristics obtained through audited GPT-4o prompting, [^1] We ask GPT-4o for typical writing characteristics for each source and manually verify them.
+verified manually (Wang et al., 2024).
+Example, for AO3, the Average Author is defined as “respecting fandom tone and style while experimenting with tropes, unconventional pairings, and alternate universes.”
+Personalization Baselines: RAG and Delta
+RAG
+The Retrieval-Augmented Generation (RAG) baseline (Salemi et al., 2024) first retrieves the most similar writing prompt and author story from the profiling set using BM25 (Robertson et al., 2009).
+Following (Wang et al., 2024), the retrieved pair is used as a one-shot demonstration to elicit role-playing behavior from the LLM, mimicking the retrieved example’s style.
+Delta
+The Delta method generates personalized story rules by contrasting the Average Author story (see Average Author) with the corresponding author-written story for each writing prompt in the profiling set.
+Following (Shashidhar et al., 2024), we then use all writing prompts in the profiling set, along with their corresponding generated story rules that are actionable instructions in direct second-person form, as few-shot demonstrations for the LLM, to generate personalized story rules for a new prompt in the generation set.
+Our Methods: Sheet and Summary
+Sheet
+Our Sheet method uses the Author Writing Sheet (Section 3) for personalization (Stage 2 of Figure 1).
 First, we prompt an LLM (LLMpersona) to generate a persona description that summarizes the author’s story-writing style as a second-person narrative, included in the system prompt (Wang et al., 2024; Jiang et al., 2024).
-Second, we prompt an LLM (LLMrule) to generate personalized story rules from the Author Writing Sheet tailored to the writing prompt in the generation set, included as constraints in the user prompt (Pham et al., 2024).
-Additionally, we include a one-shot demonstration, following the approach described in RAG (Richardson et al., 2023).
-Writing Summary:
-As an alternative to the Author Writing Sheet, the Writing Summary method leverages the LLM’s long-context capabilities (Ding et al., 2023) by packing all past stories from the profiling set as input.
-Specifically, we include them in a single input prompt to generate an Author Writing Summary in the same format as the Author Writing Sheet.
-Similar to the Writing Sheet method, persona descriptions, and story rules are derived from the Author Writing Summary and used as constraints for story generation, along with a one-shot demonstration.
-Oracle:
-The Oracle method establishes an upper bound on personalization performance, in terms of similarity to the ground truth, by using Oracle story rules derived directly from the ground-truth author-written story for each writing prompt in the generation set.
-These rules are obtained by contrasting the Average Author with the ground-truth author story, following the same process as in Delta.
-Additionally, we include a one-shot demonstration, following the approach described in RAG.
+Second, we prompt an LLM (LLMrule) to generate personalized story rules from the Author Writing Sheet tailored to the writing prompt in the generation set, included as constraints in the user prompt (Pham et al., 2024) for LLMstory.
+Additionally, following (Richardson et al., 2023), we include a one-shot demonstration using the approach described in RAG.
+Summary (Summ)
+As a variant of the Author Writing Sheet, the Summ method leverages the LLM’s long-context capabilities (Ding et al., 2023) by packing all past stories from the profiling set as input in the same prompt to generate an Author Writing Summary in the same format as the Author Writing Sheet.
+Similar to the Sheet method, persona descriptions and story rules are derived from the Author Writing Summary and used as constraints for story generation, along with a one-shot demonstration.
+Reference Method: Oracle
+The Oracle method establishes an upper bound on personalization performance by using story rules derived directly from the ground-truth author story for each writing prompt.
+These rules are obtained by contrasting the Average Author with the ground-truth story (see Delta).
+We also include a one-shot demonstration following the approach described in RAG.
 5 Experiments
-In this section, we describe the implementation details and evaluation methods.
+We now detail our experimental/evaluation settings.
 5.1 Implementation Details
-We prompt GPT-4o in a chain-of-thought manner (Wei et al., 2022; Shashidhar et al., 2024) with a temperature of 0.0 and a maximum token limit of 4096 to generate story rules for the Delta and Oracle methods, as well as for generating the Author Writing Sheet, Author Writing Summary, and their respective persona descriptions (LLMpersona) and personalized story rules (LLMrule).
-For story generation (LLMstory) across all methods, including Average Author, we evaluate three models: GPT-4o (Bubeck et al., 2023), Llama 3.1 8B, and Llama 3.1 70B (Dubey et al., 2024), using a temperature of 0.7 and top_p of 0.95 (Wang et al., 2024) (See Appendix F for results using Llama models for story generation).
-Additionally, we implement an ablation variant of our proposed method for Writing Sheet and Writing Summary that excludes the persona description from the system prompt called Writing Sheet nP and Writing Summary nP respectively.
-5.2 Automatic Evaluation
-We evaluate the performance of personalized story generation using two automated evaluation methods: First, we use GPT-4o-as-a-judge which is shown to be a scalable way to approximate human judgments for complex open-ended tasks (Zheng et al., 2023) and calculate the win-rate of our personalization methods against the non-personalized method, Average Author on two aspects:
-  • Faithfulness to Writing History: We evaluate how well the generated story aligns with the author history by measuring win-rates using Claims from the Author Writing Summary as the ground-truth reference for each narrative category (Wang et al., 2024, 2023; Yunusov et al., 2024).
+We prompt GPT-4o in a chain-of-thought manner (Wei et al., 2022) with temperature 0 and a maximum token limit of 4096 for all the methods described above (Section 4) to generate story rules and persona descriptions.
+Only for story generation (LLMstory) across all methods, including Average Author, we evaluate three different models: GPT-4o [^2] For Reddit, we only consider 5 stories per author for generation totaling 300 stories, due to API cost constraints.
+(Bubeck et al., 2023), Llama 3.1 8B, and Llama 3.1 70B (Dubey et al., 2024), using a temperature of 0.7 and top_p of 0.95 (Wang et al., 2024) (See Appendix F for results using Llama models).
+Additionally, we implement an ablation variant of our proposed method for Sheet and Summ that excludes the persona description from the system prompt called Sheet-nP and Summ-nP, respectively.
+5.2 Automated Evaluation
+We evaluate the performance of personalized story generation using two automated evaluation methods: First, we use LLM-as-a-judge, which is a scalable alternative to human evaluation for complex open-ended text generation tasks (Zheng et al., 2023) and calculate the win-rate of a story generation method against the non-personalized baseline, Average Author, on two aspects:
+  • Faithfulness to Writing History: We evaluate how well the generated story aligns with the author history by measuring win-rates using Claims for each narrative category from the Author Writing Summary as the reference (Wang et al., 2023; Yunusov et al., 2024).
   • Similarity to Author Story: We assess how closely the generated story matches the author’s ground-truth story by measuring win-rates using the ground-truth author story as the reference for each narrative category (Lyu et al., 2024; Shashidhar et al., 2024).
-For both aspects, we prompt GPT-4o in a chain-of-thought manner (Wei et al., 2022) to assign scores (1–5) to stories.
-We randomly shuffle the order of compared pairs, one generated by a personalization method and the other by the Average Author method, to avoid biases.
+For both aspects, we prompt OpenAI o4-mini [^3] https://openai.com/index/introducing-o3-and-o4-mini/ in a chain-of-thought manner (Wei et al., 2022) to assign scores (1–5) to both stories.
+We randomly shuffle the order of stories in comparison, to avoid biases.
 Scores are broken down into four narrative categories: Plot, Creativity, Development, and Language Use (Saha et al., 2024).
-The story with the higher score in a category is declared the winner, and the overall winner is based on the highest total score across all categories.
-Additionally, we evaluate personalization using traditional metrics (Xie et al., 2023) including lexical overlap, story diversity, and stylistic similarity (see Appendix D.2 for more details).
-5.3 Human Evaluation
-We conduct a human evaluation on a subset of our data to further assess subtext and identify any interesting insights that the LLM judges might have missed (Chakrabarty et al., 2024; Subbiah et al., 2024b), by comparing stories generated by personalization methods (Delta, Writing Sheet, Writing Summary) against the Average Author method for similarity to the ground-truth author story (Lyu et al., 2024).
-We ask annotators to justify their choices after making the comparison.
-Three annotators, recruited via Upwork (\$17/hour), evaluated 45 author stories in total (15 common to establish agreement and 10 exclusive to each annotator to increase coverage of annotation (Song et al., 2024)).
-Each author story had three pairs, each containing a story from one personalization method, resulting in 135 annotated story pairs.
+For each category, the story with the higher score is declared the winner, and the overall winner is based on the highest total score across all categories; two stories having the same total scores result in a tie.
+See Appendix D.1 for the prompts.
+Additionally, we evaluate personalization using traditional metrics (Xie et al., 2023), including lexical overlap, story diversity, and stylistic similarity.
+See Appendix D.2 for details.
 6 Results and Discussion
-We now discuss our experimental results and findings.
-Table 2: Percentage win-rate for Faithfulness to Writing History evaluated by GPT-4o.
-Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
-Best win-rates for each source is underlined.
-Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
-AO3 52-32 31-54 49-41 74-20 80-10 68-21 72-19
-Reddit 74-21 35-56 58-26 86-7 89-7 74-19 72-19
-Storium 50-40 42-45 45-42 68-15 75-8 70-20 68-20
-N.Magazine 71-14 57-21 50-36 79-7 71-21 71-21 71-0
-New Yorker 53-40 27-67 53-40 67-20 80-20 67-27 80-13
-Overall 60-29 38-49 51-37 75-14 79-13 70-22 73-14
-Table 3: Percentage win-rate for Similarity with Author Story evaluated by GPT-4o and Human (last column).
-Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
-Best win-rates for each source is underlined.
-Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP Human Eval
-AO3 71-18 30-54 45-32 40-41 60-30 46-38 45-40 37-33
-Reddit 91-5 37-46 49-35 54-35 49-42 47-40 42-44 52-37
-Storium 60-30 45-40 35-35 52-32 50-35 57-14 14-29 30-30
-N.Magazine 64-21 43-36 36-43 50-21 57-21 43-43 50-29 48-30
-New Yorker 67-13 27-47 40-40 27-60 27-47 40-60 47-40 70-15
-Overall 71-17 36-45 41-37 45-38 49-35 47-39 40-36 47-29
-6.1 Automatic Evaluation
-Our automatic evaluation results show that, first, Writing Sheet and Summary consistently outperform other methods.
-Second, persona descriptions improve faithfulness to writing history but improve similarity to the ground truth only for select sources like Reddit.
-Third, personalization is more effective for sources like Reddit than AO3 and Storium, and for narrative categories like Creativity, and Language Use, compared to Plot.
-6.1.1 Faithfulness to Writing History
-Table 2 shows results for faithfulness to writing history using GPT-4o as LLMstory for our personalization methods. We observe the following:
-Writing Sheet and Summary achieve best faithfulness to writing history:
-Writing Sheet and Summary achieve the highest win-rates across all sources, outperforming Average Author by 64%.
-The strong performance of the Writing Summary follows from its reliance on the Author Writing Summary as both a reference for evaluation and a generation constraint.
-Writing Sheet, despite not being explicitly conditioned on the Author Writing Summary, generalizes well, likely due to its format being the same as the Author Writing Summary.
-Oracle outperforms Delta and RAG, benefiting from direct conditioning on ground-truth story rules, which enhances adherence to the author’s writing history.
-However, Oracle underperforms the Writing Sheet and Summary, likely due to its lack of explicit conditioning on the Author Writing Summary.
-Another possible cause is that the author’s writing style possibly changes over time, which is not fully captured by the ground-truth story rules.
-See Appendix E.1 for an example.
-Reddit shows the highest win-rates among sources:
-Reddit exhibits the highest win-rates across all methods, suggesting its stories are more conducive to personalization.
-This trend likely stems from Reddit’s crowd-sourced writing prompts, where authors create prompts for others.
-Therefore, there are a range of diverse topics, ranging from power, survival, and mystery to humor and the supernatural.
-The broad thematic variety may have enabled personalization methods to capture distinct stylistic elements more effectively, resulting in higher win-rates (See the paragraph Themes in Appendix A).
-Persona descriptions improve faithfulness to writing history:
-We see that adding persona descriptions in the system prompt for Writing Sheet and Summary improves win-rates by around 5% over their nP variants.
-This result likely follows from personas enhancing the LLM’s role-playing capabilities, allowing for a more faithful representation of the author’s writing style (Jiang et al., 2024; Wang et al., 2024; Wallace et al., 2024).
-6.1.2 Similarity to Author Story
-Table 3 shows results for similarity to the author story using GPT-4o as LLMstory for our personalization methods. We observe the following:
-Writing Sheet and Summary outperform baselines of RAG and Delta:
-Oracle achieves the highest win-rates across all methods as its story rules come from the ground-truth author story.
-Reddit shows the highest Oracle win-rate (91%), followed by 65% on average for other methods, indicating easier personalization likely due to the broad thematic variety in the writing prompts.
-Writing Sheet and Summary outperform RAG and Delta by 11% and 6%, highlighting the benefit of summarizing the author’s writing history (Richardson et al., 2023).
-Persona descriptions help in sources with broad thematic variety in writing prompts:
-Writing Summary with persona descriptions outperforms its nP variant by 9% across all sources.
-However, Writing Sheet shows no significant difference from its nP variant except for Reddit, where it surpasses it by 7%.
-This result suggests that persona descriptions enhance personalization in sources that are easier to personalize, such as Reddit, but have limited impact elsewhere (Zheng et al., 2024).
-The broad thematic variety in Reddit’s writing prompts creates a greater divergence between author styles and the Average Story, leading to more informative Author Writing Sheets and persona descriptions that further aid personalization.
-See Appendix E.2 for an example.
-Creativity and Language Use are the easiest categories to personalize:
-Figure 2: Win-rate of personalization methods vs. Average Author for Reddit across four narrative categories. D: Delta, S: Writing Summary, WS: Writing Sheet.
-Table 10 (in the Appendix) shows that Delta, Writing Summary, and Writing Sheet outperform Average Author in Creativity (23% higher) and Language Use (15% higher) (Huot et al., 2024), but perform worse in Plot and slightly worse in Development (Tian et al., 2024; Xu et al., 2024).
-Creativity and Language Use are less dependent on the writing prompt, making them easier to transfer across prompts, whereas Plot and Development are more closely tied to specific prompts, making generalization difficult, especially without thematic overlap.
-To better understand personalization performance in Reddit, the easiest source to personalize, Figure 2 shows that Writing Sheet is the best-performing method across all categories, with a 35% advantage over Delta and 20% over Writing Summary in Creativity.
-Writing Sheet explicitly summarizes differences between the ground-truth author story and the Average Story, providing more informative guidance for personalization.
-This advantage is particularly evident in Creativity, which is less constrained by the writing prompt.
-See Appendix E.3 for plots of other data sources.
-6.1.3 Traditional Metrics
-Table 11 (in the Appendix) shows results with the traditional metrics.
-Lexical Overlap and Diversity Metrics yield similar values:
-We observe that Lexical Overlap and Diversity metrics yield similar values, as all methods use the same generation model, leading to overlapping lexical distributions and limiting these metrics’ ability to capture nuanced stylistic differences (Zheng et al., 2023; Xie et al., 2023).
-Writing Sheet outperforms other methods, especially for Style metrics:
-We observe that Writing Sheet nP consistently performs best across all metrics, particularly in Style metrics for both Author History and Author Story, as measured by LUAR (Rivera-Soto et al., 2021).
-This improvement likely results from the Writing Sheet explicitly summarizing an author’s stylistic deviations from an Average Author, enhancing personalization.
-6.2 Human Evaluation
-Table 3 shows win-rates from human evaluation (Section 5.3) under the Human Eval column, showing that personalization methods—Delta, Writing Sheet, and Summary—outperform the Average Author method by 18% points absolute win-rate in terms of similarity to ground-truth writings.
-See Appendix G for details on experiment design, and qualitative examples.
-Low inter-annotator agreement due to task subjectivity:
-Fleiss’ Kappa for inter-annotator agreement ranges from 0.2 to 0.4, indicating low but above-random agreement, likely due to the challenges of annotating long-form subjective texts (Subbiah et al., 2024a).
-Specifically, different annotators prioritize different narrative categories when determining which story in a pair better reflects the author’s style, leading to disagreement.
-Sources like Reddit and New Yorker have better win-rates than others:
-Reddit and New Yorker achieve the highest win-rates, with Reddit benefiting from a broad thematic variety in writing prompts and New Yorker from advanced narrative devices like subtext, which humans assess more reliably than LLMs (Subbiah et al., 2024b).
-In contrast, AO3 and Storium exhibit lower win-rates due to LLMs’ familiarity with fanfiction tropes (AO3) [^3] https://archiveofourown.org/admin_posts/25888 and the lack of strong stylistic differentiation in Storium beyond its open-ended storytelling structure (Xu et al., 2024; Tian et al., 2024).
-These findings suggest that personalization is most effective when authors engage with diverse topics or possess distinct stylistic traits (such as complex narrative structures, creative elements, and advanced language use).
-These characteristics, reflected in their past writing history, provide clear signals for us to mimic their writing style.
-Narrative categories like Creativity, and Language use are more conducive to personalization:
-Annotators favor personalized stories for Creativity, and Language Use, noting their stronger use of deeper symbolism, thematic richness, layered narratives, and expressive language.
+We now present experimental results. Our automatic evaluation shows that the proposed methods are effective, particularly for certain sources and narrative categories.
+6.1 Faithfulness to Writing History
+Table 2 shows results for faithfulness to writing history among GPT-4o-personalized stories. We observe the following:
+Method Reddit AO3 Storium N.Mag N.York All
+RAG 40 52 42 21 40 39
+Delta 56 60 50 64 53 57
+Sheet 74 76 73 85 80 78
+Sheet-nP 74 80 60 85 86 78
+Summ 84 96 87 85 80 87
+Summ-nP 76 86 85 100 100 90
+Table 2: Win-rates (%) for Faithfulness to Writing History vs.
+Average Author baseline for GPT-4o.
+‘Sheet’ (underlined) performs close to the upper bound established by the ‘Summ’ method and significantly outperforms other personalization methods.
+Sheet and Summ achieve best faithfulness to writing history:
+The high performance of Summ sets a performance upper bound, since it relies on the Author Writing Summary, the reference for evaluation, to generate the story.
+Sheet, despite not being explicitly conditioned on the Author Writing Summary, generalizes well (only a 9% gap in win rate compared to Summ), which shows that it is capable in faithfully reflecting the author’s writing history.
+RAG performs the worst since it relies on a single in-context example to mimic the author’s style.
+Delta significantly outperforms RAG (by 18%) but significantly underperforms Sheet (by 21%).
+This result suggests that explicitly summarizing authors’ stylistic characteristics systematically in the form of a Sheet is better than relying on rules extracted from all past author stories.
+See Appendix E.1.1 and Table 10 for an example.
+Persona descriptions benefit amateur sources more than expert sources for Summ:
+We see that personas have little impact on Sheet, since the overall win-rates for Sheet and Sheet-nP are identical.
+This result is likely attributed to using the Author Writing Summary, rather than the Sheet, as the reference for faithfulness evaluation.
+In contrast, for Summ, adding persona descriptions improves win-rates over Summ-nP for amateur sources such as Reddit, AO3, and Storium, but leads to lower performance for expert sources like the N.Mag and N.York.
+This result is likely due to personas for expert authors failing to capture the nuanced and subtle characteristics of their writing styles, which are reflected in their Author Writing Summary.
+See Appendix E.1.2 and Tables 11, 12 for examples.
+6.2 Similarity to Author Story
+Method Reddit AO3 Storium N.Mag N.York All
+Oracle 86 84 78 79 87 83
+RAG 38 56 50 43 27 43
+Delta 41 60 48 43 33 45
+Sheet 62 66 40 71 54 59
+Sheet-nP 47 50 33 64 33 45
+Summ 45 50 38 50 40 45
+Summ-nP 43 53 45 64 20 45
+Table 3: Win-rates (%) for Similarity to Author Story vs. Average Author baseline for GPT-4o. Best methods per source are bolded.
+Table 3 shows results on similarity to the author story among GPT-4o-personalized stories. We observe the following:
+Sheet outperforms all other methods:
+The Sheet method achieves the highest overall win-rate of 59%, outperforming Summ by 14%, Delta by 14%, and RAG by 16%.
+The advantage is more pronounced on Reddit, by far the largest subset of our dataset, where Sheet outperforms Summ by 17%, Delta by 21%, and RAG by 24%.
+These results suggest that relying on a single in-context example (RAG) is insufficient for capturing personalized writing styles, particularly when authors have more varied profiles, as in Reddit.
+In contrast, for sources like Storium, where authors have fewer stories and less consistent styles, RAG is more competitive.
+The Delta method, which uses only story rules, struggles to capture subtle, repeated writing patterns that are better modeled through the longer and more detailed Author Writing Sheet.
+Finally, Summ’s approach of identifying writing characteristics from all profiling stories at once, without contrasting against Average Author stories, is limited by LLMs’ long-context reasoning capabilities (Shashidhar et al., 2024).
+See Appendix E.2.1 and Table 13 for an example.
+Sheet still lags behind the upper-bound Oracle method:
+Although the Sheet method performs best, there remains a substantial overall gap (of 24%) compared to the Oracle method, which highlights the challenging nature of personalization in story generation.
+This gap likely arises because the Sheet cannot capture all of an author’s writing characteristics within the profiling set, due to generation length constraints (4096 tokens).
+Moreover, the Sheet does not explicitly rank characteristics based on their relevance to a new writing prompt, which limits its ability to adaptively emphasize the most salient aspects during generation.
+See Appendix E.2.2 and Table 14 for an example.
+Persona descriptions benefit Sheet:
+We see that Sheet outperforms its ‘nP’ variant overall by 14%, with a particular gain of 15% on Reddit, the largest subset of our data.
+Persona descriptions, provided in the system prompt, offer additional control for personalized story generation by summarizing the author’s general writing style independently of the specific writing prompt, complementing the story rules used in the user prompt.
+See Appendix E.2.3 and Table 15 for an example.
+In contrast, persona descriptions have limited impact on the Summ method, likely because the Author Writing Summary is less effective than the Author Writing Sheet at capturing subtle author characteristics.
+Creativity and Language Use are easier to personalize:
+Method Plot Creativity Development Language Use
+Oracle 74 82 69 88
+Delta 44 55 43 57
+Sheet 49 74 51 62
+Summ 44 51 50 51
+Table 4: Win-rates (%) across narrative categories for Similarity to Author Story, averaged across sources. Best numbers for each method are bolded.
+Figure 2: Win-rates proportions across narrative categories for Similarity to Author Story for Reddit. O: Oracle, D: Delta, Sh: Sheet, Su: Summ.
+Table 4 shows that all methods have higher respective win-rates on Creativity and Language Use than Plot and Development.
+This result likely arises because Creativity (e.g., unique elements introduced by the author) and Language Use (e.g., specific diction or use of dialogue) are less dependent on the writing prompt and thus more generalizable across stories (Huot et al., 2024).
+In contrast, Plot (e.g., conflict introduction or resolution) and Development (e.g., character growth) are closely tied to the writing prompt, making generalization difficult, particularly when there is limited thematic overlap (Tian et al., 2024).
+Figure 2 shows that, for our largest story source, Reddit, the Sheet method achieves the highest win-rates across all categories except Development, where it is slightly outperformed by Summ.
+For Creativity, Sheet performs nearly as well as the Oracle, with only a 5% win-rate gap between them.
+However, substantial gaps remain between Sheet and Oracle for other categories, which highlights the difficulty of the task and room for future improvement.
+See Appendix E.2.4 and Figure 22 for plots from other data sources.
+6.3 Traditional Metrics
+Table 16 in the Appendix shows results with traditional metrics.
+Overall, we see that the lexical overlap and diversity scores are similar across methods, since all generations use the same LLM, resulting in overlapping lexical distributions that limit these metrics’ sensitivity to stylistic differences (Zheng et al., 2023; Xie et al., 2023).
+We also see that the Sheet achieves the best stylistic similarity to both Author History and Author Story, measured by LUAR (Rivera-Soto et al., 2021).
+This result validates the Sheet’s ability to explicitly summarize stylistic deviations from an Average Author, leading to better personalization.
+7 Human Evaluation
+7.1 Setting
+We conduct a human evaluation on a subset of our data to assess narrative subtext and identify insights potentially missed by LLM judges (Chakrabarty et al., 2024; Subbiah et al., 2024b).
+Annotators compare stories generated by personalization methods (Delta, Sheet, and Summ) against the Average Author method on similarity to the ground-truth author story (Lyu et al., 2024).
+For each comparison, annotators choose an overall winner by evaluating both stories on all four narrative categories, and provide justifications as free-form responses.
+Three annotators, recruited via Upwork, evaluated 45 author stories in total.
+There are 15 shared stories across annotators to assess agreement, and 10 were unique to each annotator to expand coverage (Song et al., 2024).
+Each author story appeared in three different pairs (one per personalization method), resulting in 135 annotated story pairs.
+See Appendix G for details on experiment design and the interface (Figure 23).
+7.2 Results
+Method Reddit AO3 Storium N.Mag N.York All
+Delta 44-44 33-33 33-33 33-33 67-11 42-31
+Sheet 67-33 44-22 22-33 67-33 78-11 56-27
+Summ 44-33 33-44 33-22 44-22 67-22 44-29
+Overall 52-37 37-33 30-30 48-30 70-15 47-29
+Table 5: Win-rates (%) for Similarity between GPT-4o-generated and Author stories, evaluated by Humans.
+Each cell (‘X-Y’) shows the method’s (‘X’) and Average Author’s (‘Y’) win-rates, with ties as remainder.
+Best methods per source are bolded.
+Table 5 shows results from our human evaluation.
+Sheet is clearly the best method, outperforming Average Author by 29% across all sources.
+On Reddit, our largest source, it reaches a 67% win-rate, further confirming its advantage.
+Fair inter-rater and model-human agreement:
+Fleiss’ Kappa among human annotators and Cohen’s Kappa between human annotations and the LLM judge (o4-mini) both range from 0.2 to 0.4, indicating fair agreement (Landis and Koch, 1977).
+This result is likely due to the inherent subjectivity of evaluating long-form texts (Subbiah et al., 2024a), where annotators prioritize different narrative aspects when choosing the overall winner.
+The scale of human evaluation, constrained by annotation costs (\$639), also contributes to variability.
+See Appendix G.3 for detailed examples.
+Creativity, and Language use are easier to personalize:
+Same as LLM-as-a-judge, human annotators also favor personalized stories for Creativity and Language Use, noting their stronger use of deeper symbolism, thematic richness, layered narratives, and expressive language.
 However, Average Author was preferred when personalized methods incorrectly introduce elements not present in the author’s ground-truth story, suggesting deviations from the author’s usual style.
+See Table 22 in the Appendix for detailed analysis of annotator free-form comments in each narrative category.
+Reddit and New Yorker are easier to personalize than other sources:
+Reddit and New Yorker show the highest overall win-rates for the personalized methods (52% and 70%, respectively).
+Reddit benefits from the broad thematic variety of its crowd-sourced writing prompts (see Figure 3 in the Appendix), which leads to diverse author writing styles and makes authors easier to distinguish from the Average Author (see Table 23).
+New Yorker, with stories written by experts, benefits from the use of advanced narrative devices such as subtext, which humans can evaluate more reliably than LLMs (Subbiah et al., 2024b) (see Table 24).
+AO3 and Storium show lower overall win-rates for personalized methods.
+For AO3, LLMs’ familiarity with common fanfiction tropes [^4] https://archiveofourown.org/admin_posts/25888 strengthens the Average Author baseline (see Table 25).
+For Storium, the small profile size and weak stylistic differentiation among authors beyond open-ended endings (Xu et al., 2024; Tian et al., 2024) limit the effectiveness of personalization methods (see Table 26).
+These findings suggest that personalization is most effective when authors explore diverse topics or display distinct traits, such as complex narratives, creativity, or advanced language use.
+Such traits, reflected in past writings, provide clearer signals for personalizing their style.
 8 Related Work
 8.1 Personalization and Role-Playing
-Recent works have introduced benchmark datasets for personalizing LLM outputs in tasks like email, abstract, and news writing, focusing on shorter outputs (e.g., 300 tokens for product reviews (Kumar et al., 2024) and 850 for news writing (Shashidhar et al., 2024)).
+Recent works have introduced benchmark datasets for personalizing LLM outputs in tasks like email and news writing, focusing on shorter outputs (e.g., 300 tokens for product reviews (Kumar et al., 2024), 850 for news writing (Shashidhar et al., 2024)).
 These methods infer user traits from history for task-specific personalization (Sun et al., 2024, 2025; Pal et al., 2024; Li et al., 2023; Salemi et al., 2025).
-In contrast, we tackle the more subjective problem of long-form story writing, with author stories averaging 1500 tokens.
-Unlike prior role-playing approaches that use predefined personas (e.g., Tony Stark, Confucius) (Wang et al., 2024; Sadeq et al., 2024; Tu et al., 2023; Xu et al., 2023), we propose a novel method to infer story-writing personas from an author’s history to guide role-playing.
+In contrast, we address the more subjective task of long-form story writing, with author stories averaging 1500 tokens.
+Unlike prior role-playing approaches that rely on predefined personas (e.g., Tony Stark, Confucius) (Wang et al., 2024; Sadeq et al., 2024; Tu et al., 2023; Xu et al., 2023), we propose a novel method to infer story-writing personas from an author’s history to guide role-playing.
 8.2 Story Understanding and Generation
-Prior work on persona-aware story generation (Yunusov et al., 2024; Bae and Kim, 2024; Zhang et al., 2022; Chandu et al., 2019) defines personas using discrete attributes like personality traits, demographics, or hobbies.
-Similarly, (Zhu et al., 2023) explore story style transfer across pre-defined domains (e.g., fairy tales, martial arts, Shakespearean plays).
-In contrast, we mimic an individual author’s writing style based on their history.
-Our approach differs by (1) inferring long-form author personas—descriptions of an author’s style from their past works, rather than relying on demographics, and (2) handling long-form story generation, averaging 1500 tokens per output, exceeding typical story lengths in prior research.
-8 Conclusion
-We introduce the task of personalizing authors’ story-writing styles across five sources with a two-stage approach that infers implicit characteristics as an Author Writing Sheet to guide LLMs via persona descriptions and story rules.
-Evaluations show that summarizing an author’s history via a Writing Sheet or Summary improves faithfulness to writing history and similarity to the ground truth.
-Personalization is more effective for sources with diverse writing prompts and distinctive author traits (e.g., Reddit, New Yorker) and for narrative categories like Creativity and Language Use, which are less writing prompt-dependent.
+Prior work on persona-aware story generation (Yunusov et al., 2024; Zhang et al., 2022; Chandu et al., 2019) defines personas using discrete attributes like personality traits, demographics, or hobbies.
+Similarly, (Zhu et al., 2023) explore style transfer across predefined domains (e.g., fairy tales, martial arts, Shakespearean plays).
+In contrast, we mimic an individual author’s style based on their history.
+Our approach differs by (1) inferring long-form author personas descriptions of style from past works, rather than relying on demographics, and (2) handling long-form story generation, averaging 1500 tokens per output, beyond typical lengths in prior work.
+9 Conclusions and Future Work
+In this paper, we introduced the task of personalizing authors’ story-writing styles across five sources by proposing a new dataset, Mythos, containing 3.6k stories.
+We also proposed a two-stage approach that infers implicit characteristics as a high-quality Author Writing Sheet, to guide LLMs via persona descriptions and story rules.
+Extensive automated and human evaluation showed that our method outperforms baselines for both faithfulness to the author’s history and similarity to the ground-truth author story.
+However, our methods still significantly lagged behind an Oracle method, by as much as 24% in some ways, which highlights the difficulty of the task and the need for future work.
+We also found that personalization is more effective for sources with diverse writing prompts and distinctive author traits (e.g., Reddit, New Yorker), and for narrative categories like Creativity and Language Use, which are less prompt-dependent.
+Future work can address the limitations of the Author Writing Sheet by improving its coverage of relevant stylistic information despite context length constraints, and by retrieving or re-ranking information relevant to the given writing prompt during generation.
+Another promising direction is to explore multi-agent systems for story-writing personalization, with specialized agents handling different narrative categories to improve the extraction and application of author stylistic traits especially for categories like Plot and Development that are hard to personalize.
 10 Limitations
-Limited Author History Size:
-Table 4 (70% of the Stories/Author column) shows that N.Magazine and New Yorker authors have an average of 3 stories in their profiling set, whereas Reddit and AO3 authors have around 10.
-A larger profiling set likely enables more accurate inference of an author’s story-writing characteristics, leading to improved personalization.
-However, the cognitive demands of writing long stories result in fewer stories per author available online, thereby limiting the scalability of personalized story generation research.
+Limited size of Profiling Set for some sources:
+Table 6 shows that when using 70% of each author’s available stories for profiling (from Stories/Author column), authors from Narrative Magazine and New Yorker have only about 2 stories on average, compared to 5 for Storium and around 10 for AO3.
+Reddit is the exception, with authors having an average of 38 stories in their profiling set.
+A larger profiling set likely enables more accurate inference of an author’s story-writing characteristics, leading to better personalization.
+However, our data collection is constrained by practical limitations.
+For all sources except Reddit, we manually collect publicly available stories without web scraping, which restricts the number of samples we can obtain.
+Furthermore, for expert sources like Narrative Magazine and New Yorker, fewer stories per author are available online, limiting the scalability of personalized story generation.
+For Reddit, we adhere to Reddit’s data collection policies and use the official API for dataset construction.
 Author Writing Sheet Recall:
-We validate the precision of our Author Writing Sheet through human evaluation, ensuring that the Claim-Evidence pairs accurately describe the author’s story-writing style.
-However, we do not assess recall, i.e., whether all of the author’s story-writing characteristics are captured.
-Measuring recall is challenging due to the inherent subjectivity of the task, making it impractical to ask humans to independently construct Writing Sheets for comparison.
+Our method for generating the Author Writing Sheet is limited in its ability to capture all of an author’s story-writing characteristics.
+This limitation arises from practical context length constraints and the challenges of long-context reasoning during the iterative merging step, which forces us to restrict the Author Writing Sheet to 10 characteristics per narrative category.
+Regarding human evaluation, we validate only the precision of the Author Writing Sheet—ensuring that the extracted Claim-Evidence pairs accurately reflect the author’s style—but do not assess recall, i.e., whether all relevant characteristics are captured.
+Measuring recall is challenging due to the inherent subjectivity of the task, making it impractical to ask annotators to independently construct Writing Sheets for comparison.
 Multi-Agent Systems and Fine-Tuning LLMs:
 Our work explores role-playing for personalized story generation by prompting an LLM with tailored persona descriptions and story rules.
-Future research can extend this by incorporating multi-agent systems, where specialized agents focus on different narrative categories to enhance personalization (Huot et al., 2024; Bae and Kim, 2024).
-Additionally, fine-tuning LLMs to explicitly reason over author history could be investigated as a method for better aligning generated content with individual author preferences (Salemi et al., 2025; Shashidhar et al., 2024; Shaikh et al., 2024).
-Automatic Evaluation Methods:
-Evaluating long-form text remains challenging for LLMs, particularly for creative outputs that involve nuanced aspects like subtext, which are difficult to assess automatically and often require human evaluation (Subbiah et al., 2024a).
+Future research can extend this by incorporating multi-agent systems, where specialized agents focus on different narrative categories to enhance personalization (Huot et al., 2024; Bae and Kim, 2024), especially for categories like Plot and Development, which are less conducive to personalization.
+Additionally, while our experiments show that fine-tuning improves personalization over prompting for LLaMA 3.1 8B (Section F.2 and Table 21), our work sets a precedent for using fine-tuning to better align LLMs with individual author preferences.
+Future work can build on this by incorporating explicit reasoning over author history to further improve personalization (Salemi et al., 2025; Shashidhar et al., 2024; Shaikh et al., 2024).
+Subjectivity in Evaluation:
+Evaluating long-form creative text remains challenging for LLMs due to the involvement of nuanced aspects like subtext (Subbiah et al., 2024a).
+Despite this subjectivity, we obtain fair Cohen’s Kappa agreement for both human-human and human-model evaluations.
+Future work can explore improving evaluation methods by using aspect-based checklists that capture finer-grained narrative details (Lee et al., 2025).
 11 Ethical Considerations
+Human Evaluation:
 Both human evaluation tasks were approved by an institutional review board (IRB).
@@
 All scientific artifacts, including models and datasets, were used in accordance with their intended purpose to ensure ethical and responsible research practices.
-All data used in this study were manually collected from publicly available sources [^4] Permission was obtained from the authors of the Storium dataset for its inclusion in Mythos., adhering to the data usage and crawling policies of the respective websites.
-To comply with copyright constraints, we will release only the links to the stories used in this research, rather than the stories themselves, following the approach of (Chakrabarty et al., 2024).
-Fully anonymizing our dataset is not feasible [^5] Storium authors already anonymize their dataset.
-as the research involves mimicking specific authors’ writing behaviors.
-We note that even if the usernames are anonymized, it is still possible to infer an author’s publicly available identity from the story link.
-Therefore, for each author for every source, we release their usernames along with corresponding writing prompts, and author-written stories in the profiling and generation set of our dataset.
-However, the dataset remains limited in scale, and no additional metadata, including demographic information, is included, thereby reducing potential risks.
+Dataset Collection:
+All data used in this study were manually collected from publicly available sources, [^5] Permission was obtained from the authors of the Storium dataset for its inclusion in Mythos.
+adhering to the data usage and crawling policies of the respective websites.
+For Reddit, we collected data exclusively using the official Reddit API, following Reddit’s data access policies.
+[^6] https://www.redditinc.com/policies/data-api-terms To comply with copyright constraints, we will release only the URLs linking to the original stories, rather than the story texts themselves, following the approach of (Chakrabarty et al., 2024).
+Data Anonymization and Licensing:
+Fully anonymizing our dataset is not feasible, [^7] Storium authors already anonymize their dataset.
+as the research involves mimicking specific authors’ story writing styles.
+Even if usernames were anonymized, authors’ publicly available identities could still be inferred through the story links.
+Therefore, for each author and source, we release their usernames along with the corresponding writing prompts and author-written stories in the profiling and generation sets.
+No additional metadata, including demographic information, is included, minimizing potential privacy risks.
 To further mitigate unforeseen harms, the dataset will be released under an **Educational or Academic Research, Non-Commercial Use** license, following (Akoury et al., 2020).
 Appendix A Dataset
-In this section, we describe the details of the construction of our dataset which includes five distinct story-writing sources, Reddit [^6] https://www.reddit.com/r/WritingPrompts/, AO3777https://archiveofourown.org/, Storium [^8] https://storium.cs.umass.edu/, N.Magazine [^9] https://www.narrativemagazine.com/, and New Yorker [^10] https://www.newyorker.com/.
+In this section, we describe the details of the construction of our dataset which includes five distinct story-writing sources, Reddit [^8] https://www.reddit.com/r/WritingPrompts/, AO3999https://archiveofourown.org/, Storium [^10] https://storium.cs.umass.edu/, Narrative Magazine [^11] https://www.narrativemagazine.com/, and New Yorker [^12] https://www.newyorker.com/.
 Selection Constraints:
@@
 Stories are limited to a length of 500 to 1500 words.
-To avoid contamination by LLM pretraining data, only recent stories published after November 2023 are included for Reddit and AO3 (Zhou et al., 2023; Magar and Schwartz, 2022).
+To avoid contamination by LLM pretraining data, only recent stories published after the release of ChatGPT [^13] https://help.openai.com/en/articles/6825453-chatgpt-release-notes November 2022 are included for Reddit and AO3 (Zhou et al., 2023; Magar and Schwartz, 2022).
 Not-Safe-For-Work (NSFW) and explicit content is excluded using automatic tagging and manual verification.
@@
 Statistics and Comparisons:
-The dataset contains 590 stories from 64 authors, with an average story length of 1500 tokens.
+The dataset contains 3.6k stories from 112 authors, with an average story length of 1500 tokens.
+Reddit forms the largest subset of our dataset with 3.2k stories given the larger availability of stories on the platform and the ease of data selection using their API  [^14] https://www.reddit.com/r/reddit/comments/145bram/addressing_the_community_about_changes_to_our_api/.
 Detailed statistics are provided in Table 6.
@@
 Themes:
-Figure 4 shows the plots depicting the percentage-normalized frequency of the top 10 themes for each source.
-To derive these themes, we prompt GPT-4o with each writing prompt to generate up to five one-word themes.
-We then aggregate the themes across all writing prompts and compute the normalized frequency by dividing the count of each theme by the total number of themes for that source.
-Overall, we observe that Reddit (Figure 4(a)), AO3 (Figure 4(b)), and Storium (Figure 4(e)) contain a higher prevalence of “conflict” themes, whereas N.Magazine (Figure 4(c)) and New Yorker (Figure 4(d)) focus more on family, loss, and identity.
-This distinction reflects the nature of story-writing settings across these sources: amateur stories tend to emphasize direct conflict resolution in their prompts, while expert-written stories explore deeper aspects of human relationships.
-Among Reddit, AO3, and Storium, Reddit exhibits a flatter distribution with greater thematic diversity, spanning identity, power, survival, humor, adventure, magic, and the supernatural.
-In contrast, AO3 has a sharper distribution, dominated by recurring themes of relationships, friendships, love, family, and secrets, reflecting a more domestic story-writing setting.
-Storium also shows a sharper distribution but favors themes such as space, danger, leadership, and war, highlighting a preference for war-centric narratives in collaborative storytelling.
+Figure 3: Story-writing themes for the Reddit source covered in Mythos.
+The pie chart Figure 3 shows the distribution of topic keywords (story-writing themes) obtained from the Reddit source of Mythos using BERTopic Grootendorst (2022) on the writing prompts.
+We see that our dataset covers a wide range of themes including Magic, Detectives, and Aliens exploring various genres of story-writing.
 Prompt for Writing Prompt Generation
@@
 Figure 4: Prompt for generating writing prompts for the stories in our dataset.
-Table 4: Dataset statistics, including the total number of stories, profiling (Profile) and generation (Gen) set distribution, number of authors, average stories per author, and token length in words.
 Dataset Total
@@
 Token Length
+Reddit 3262 2255 1007 60 54 1210
 AO3 239 159 80 17 14 1220
-Reddit 170 113 57 12 14 1210
 Storium 111 71 40 15 7 1120
-N.Magazine 32 18 14 10 3 1745
-NewYorker 38 22 16 10 4 1950
-Overall 590 383 207 64 8 1517
-(a) Reddit
-(b) AO3
-(c) N.Magazine
-(d) New Yorker
-(e) Storium
-Figure 4: Source-wise Percentage Normalized Frequency of the top 10 themes.
+N.Mag 32 18 14 10 3 1745
+N.York 38 22 16 10 4 1950
+Overall 3682 2525 1157 112 16 1517
+Table 6: Dataset statistics, including the total number of stories, profiling (Profile) and generation (Gen) set distribution, number of authors, average stories per author, and token length in words.
 Appendix B Author Writing Sheet
@@
 return $A_{|P|}$
-Algorithm 1 Constructing the Author Writing Sheet from a profiling set of stories
+Algorithm 1 Constructing the Author Writing Sheet from the profiling set
 Average Author Prompt for AO3 Generation
@@
 For profiling set sizes where $|P|>30$, generating the Author Writing Sheet is significantly more cost-effective than the Author Writing Summary.
-In this section, we analyze the cost of generating the Author Writing Sheet (see Section 3) and compare it with the cost of generating the Author Writing Summary (see Writing Summary) as a function of the profiling set size ($|P|$) using GPT-4o.
+In this section, we analyze the cost of generating the Author Writing Sheet (see Section 3) and compare it with the cost of generating the Author Writing Summary (see Summary) as a function of the profiling set size ($|P|$) using GPT-4o.
 We calculate the average cost per sample (a writing prompt and its corresponding story) as 0.06\$ for the Author Writing Sheet and 0.02\$ for the Author Writing Summary considering 2.50\$ per 1M tokens for input, and 10.00\$ per 1M tokens for output [^16] https://openai.com/api/pricing/.
@@
 Figure 19: Prompt for generating Story Rules based on the Author Writing Sheet and Writing Prompt.
-Persona Description
-As a storyteller, you craft narratives that are both humorous and transformative, often centering around light-hearted approaches to conflict resolution.
-Your plots are ingeniously structured, whether through the comedic misadventures of a rat named Pip or the empowering journey of Margaret Rose in a fantastical court.
-You delight in creating layered narratives, such as those told through fictional news articles, where humor and misunderstanding play pivotal roles.
-Your creativity knows no bounds, as you anthropomorphize animals to add whimsy and employ meta-satire to critique both human and alien perspectives.
-You blend historical and fantastical elements seamlessly, setting tales in places like Windsor Castle with characters that include both real-world figures and mythical beings.
-Your imaginative approach often intertwines religious and financial themes, crafting narratives that are as thought-provoking as they are entertaining.
-Character and setting development are your forte, with characters coming to life through their humorous dialogues and interactions.
-You explore their growth and empowerment, crafting relatable figures who navigate their internal struggles and aspirations.
-Your settings are vivid and engaging, providing a rich backdrop for the characters’ journeys.
-Your language use is playful and imaginative, enhancing the whimsical tone of your stories.
-You wield humor and irony with skill, using them to convey themes of misunderstanding and cultural critique.
-Whether through whimsical phrases or a blend of formal and playful language, your narratives are infused with a tone that is both magical and sophisticated, inviting readers into a world where the absurd and the profound coexist harmoniously.
-Table 6: Persona description for a Reddit author, structured into four paragraphs, each corresponding to a distinct narrative category: Plot, Creativity, Development (Character and Setting), and Language Use.
 Writing Prompt
@@
 Table 8: Personalized Story Rules obtained using the Author Writing Sheet for a Writing Prompt for a Reddit author.
+Persona Description
+As a storyteller, you craft narratives that are both humorous and transformative, often centering around light-hearted approaches to conflict resolution.
+Your plots are ingeniously structured, whether through the comedic misadventures of a rat named Pip or the empowering journey of Margaret Rose in a fantastical court.
+You delight in creating layered narratives, such as those told through fictional news articles, where humor and misunderstanding play pivotal roles.
+Your creativity knows no bounds, as you anthropomorphize animals to add whimsy and employ meta-satire to critique both human and alien perspectives.
+You blend historical and fantastical elements seamlessly, setting tales in places like Windsor Castle with characters that include both real-world figures and mythical beings.
+Your imaginative approach often intertwines religious and financial themes, crafting narratives that are as thought-provoking as they are entertaining.
+Character and setting development are your forte, with characters coming to life through their humorous dialogues and interactions.
+You explore their growth and empowerment, crafting relatable figures who navigate their internal struggles and aspirations.
+Your settings are vivid and engaging, providing a rich backdrop for the characters’ journeys.
+Your language use is playful and imaginative, enhancing the whimsical tone of your stories.
+You wield humor and irony with skill, using them to convey themes of misunderstanding and cultural critique.
+Whether through whimsical phrases or a blend of formal and playful language, your narratives are infused with a tone that is both magical and sophisticated, inviting readers into a world where the absurd and the profound coexist harmoniously.
+Table 9: Persona description for a Reddit author, structured into four paragraphs, each corresponding to a distinct narrative category: Plot, Creativity, Development (Character and Setting), and Language Use.
 C.1 Prompts
@@
 Figure 16 shows the prompt for generating story rules for a new writing prompt in the generation set using the story rules of the profiling set as few-shot demonstrations.
-Writing Sheet and Summary:
+Sheet and Summ:
 Figure 17 shows the prompt for generating the Author Writing Summary using all the stories in the author history.
@@
 Appendix D Experiments
-D.1 GPT4o-as-a-judge evaluation
+D.1 LLM-as-a-judge evaluation
 Prompt for evaluating Faithfulness to Writing History
@@
 Figure 20: Prompt for evaluating story alignment with an author’s writing style using the Author Writing Sheet.
-Prompt for Evaluating AI-Generated Stories Against Human-Written References
+Prompt for Evaluating Similarity to Author Story
 System Prompt
@@
 Figure 21: Prompt for evaluating AI-generated stories against a Human-Written reference based on a specified storytelling aspect.
-Figure 20 shows the prompt for GPT4o-as-a-judge evaluation for Faithfulness to Writing History using the Author Writing Summary as the reference.
-Figure 21 shows the for GPT4o-as-a-judge evaluation for Similarity to Author Story.
-We prompt GPT4o separately with each narrative category and its description to obtain the winner among the pair of stories (Saha et al., 2024).
+Figure 20 shows the prompt for LLM-as-a-judge evaluation for Faithfulness to Writing History using the Author Writing Summary as the reference.
+Figure 21 shows the for LLM-as-a-judge evaluation for Similarity to Author Story.
+We prompt OpenAI o4-mini separately with each narrative category and its description to obtain the winner among the pair of stories (Saha et al., 2024).
 D.2 Traditional Metrics
-We complement GPT-4o-as-a-judge evaluations with traditional metrics (Ivey et al., 2024; Huot et al., 2024) across three criteria: (1) lexical overlap, measured using BLEU and ROUGE-L (Xie et al., 2023); (2) story diversity, assessed via homogenization scores (Padmakumar and He, 2023); and (3) stylistic similarity, evaluated separately for the author history and the ground-truth author story using LUAR (Rivera-Soto et al., 2021).
+We complement LLM-as-a-judge evaluations with traditional metrics (Ivey et al., 2024; Huot et al., 2024) across three criteria: (1) lexical overlap, measured using BLEU and ROUGE-L (Xie et al., 2023); (2) story diversity, assessed via homogenization scores (Padmakumar and He, 2023); and (3) stylistic similarity, evaluated separately for the author history and the ground-truth author story using LUAR (Rivera-Soto et al., 2021).
 Specifically, we compute two stylistic similarity scores: one measuring the cosine similarity between the LUAR embeddings of the generated story and the mean embeddings of the stories in the author’s profiling set (author history), and another comparing the generated story to the ground-truth author story.
@@
 E.1 Faithfulness to Writing History
-Claims from the Author Writing Summary for Language Use
-1. A conversational and informal tone is employed, using humor and sarcasm.
-2. Dialogue reflects characters’ distinct voices and contributes to the comedic tone.
-3. Vivid and sometimes crude language adds to the immersive nature of the stories.
+E.1.1 Sheet and Summ achieve best faithfulness to writing history:
 Writing Prompt
-"At the Library at the End of the Universe, where dawn and dusk are the only times that exist, the Librarians are tasked with venturing into Bookworlds to fix stories that have gone astray—but what happens when the Chronicle itself begins to rewrite *their* story?"
-Oracle Story
-In the grand cosmic tapestry, where all threads converge into the singular, glittering endpoint known as the Library at the End of the Universe, there lies a place so vast, so staggeringly infinite, that it defies all comprehension.
-You just won’t believe how vastly, hugely, mind-bogglingly big it was.
-Imagine a labyrinth of endless corridors, each lined with shelves that stretch beyond the horizon, filled with books of every conceivable and inconceivable nature.
-In this whimsical realm, the concept of time is a mere suggestion, a faint whisper of dawn and dusk that play tag across the vaulted ceilings, casting elongated shadows that danced like giddy specters…….With books in hand and hearts alight with curiosity, they stepped forward into the unknown, ready to write the next chapter in the endless, wondrous saga of the Library at the End of the Universe.
-Writing Summary Story
-In the Library at the End of the Universe, ….
-But today, chaos decided to go on a caffeine-fueled rampage.
-The dawn had decided to join forces with dusk, creating a bizarre twilight that left the Librarians squinting like moles in a disco…Amongst the squad of book-botherers, there was Reginald, a Librarian so obsessed with order that he arranged his socks by the Dewey Decimal System.
-…Gertie grinned, her parrot squawking yet another Shakespearean soliloquy.
-Äye, who knew we’d end up in a romance novel, eh?
-But hey, at least we’re not in a tragedy!
-Average Author Story
-In the Library at the End of the Universe, where dawn and dusk were the only times that existed, a peculiar light filtered through the stained glass windows, casting kaleidoscopic shadows across endless shelves of books…..Livia, a seasoned Librarian with silver streaks in her auburn hair, stood before the Chronicle, her heart pounding with an uneasy rhythm….
-"This isn’t right," she murmured, her voice swallowed by the cavernous room.
-"The Chronicle should not change on its own." Beside her, Eamon, a younger Librarian with keen eyes and an inquisitive mind, frowned.
-"Perhaps it’s a reflection of a story gone wrong?
-One of the Bookworlds may be bleeding into our reality."
-Table 8: Example illustrating a case where the Oracle story deviates from the author’s documented story-writing characteristics in the Language Use category of the Author Writing Summary for a Storium author.
-Highlighted text in the Average Author Story shows partial alignment with the Author Writing Summary, while highlighted text in the Writing Summary story demonstrates stronger alignment.
-In contrast, the Oracle story lacks alignment, leading to the Average Author Story being preferred over the Oracle story and the Writing Summary story being preferred over the Average Author story.
-This explains the lower win-rate of the Oracle method compared to the Writing Summary method.
-Table 8 shows an example where the Oracle story deviates from the author’s documented story-writing characteristics as described in the Author Writing Summary for a Storium author.
-The author’s profiling set consists of three stories that consistently employ informal and colloquial language using humor and sarcasm, often conveyed through dialogue.
-However, for the new writing prompt in the generation set, the author’s style shifts to a more serious and mystic tone, diverging from their previously established preferences.
-In contrast, the Average Author method partially aligns with the Author Writing Summary by incorporating dialogue, while the Writing Summary method exhibits stronger alignment by more accurately capturing the author’s use of informal language and comedic tone.
-The Oracle story, however, fails to align with the author’s established style.
-This phenomenon results in the Average Author story being preferred over Oracle, and the Writing Summary story being preferred over the Average Author story, explaining the lower win-rate of the Oracle method compared to the Writing Summary method in GPT-4o-as-a-judge evaluation for Faithfulness to Author History.
-E.2 Effect of Persona on Similarity to Author Story
+"Welcome to Hell. To leave, all you must do is finish seven impossible tasks decided by your seven greatest sins. Four people have made it out so far. Will you be the fifth?"
+Author Writing Summary (Language Use)
+The author frequently uses dialogue to reveal character traits and dynamics.
+The author often employs vivid imagery to create an immersive atmosphere.
+The author often uses rhetorical questions to convey a character’s internal conflict or uncertainty.
+Average Author
+The air was thick with an acrid, sulfurous scent as Lucy stood at the threshold of Hell, a place she’d never believed in until now.
+A figure emerged from the shadows, tall and imposing, yet with a surprisingly warm smile.
+"Welcome to Hell," he said, his voice smooth and oddly comforting.
+"To leave, all you must do is finish seven impossible tasks decided by your seven greatest sins.
+Four people have made it out so far.
+Will you be the fifth?"
+Sheet
+I awoke on a bed of ash and cinders, the air around me thick with the scent of sulfur and regret. A laugh, echoing and absurd, filled the air—a demon guide, amused by my struggle, appeared at my side.
+"All this for a toy soldier?" it chortled, eyes gleaming with mischief. "You must learn to let go, or you’ll never find the exit."
+As I navigated the shifting corridors, I pondered the truth in its words. Each step forward was a step away from the past, from the child clutching stolen dreams.
+Delta
+Flames licked the edges of a crimson sky, and the ground beneath my feet felt alive, pulsing with a malevolent energy.
+I stood on the precipice of an endless abyss, the cacophony of wailing souls echoing in my ears.
+Emerging from the maze, I found myself in a grand hall, gilded and opulent, dripping with wealth beyond imagination. Coins rained from the ceiling, and jewels sparkled like fallen stars.
+Greed beckoned me to take, to hoard, to drown in the riches that lay at my feet. But I knew the price of succumbing; this treasure would bury me alive.
+Table 10: Example where Sheet outperforms Delta for Faithfulness to Author History.
+Sheet reflects the Author Writing Summary through vivid imagery, character-driven dialogue, and rhetorical questions, whereas Delta emphasizes setting and symbolism but lacks interactive dialogue and internal reflection.
+Table 10 shows an example for a Reddit author where the Sheet method outperforms Delta in terms of Faithfulness to Author History for the Language Use category.
+The Author Writing Summary highlights stylistic traits such as vivid imagery, dialogue, and rhetorical questions.
+The Sheet captures these through immersive narration and reflective exchanges with a demon guide.
+In contrast, Delta focuses on abstract setting descriptions and symbolic temptation but omits interactive dialogue and rhetorical structure, making it less faithful to the author’s characteristic style.
+E.1.2 Persona descriptions benefit amateur sources more than expert sources for Summ
+Author Writing Summary for Language Use
+The author employs conversational and accessible language to engage the reader.
+The author uses humor and wit in dialogue to add depth to characters and situations.
+The author frequently uses rhetorical questions to engage the reader and provoke thought.
+The author uses repetition for emphasis and to reinforce key themes or ideas.
+Persona Description for Language Use
+Your language use is conversational and accessible, engaging readers with humor and wit.
+You employ vivid imagery to create immersive settings and use rhetorical questions to provoke thought.
+Repetition reinforces key themes, while a mix of short and long sentences controls pacing and builds tension.
+Writing Prompt
+The reason humanity has never experienced any world-ending events is that future humans keep travelling back in time to stop them.
+You know this because there’s one of them in front of you now, telling you that you must absolutely not go into work today.
+Summ
+Evelyn’s mornings were a symphony of simplicity… "Whoa, what year is it?" he blurted out, eyes wide with what could only be described as temporal jet lag… "Listen, Evelyn, you absolutely cannot go to work today." Evelyn blinked, her skepticism as thick as the cream in her coffee.
+"And why is that?
+Did the universe finally decide to give me a spontaneous day off?"… Evelyn took a slow sip of her coffee, considering her options.
+Call the police?
+No, too much paperwork.
+Wake up from this bizarre dream?
+Preferable, but unlikely.
+Summ-nP
+The day started like any other.
+The alarm buzzed obnoxiously at 6:30 AM, shaking me from a dreamless sleep.
+I fumbled to silence it, already dreading the dreary commute to the office… The broadcast droned on about the usual—traffic jams, a potential storm, and the mayor’s latest scandal… until a peculiar sound interrupted my routine: a soft hum, followed by a voice behind me.
+"Good morning, citizen of the past!"
+Table 11: Example showing the role of persona descriptions in enhancing Language Use for a Reddit author for Faithfulness to Author History.
+The Summ method captures rhetorical engagement, and repetition more effectively than its counterpart without persona descriptions - Summ-nP.
+Author Writing Summary for Language Use
+Vivid and sensory language creates immersive atmospheres. Dialogue effectively reveals character emotions and tensions. The narrative voice carries a reflective and introspective tone.
+Persona Description for Language Use
+Your language use is vivid and sensory, painting immersive atmospheres that draw readers into your world.
+Through dialogue, you reveal the emotions and tensions that simmer beneath the surface, adding authenticity and depth to your characters.
+Your narrative voice carries a reflective and introspective tone, inviting readers to engage with the deeper themes of your stories and reflect on their own experiences.
+Writing Prompt
+After moving into their dream apartment, Romi witnesses a neighbor’s tragic death, but as the days pass, she begins to suspect the act was somehow meant to shatter her own carefully constructed life.
+Average Author
+The apartment was a sanctuary of light.
+Sunlight spilled through the floor-to-ceiling windows, pooling onto the wide-plank oak floors and casting soft, dappled shadows across the walls.
+Romi had waited years for this—an apartment with high ceilings, a balcony overlooking the park, and an airy, open layout that made her feel as though she were floating above the city.
+Summ
+The apartment on the top floor had a view that stretched far beyond the horizon.
+Every evening, when the sun dipped below the skyline, Romi would stand by the window, tracing the fading line of light with her eyes, feeling the world shrink to just her and this slice of heaven.
+The move had been a leap, a daring escape from the mundane into the realm of dreams where everything seemed possible.
+Summ-nP
+The apartment was everything Romi had ever imagined for herself… perched high above the city with its sprawling windows that captured the sunrise like a secret whispered just for her.
+The floors were a deep mahogany that gleamed under the soft light… The creaking of the floors became a language of its own, a series of codes and messages that she was determined to decipher.
+Table 12: Example showing that persona descriptions do not always help improve Faithfulness for expert sources like New Yorker.
+The Summ-nP captures vivid sensory details more effectively than the Summ when compared against the Average Story.
+Table 11 shows an example for a Reddit author where the Summ method performs better that its counterpart without persona descriptions (Summ-nP) in terms of Faithfulness to Author History, evaluated using claims from the Author Writing Summary for the Language Use category.
+Specifically, we see that the Summ method includes rhetorical questions and repetition for emphasis that the Summ-nP method misses that make the Summ method better aligned to the Author Writing Summary.
+Table 12 shows an example for a New Yorker author where the Summ method performs worse than its counterpart without persona descriptions (Summ-nP) in terms of Faithfulness to Author History, evaluated using claims from the Author Writing Summary for the Language Use category.
+Both the Average Author and Summ-nP better capture vivid sensory details compared to Summ, which instead emphasizes reflective and introspective tone, likely because the persona description places greater focus on this aspect.
+E.2 Similarity to Author Story
+E.2.1 Sheet outperforms all other methods
+Writing Prompt
+The world’s greatest hero has turned traitor and joined the forces of evil. Now, everyone wants to know why.
+Author Story
+"You were the chosen one!" the paladin screamed, pointing an accusatory finger at the prior hero.
+"You were to save us from evil, not join them!" The prior hero paused for a moment before nodding.
+"They had cookies."
+The paladin was dumbfounded. "You sold away everything you were, sold out the mortal race…for *confections?*"
+It is not that I am fed confections; it is that I make them, with a monster who treats me as their child, and I to her, as my mother—something I’ve never had.
+It is not that I am promised riches; it is that I earn them, and am encouraged to grow by a leader who *cares* for their subjects.
+Average Author
+The city held its breath. The world’s former protector, now veiled in darkness, stood at the gates of ruin. "Why?" the crowd asked. But no answer came—only silence, heavy with betrayal.
+His eyes, once symbols of hope, now flickered with uncertainty. A fallen hero. A world left wondering.
+Sheet
+In the shadow of my former self, I find clarity. "Not truth, perhaps," I admit, "but a different perspective. Perhaps the true villain is the one who blindly follows orders."
+Our conversation is interrupted by a comedic spectacle—a minion, oblivious to the tension, tries to carry a stack of forms taller than himself.
+He stumbles, papers flying everywhere, and I can’t help but laugh.
+The minion returns, this time with a clipboard. "New recruit orientation," he says sheepishly. "Welcome packet and all."
+I take the clipboard, glancing over the mundane details. "Orientation, really? I thought I’d at least get a dramatic initiation ceremony."
+Summary
+In the heart of the once-vibrant city of Elysium, now shrouded in shadows, the air buzzed with whispers of betrayal.
+Streets that once echoed with laughter were silent, as newspapers screamed of the hero’s fall from grace.
+"Will you stand with me, Alaric?
+Not as an ally of light or shadow, but as a seeker of truth?" He hesitated.
+Yet in the end, the promise of redemption proved irresistible.
+"I will," he declared.
+"For the sake of the city.
+For the sake of what we once stood for."
+Table 13: Example where Sheet outperforms both Average Author and Summary for Similarity to Author Story in the Creativity category.
+Sheet captures the Author Story’s blend of humor and imaginative reinterpretation of the prompt, while Summary follows a more conventional hero-fall narrative and lacks the light-hearted tone.
+Table 13 shows an example from the Reddit source where the Sheet method outperforms both the Average Author and Summary in terms of Similarity to the Author Story for the Creativity category.
+The Author Story takes a humorous and unconventional approach to the prompt, blending light-hearted banter and emotional subtext.
+The Sheet captures this tone through witty dialogue and imaginative reinterpretation (e.g., minion mishaps and onboarding rituals), while the Summary adopts a more serious, dramatic tone typical of heroic betrayal arcs.
+The Average Author response remains generic and somber, lacking the distinct creative flair that characterizes the Author Story.
+This example illustrates the Sheet method’s advantage in preserving the author-specific stylistic choices (unconventional re-interpretation of the prompt) beyond surface-level lexical similarity.
+E.2.2 Sheet still underperforms upper-bound Oracle method
+Writing Prompt
+You have been in a long, lonely coma.
+One day, Life and Death visit your consciousness, and get into a heated debate about what should be done with you.
+Indifferent to your fate, you like hearing them argue like an old married couple, and decide to try and set them up with each other.
+Author Story
+I’ve been drifting in this vast, empty void for so long… Akin to a leaf drifting away and about in a still pond, time doesn’t really exist here, and the silence is so deep that it almost feels like it has weight… I’m trapped in some sort of elevator that not only goes up and down, but also left, right, and diagonally… And hey, if it works out, beyond recovering from my coma, I might just score a very special invite to the universe’s most unlikely wedding.
+Average Author
+In the quiet echo of my mind, where only the soft hum of forgotten dreams lingered… It was like waking up, but not quite; it was more like tuning into a radio station I hadn’t realized existed… It was like listening to an old married couple who’d had the same argument for centuries… I imagined them, Life with her radiant glow, hands on her hips, eyes sparkling with the passion of a thousand dawns, and Death, calm and composed, a gentle shadow with a hint of a smile playing on his lips.
+Sheet
+Consciousness is a peculiar thing.
+It floats somewhere between the tangible and the illusory, like a whisper that hangs in the air, just out of reach.
+In the coma world, where I have resided for what feels like an eternity, time has no meaning… "Wouldn’t it be easier if you both just worked together?" I muse, my thoughts a ripple in the ether.
+Life laughs, a sound like tinkling bells.
+"Oh, wouldn’t that be something?
+But we are not meant to intertwine.
+We serve different purposes."… And in this place of dreams and whispers, I choose to live, fully and freely, in whatever form that may take.
+Oracle
+Floating in the vast, ethereal void of my own consciousness was akin to being a leaf drifting in a still pond—untouched by time, unanchored by space, and entirely alone.
+"Have you ever considered that maybe all this arguing is just… romantic tension?"
+Life blinked, momentarily at a loss for words. Death’s lips curled into a faint smile, his eyes twinkling with amusement. "Romantic tension?"
+Table 14: Example for a Reddit author where Sheet underperforms Oracle in capturing the Author Story’s lively and light-hearted tone and dialogue and instead leans toward introspection and philosophical reflection.
+Table 14 shows an example for a Reddit author where the Sheet method loses to the Average Author, while the Oracle method achieves the best performance for the Language Use category.
+The Sheet’s primary limitation is its lack of lively, humorous dialogue exchanges present in both the Author Story and the Oracle.
+Instead, it leans toward introspection and philosophical reflection.
+In contrast, the Average Author includes light-hearted descriptions that bring it closer to the Author Story, though it still falls short of the Oracle, which captures the tone and dialogue more effectively.
+E.2.3 Effect of Persona on Similarity to Author Story
 Writing Prompt
@@
 Leave now, and I will spare your lives…."
-Writing Sheet
+Sheet
 The castle itself seemed to rise in defiance, its ancient stones humming with a silent strength.
@@
 It was as if the castle was alive, an ally in his solitary defense.
-Writing Sheet nP
+Sheet nP
 The first attack came at dawn.
@@
 Table 15: Example showing the role of persona descriptions in determining similarity to the Author Story for a Reddit author.
-The Writing Sheet creatively blends fantasy and military strategy, making it more similar to the Author Story than the Average Author, while the Writing Sheet nP focuses primarily on direct conflict resolution.
+The Sheet creatively blends fantasy and military strategy, making it more similar to the Author Story than the Average Author, while the Sheet nP focuses primarily on direct conflict resolution.
 Table 15 highlights the importance of persona descriptions in achieving similarity to the Author Story for a Reddit author.
-The Author Story creatively blends military strategy and fantasy, exemplified by phrases like “the weapons and armors of the enemy shifted by his will.” Similarly, the Writing Sheet integrates both elements, as seen in “It was as if the castle was alive, an ally in his solitary defense.” In contrast, the Average Author and Writing Sheet nP primarily focus on directly resolving the writing prompt through descriptions of battle.
-The Average Author incorporates minor fantasy elements, such as “amplified by some unseen force,” whereas the Writing Sheet nP adheres strictly to conventional military tactics.
-As a result, the Writing Sheet is preferred over the Average Author, which in turn is preferred over the Writing Sheet nP.
-This outcome demonstrates that following persona descriptions enhances alignment with the author’s style, as the Writing Sheet successfully captures their characteristic genre-blending approach of including elements of fantasy in the story.
-The broader thematic variety of Reddit, i.e., war-themed prompts, led to a more descriptive Author Writing Sheet capturing the author’s style of combining fantasy with the writing prompt, which helped construct a useful persona to enable better personalization using the Writing Sheet method.
-E.3 Category-Wise Results for Similarity to Author Story
+The Author Story creatively blends military strategy and fantasy, exemplified by phrases like “the weapons and armors of the enemy shifted by his will.” Similarly, the Sheet integrates both elements, as seen in “It was as if the castle was alive, an ally in his solitary defense.” In contrast, the Average Author and Sheet nP primarily focus on directly resolving the writing prompt through descriptions of battle.
+The Average Author incorporates minor fantasy elements, such as “amplified by some unseen force,” whereas the Sheet nP adheres strictly to conventional military tactics.
+As a result, the Sheet is preferred over the Average Author, which in turn is preferred over the Sheet nP.
+This outcome demonstrates that following persona descriptions enhances alignment with the author’s style, as the Sheet successfully captures their characteristic genre-blending approach of including elements of fantasy in the story.
+The broader thematic variety of Reddit, i.e., war-themed prompts, led to a more descriptive Author Writing Sheet capturing the author’s style of combining fantasy with the writing prompt, which helped construct a useful persona to enable better personalization using the Sheet method.
+E.2.4 Category-Wise Results for Similarity to Author Story
 (a) AO3
 (b) Storium
-(c) N.Magazine
-(d) New Yorker
-Figure 22: Win-rate proportion of our personalization methods against Average Author for AO3, Storium, N.Magazine, and New Yorker across four narrative categories for similarity to author story.
-D stands for Delta, S for Writing Summary, and WS for Writing Sheet.
-Table 10: Win-rate percentages across narrative categories for similarity to the author story, averaged across sources.
-Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
-Category Delta Writing Sheet Writing Summary
-Plot 24-37 32-50 34-45
-Creativity 43-24 52-32 53-24
-Development 32-39 37-39 40-43
-Language Use 45-31 47-37 54-32
-Table 10 shows that Delta, Writing Summary, and Writing Sheet outperform Average Author in Creativity (23% higher) and Language Use (15% higher) (Huot et al., 2024), but perform worse in Plot and slightly worse in Development (Tian et al., 2024; Xu et al., 2024).
-Creativity and Language Use are less dependent on the writing prompt, making them easier to transfer across prompts, whereas Plot and Development are more closely tied to specific prompts, making generalization difficult, especially without thematic overlap.
-To better understand narrative category-wise personalization performance, Figure 22 shows the win-rates for sources - AO3, Storium, N.Magazine, and New Yorker similar to Figure 2 in the main paper.
-Overall, across all sources, we see that our proposed methods of Writing Sheet and Summary perform better than Delta across all four narrative categories.
-We see that Writing Summary generally performs better than Writing Sheet for sources like AO3 indicating that for the fanfiction domain explicitly contrasting the author-written story with the Average Story does not help much as the authors demonstrate limited stylistic variations among each other.
-In general, we observe that the win-rates of our personalization methods are higher than the Average Author for narrative categories of Creativity and Language Use as compared to Plot and Development confirming with prior results likely because Creativity and Language Use are less tied to the writing prompt than Plot and Development.
-E.4 Traditional Metrics
-Table 11: Comparison of methods across Lexical similarity, Diversity, and Style metrics.
+(c) N.Mag
+(d) N.York
+Figure 22: Win-rates proportions across narrative categories for Similarity to Author Story for AO3, Storium, N.Mag, and N.York across four narrative categories for similarity to author story.O: Oracle, D: Delta, Sh: Sheet, Su: Summ.
+Figure 22 shows win-rate proportions across narrative categories for Similarity to Author Story on AO3, Storium, N.Mag, and N.York.
+Similar to Reddit, Creativity and Language Use show higher personalization performance than Plot and Development, because the latter are more dependent on the specific writing prompt.
+E.3 Traditional Metrics
 Category Lexical Overlap Diversity (Homogenization) Style
 Method BLEU ROUGE-L ROUGE-L BERTScore Author History Author Story Length
-Ground Truth 0.122 0.698 0.8882 1516
-Oracle 0.0306 0.1577 0.1436 0.7132 0.7662 0.7712 1300
-Average Author 0.0123 0.1455 0.1412 0.7146 0.7378 0.7289 1383
-RAG 0.0118 0.1471 0.1432 0.7176 0.7393 0.7333 1326
-Delta 0.0115 0.146 0.147 0.7182 0.7462 0.7388 1298
-Writing Sheet 0.0101 0.1465 0.1452 0.7182 0.7735 0.7545 1344
-Writing Summary 0.0109 0.1463 0.1474 0.7202 0.7589 0.7418 1353
-Writing Sheet nP 0.0116 0.1475 0.1442 0.7168 0.7741 0.7563 1331
-Writing Summary nP 0.0112 0.1461 0.1456 0.7176 0.768 0.7465 1313
+Ground Truth 0.1222 0.7015 0.7402 1074
+Oracle 0.0311 0.1577 0.1427 0.7167 0.6385 0.7707 930
+Average Author 0.0122 0.1448 0.1415 0.7190 0.6149 0.7306 985
+RAG 0.0117 0.1463 0.1428 0.7212 0.6161 0.7359 947
+Delta 0.0114 0.1450 0.1462 0.7217 0.6218 0.7391 929
+Sheet 0.0102 0.1454 0.1440 0.7215 0.6446 0.7532 959
+Sheet-nP 0.0114 0.1464 0.1433 0.7208 0.6451 0.7544 950
+Summ 0.0108 0.1451 0.1460 0.7237 0.6324 0.7432 965
+Summ-nP 0.0111 0.1450 0.1445 0.7222 0.6400 0.7459 939
+Table 16: Comparison of methods across Lexical similarity, Diversity, and Style metrics for GPT-4o generated stories.
 Table 16 reports results using traditional metrics (Section D.2).
 Lexical Overlap and Diversity metrics yield similar values, as all methods use the same generation model, leading to overlapping lexical distributions and limiting these metrics’ ability to capture nuanced stylistic differences (Zheng et al., 2023; Xie et al., 2023).
-However, Writing Sheet nP consistently outperforms other methods, particularly in Style metrics for both similarity to Author History and Author Story, as measured by LUAR (Rivera-Soto et al., 2021).
-This improvement likely results from the Writing Sheet explicitly summarizing an author’s stylistic deviations from an Average Author, enhancing personalization.
+However, Sheet-nP consistently outperforms other methods, particularly in Style metrics for both similarity to Author History and Author Story, as measured by LUAR (Rivera-Soto et al., 2021).
+This improvement likely results from the Sheet explicitly summarizing an author’s stylistic deviations from an Average Author, enhancing personalization.
 Homogenization scores are slightly worse than the Average Author, likely explained by the increased similarity among the generated stories for the same author.
 Appendix F LLama Results
-Table 12: Percentage win-rate for Faithfulness to Author History for LLama 3.1 8B.
-Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
-Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
-AO3 34-59 30-59 54-40 51-34 45-46 51-40 50-36
-Reddit 39-40 39-54 46-44 54-30 56-35 47-46 54-40
-Storium 22-65 32-55 45-42 40-40 45-45 42-45 38-42
-N.Magazine 14-64 21-71 29-43 57-43 57-29 29-64 21-57
-New Yorker 12-80 20-73 40-60 53-33 53-27 40-47 47-40
-Overall 23-62 28-62 43-46 51-36 51-36 42-48 42-43
-Table 13: Percentage win-rate for Faithfulness to Writing History for LLama 3.1 70B.
-Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder
-Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
-AO3 41-48 44-42 35-52 41-36 57-29 50-36 64-26
-Reddit 42-46 28-65 40-44 53-37 53-35 53-37 67-21
-Storium 30-57 52-32 38-50 35-40 40-38 25-60 48-35
-N.Magazine 50-29 29-64 43-57 50-36 50-50 64-14 64-29
-New Yorker 53-47 33-67 27-60 67-20 40-53 73-13 60-27
-Overall 43-45 37-54 37-53 49-34 48-41 53-32 61-28
-Table 14: Percentage win-rate for Similarity to Author Story for LLama 3.1 8B. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
-Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
-AO3 66-26 31-51 52-34 48-38 38-50 52-34 51-32
-Reddit 75-23 32-51 51-37 42-49 37-40 39-49 46-44
-Storium 42-50 25-60 30-57 18-68 30-52 20-57 28-60
-N.Magazine 43-50 36-43 43-57 21-50 36-64 43-57 36-43
-New Yorker 27-60 07-80 13-60 20-67 13-67 0-93 20-73
-Overall 51-42 26-57 38-49 30-54 31-55 31-58 36-50
-Table 15: Percentage win-rate for Similarity to Author Story for LLama 3.1 70B. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
-Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
-AO3 62-24 35-50 35-55 36-46 44-40 46-39 39-40
-Reddit 74-21 37-49 40-42 47-39 49-37 42-44 44-33
-Storium 50-42 30-57 32-52 18-57 38-52 28-62 42-40
-N.Magazine 79-7 43-21 36-50 29-43 36-50 21-43 50-29
-New Yorker 60-27 20-60 27-60 27-67 20-80 20-60 27-67
-Overall 65-24 33-47 34-52 31-50 37-52 31-50 40-42
 Below, we discuss results using Llama 3.1 8B and Llama 3.1 70B (Dubey et al., 2024) as story generation models, conditioned on GPT-4o-generated story rules and persona descriptions.
 Each method is evaluated by GPT-4o against the Average Author story generated by the respective Llama model, not GPT-4o.
+F.1 Faithfulness to Author History
+Method Reddit AO3 Storium N.Mag N.York All
+RAG 39 30 32 21 20 28
+Delta 46 54 45 29 40 43
+Sheet 54 51 40 57 53 51
+Sheet-nP 47 51 42 29 40 42
+Summ 56 45 45 57 53 51
+Summ-nP 54 50 38 21 47 42
+Table 17: Win-rates (%) for Faithfulness to Writing History vs. Average Author baseline for LLaMA 3.1 8B evaluated by GPT-4o.
+Method Reddit AO3 Storium N.Mag N.York All
+RAG 28 44 52 29 33 37
+Delta 40 35 38 43 27 37
+Sheet 53 41 35 50 67 49
+Sheet-nP 53 50 25 64 73 53
+Summ 53 57 40 50 40 48
+Summ-nP 67 64 48 64 60 61
+Table 18: Win-rates (%) for Faithfulness to Writing History vs. Average Author baseline for LLaMA 3.1 70B evaluated by GPT-4o.
 Larger models improve instruction-following and personalization:
-Tables 12 and 13 show Faithfulness to Writing History for Llama 3.1 8B and 70B, while Tables 14 and 15 report similarity to the author story.
+Tables 17 and 18 show Faithfulness to Writing History for Llama 3.1 8B and 70B, while Tables 19 and 20 report similarity to the author story.
 Across both criteria, Oracle achieves higher win-rates with Llama 3.1 70B, indicating that larger models enhance instruction-following and personalization (Chung et al., 2024).
 Cross-Model Persona Descriptions Aid Smaller Models but Not Larger Ones for Faithfulness to Writing History:
-For Llama 3.1 8B (Table 12), Writing Sheet and Writing Summary achieve the highest scores (9% higher than nP), demonstrating the benefit of persona descriptions.
-However, for Llama 3.1 70B (Table 13), Writing Sheet nP and Writing Summary nP outperform their persona-based counterparts by approximately 9%, suggesting that GPT-4o-generated personas provide no benefit to other models with stronger instruction-following capabilities (McKenzie et al., 2023).
+For Llama 3.1 8B (Table 17), Sheet and Summ achieve the highest scores (9% higher than nP), demonstrating the benefit of persona descriptions.
+However, for Llama 3.1 70B (Table 18), Sheet nP and Summ nP outperform their persona-based counterparts by approximately 9%, suggesting that GPT-4o-generated personas provide no benefit to other models with stronger instruction-following capabilities (McKenzie et al., 2023).
 This is likely because persona descriptions of an author’s writing style are model-dependent and do not transfer well to other models with comparable performance (Shashidhar et al., 2024).
+F.2 Similarity to Author Story
+Method Reddit AO3 Storium N.Mag N.York All
+Oracle 75 66 42 43 27 51
+RAG 32 31 25 36 7 26
+Delta 51 52 30 43 13 38
+Sheet 42 48 18 21 20 30
+Sheet-nP 39 52 20 43 0 31
+Summ 37 38 30 36 13 31
+Summ-nP 46 51 28 36 20 36
+Table 19: Win-rates (%) for Similarity to Author Story vs. Average Author baseline for LLaMA 3.1 8B evaluated by GPT-4o.
+Method Reddit AO3 Storium N.Mag N.York All
+Oracle 74 62 50 79 60 65
+RAG 37 35 30 43 20 33
+Delta 40 35 32 36 27 34
+Sheet 47 36 18 29 27 31
+Sheet-nP 42 46 28 21 20 31
+Summ 49 44 38 36 20 37
+Summ-nP 44 39 42 50 27 40
+Table 20: Win-rates (%) for Similarity to Author Story vs. Average Author baseline for LLaMA 3.1 70B evaluated by GPT-4o.
+Method Sheet Sheet-ft Summ Summ-ft
+Reddit (%) 42 62 37 60
+Table 21: Win-rates (%) for Similarity to Author Story with and without fine-tuning for LLaMA 3.1 8B evaluated by GPT-4o for the Reddit subset of our dataset.
+Fine-tuned (ft) models show significant improvements over their counterparts with just prompting.
 Limited Cross-Model Generalization of the Author Writing Sheet Compared to the Author Writing Summary for Similarity to Author Story:
-For both Llama 3.1 8B (Table 14) and Llama 3.1 70B (Table 15), Writing Summary outperforms Writing Sheet by 5% and 9%, respectively (Shashidhar et al., 2024).
+For both Llama 3.1 8B (Table 19) and Llama 3.1 70B (Table 20), Summ outperforms Sheet by 5% and 9%, respectively (Shashidhar et al., 2024).
 This outcome likely stems from the construction process: the Author Writing Sheet contrasts an LLM-generated Average Story with the author-written story, making it more model-dependent.
 In contrast, the Author Writing Summary does not rely on the LLM-generated Average Author story, allowing for greater generalization across models.
+Fine-tuning on the Author Writing Sheet/ Summary boosts performance for LLaMA 3.1 8B:
+We experiment with fine-tuning the LLaMA 3.1 8B model on the Author Writing Sheet and Summary for the Reddit subset by prepending the Author Writing Sheet/ Summary to the writing prompt during story generation, in addition to prompting alone (see Table 21).
+We compare the fine-tuned method against an equivalent Average Author baseline, fine-tuned on the writing prompt alone without the Author Writing Sheet/ Summary.
+Our method with Sheet achieves a 62% win-rate against the baseline, compared to 42% (‘Sheet’ row of Table 19) for the prompting-only approach, yielding a 20% improvement.
+Similarly, for Summ, we achieve a 60% win-rate against the baseline, compared to 37% (‘Summ’ row of Table 19) for the prompting-only approach, yielding a 23% improvement.
+These results highlight the potential of Author Writing Sheet/ Summary generated by a different model (GPT-4o) to effectively guide another model (LLaMA 3.1 8B) when fine-tuned rather than just prompted.
 Appendix G Human Evaluation
@@
 The annotators provided consent by signing a Google form that outlined the instructions of the task including the purpose of the research study.
-Each annotator evaluated 25 author stories and 75 story pairs, corresponding to the three personalization methods: Delta, Writing Sheet, and Writing Summary.
+Each annotator evaluated 25 author stories and 75 story pairs, corresponding to the three personalization methods: Delta, Sheet, and Summ.
 Collectively, our annotators evaluated 45 unique author stories (135 story pairs), of which 15 author stories were annotated by all three annotators to assess inter-annotator agreement.
@@
 The sky was painted in hues of orange and pink, a beautiful contrast to the sorrow in my heart.
-I sat there, watching the sun dip below the horizon, and for the first time, I felt truly mortal.
-I realized that even heroes are not immune to the pain of loss, the fragility of life…….And as the night embraced the world, I knew that Ethan’s spirit would live on, a beacon of hope and courage in the heart of the world’s greatest hero.
-Writing Sheet
+Sheet
 The Hero’s Journey: A Tale of Capes and Cereal In the heart of the mythical city of Avalon Heights, where ancient heroes once strode and the skies shimmered with the glow of enchanted skyscrapers, the world’s greatest hero faced their most daunting foe: the frailty of life itself.
@@
 Max, with dreams and hopes as boundless as the sky, taught her that heroism wasn’t defined by powers but by the courage to embrace life’s uncertainties.
-Table 17: Qualitative Example where Writing Sheet method is preferred over Average Author by all three annotators for similarity to the ground-truth author story for a Reddit author.
-Writing Prompt
-"Late at night, the King of Rohan and his queen sneak into the pantry for a forbidden snack, only to find that stealing cured venison is far more complicated—and far more amusing—than they anticipated."
-Ground-truth Author Story
-If anyone—a kitchen boy, a maid, or, Eru forbid, even the housekeeper Wynhild—had wandered near the pantry in the quiet hour just before dawn, they would have heard a muffled exchange from the depths of the storeroom: “Try shifting this bowl here…” “Easy, you’re on me.” “Oh, does that bother you, my lady?” “Get off me; I can’t pull it out!……Lothíriel watched her royal husband with a half-incredulous, half-amused smile, as he held the large salami with the pride of a merchant at the autumn fair……To Lothíriel, he looked like a great golden-haired squirrel with the morsel stuffed in one cheek.
-How he could be both a fierce warrior and a wise sovereign as well, was a mystery she had yet to solve.
-His voice came out muffled as he swallowed.
-“You’re right, but if it ever gets out, I’ll say the sweet Queen of the Mark was struck by an irresistible craving for spiced meat in the dead of night—and who would deny anything to the future mother of the next royal heir?” He crossed his arms, looking quite pleased with his idea, but Lothíriel’s eyes widened at the bold lie, and she covered her face with her hand, foreseeing trouble ahead.
-Average Author
-Éomer, King of Rohan, tiptoed through the dimly lit corridors of Meduseld, holding his breath as he reached the door to the pantry.
-He glanced back at Lothíriel, his queen, who followed close behind, stifling a giggle.
-The flickering torchlight cast playful shadows on the stone walls, and Éomer couldn’t help but feel a rush of exhilaration.
-“We shouldn’t be doing this,” Lothíriel whispered, though her eyes sparkled with mischief.
-“Ah, but where’s the fun in being King and Queen if we can’t indulge in a little late-night adventure?…..
-“You’ve always had a knack for adventure.
-It’s one of the reasons I fell in love with you.” Lothíriel’s expression softened, and she leaned in to kiss him, the taste of venison still on her lips.
-“And I love you for indulging my whims, even when they lead us to pilfering from our own pantry.”…..
-Éomer took Lothíriel’s hand, and they slipped out of the pantry, leaving behind only the echo of their laughter and the memory of a stolen moment in the heart of Rohan.
-“Next time,” Éomer whispered as they crept back to their chambers, “we’ll try the cheese.”
-Writing Sheet
-The moon cast a silvery glow over Meduseld, its light filtering through the high windows to illuminate the grand halls of the Golden Hall.
-Éomer, the formidable King of Rohan, found himself tiptoeing like a child on a dare, his hand clasped firmly in Lothíriel’s, whose laughter was a soft melody in the silence…..
-“Are you sure this is wise, my love?” “Wise?
-No.
-Fun?
-Absolutely,” she replied, her laughter a soft ripple in the silence.
-Éomer shook his head, a reluctant smile tugging at his lips as he watched her nimble fingers work to untie the savory prize.
-He was a warrior, a king, yet here he was, caught in the delightful absurdity of this secretive adventure, his heart light with the thrill of it all.
-“Besides,” Lothíriel continued, handing him a piece of the cured venison with a flourish, “what’s the worst that could happen?
-A tale of the gallant King brought low by his own appetite?” Éomer took the offered morsel, his grin widening as he tasted the rich, smoky flavor.
-“Perhaps the bards will sing of our escapades instead of my battles,” he mused, the thought surprisingly pleasing.
-…..As they slipped back into the corridors, leaving the pantry and its shadows behind, Éomer realized that it was not the venison he would remember, but the adventure, the laughter, and the love that had made the night truly unforgettable.
-Table 18: Qualitative Example where the Average Author method is preferred over Writing Sheet by all three annotators for similarity to the ground-truth author story for an AO3 author.
-Writing Prompt
-"In a desert town under the watchful eyes of royal guards and oppressive banners, you and your team have five minutes to locate a mysterious contact in a pink house before chaos erupts—failure is not an option."
-Ground-truth Author Story
-# Chapter 1: Danger Close *La Makan, 12:54 standard time* Wind whistled through the close streets of *La Makan*, stirring grains of sand with all the effect of a man attempting to blow salt through eternity.
-It whipped around the corners of the alleys, rustling banners hanging from balconies and glassless windows.
-It looked for something, a thing it would never find…..
-Either way wouldn’t end well for them.
-Our contact gave us an area, but not an address.
-We have, he checked his watch, five minutes until the distraction starts.
-We need to find them and get them out.
-Problem is, we don’t know what they look like, only that their house is pink.
-He glanced over his shoulder at the multicolored buildings.
-Any one of the scoured pink buildings could’ve been their one.
-“Vic, you knew Xitor the best of any of us.
-I’m trusting you to get us there.
-Everyone else, let’s keep our hats on, kay?"
-Average Author
-The sun hung low in the sky, casting long shadows across the dusty streets of the desert town of Vespera.
-Time was of the essence; you had five minutes to locate your mysterious contact before chaos erupted.
-Failure was not an option.
-“Stay sharp,” you whispered to your team, your voice barely audible over the din of the market.
-"But time is short.
-We must act quickly." The weight of the moment settled over you like a heavy mantle.
-The fate of Vespera rested on the success of this meeting.
-As the figure began to speak, outlining the plan that would either save or doom the town, you couldn’t shake the feeling that this was only the beginning.
-The real challenge was yet to come.
-Writing Sheet
-The sun beat down mercilessly, casting a harsh glare on the town’s whitewashed walls.
-The narrow alleys were alive with whispers and hurried footsteps, each person a part of the unfolding drama.
-My team and I moved through the bustling streets, the oppressive presence of royal guards a constant reminder of the stakes.
-We had five minutes—no more, no less—to locate the contact in the elusive pink house before chaos erupted.
-The memory of a previous mission flickered in my mind: the sun setting behind the dunes, the sound of distant drums, and the rush of adrenaline as we narrowly escaped a similar fate.
-That mission had gone wrong, leaving scars that still haunted my dreams.
-Failure was not an option today… Outside, the tension remained, a silent storm waiting to break.
-But for now, we had succeeded.
-The town’s frenetic energy continued, unaware of our brief triumph.
-The banners flapped, the guards watched, and the countdown to chaos continued, but we had found our mark.
-The mission was still on course, and the desert town held its breath, waiting for the next move.
-Table 19: Qualitative Example where all three annotators select ‘Tie’ for similarity to the author story of a Storium author.
+Table 23: Qualitative example where the Sheet method is preferred over Average Author by all annotators for a Reddit author.
+Sheet captures the author’s creative use of humor, satire, and unconventional programming jargon, which the Average Author misses.
 Writing Prompt
@@
 She continues to read to her mother, to sit by her side, and to hold her hand through the long nights.
-One crisp morning, as the first frost dusts the garden in a delicate shimmer, Marjorie slips away.
-Clara is by her side, her hand still entwined with her mother’s, and she feels a profound stillness settle over her…..
-Writing Sheet
+Sheet
 Every morning, Nurse Clara began her rounds at St.
@@
 With each cycle of her rounds, Clara realized that the light had always been there, waiting to be uncovered, and that she had the power to help others see it too.
-Table 20: Qualitative Example where all annotators select different winners for similarity to the author story of a New Yorker author.
+Table 24: Qualitative Example where annotators select different winners for a New Yorker author.
+The Author Story uses metaphor and subtext to reframe hospice as both a physical and emotional space.
+While Sheet captures these elements through cyclic structure and introspective tone, the subjective nature of subtext leads annotators to different interpretations of similarity.
+Writing Prompt
+"Late at night, the King of Rohan and his queen sneak into the pantry for a forbidden snack, only to find that stealing cured venison is far more complicated—and far more amusing—than they anticipated."
+Ground-truth Author Story
+If anyone—a kitchen boy, a maid, or, Eru forbid, even the housekeeper Wynhild—had wandered near the pantry in the quiet hour just before dawn, they would have heard a muffled exchange from the depths of the storeroom: “Try shifting this bowl here…” “Easy, you’re on me.” “Oh, does that bother you, my lady?” “Get off me; I can’t pull it out!……Lothíriel watched her royal husband with a half-incredulous, half-amused smile, as he held the large salami with the pride of a merchant at the autumn fair……To Lothíriel, he looked like a great golden-haired squirrel with the morsel stuffed in one cheek.
+How he could be both a fierce warrior and a wise sovereign as well, was a mystery she had yet to solve.
+His voice came out muffled as he swallowed.
+“You’re right, but if it ever gets out, I’ll say the sweet Queen of the Mark was struck by an irresistible craving for spiced meat in the dead of night—and who would deny anything to the future mother of the next royal heir?” He crossed his arms, looking quite pleased with his idea, but Lothíriel’s eyes widened at the bold lie, and she covered her face with her hand, foreseeing trouble ahead.
+Average Author
+Éomer, King of Rohan, tiptoed through the dimly lit corridors of Meduseld, holding his breath as he reached the door to the pantry.
+He glanced back at Lothíriel, his queen, who followed close behind, stifling a giggle.
+The flickering torchlight cast playful shadows on the stone walls, and Éomer couldn’t help but feel a rush of exhilaration.
+“We shouldn’t be doing this,” Lothíriel whispered, though her eyes sparkled with mischief.
+“Ah, but where’s the fun in being King and Queen if we can’t indulge in a little late-night adventure?…..
+“You’ve always had a knack for adventure.
+It’s one of the reasons I fell in love with you.” Lothíriel’s expression softened, and she leaned in to kiss him, the taste of venison still on her lips.
+“And I love you for indulging my whims, even when they lead us to pilfering from our own pantry.”…..
+Éomer took Lothíriel’s hand, and they slipped out of the pantry, leaving behind only the echo of their laughter and the memory of a stolen moment in the heart of Rohan.
+“Next time,” Éomer whispered as they crept back to their chambers, “we’ll try the cheese.”
+Sheet
+The moon cast a silvery glow over Meduseld, its light filtering through the high windows to illuminate the grand halls of the Golden Hall.
+Éomer, the formidable King of Rohan, found himself tiptoeing like a child on a dare, his hand clasped firmly in Lothíriel’s, whose laughter was a soft melody in the silence…..
+“Are you sure this is wise, my love?” “Wise?
+No.
+Fun?
+Absolutely,” she replied, her laughter a soft ripple in the silence.
+Éomer shook his head, a reluctant smile tugging at his lips as he watched her nimble fingers work to untie the savory prize.
+He was a warrior, a king, yet here he was, caught in the delightful absurdity of this secretive adventure, his heart light with the thrill of it all.
+“Besides,” Lothíriel continued, handing him a piece of the cured venison with a flourish, “what’s the worst that could happen?
+A tale of the gallant King brought low by his own appetite?” Éomer took the offered morsel, his grin widening as he tasted the rich, smoky flavor.
+“Perhaps the bards will sing of our escapades instead of my battles,” he mused, the thought surprisingly pleasing.
+…..As they slipped back into the corridors, leaving the pantry and its shadows behind, Éomer realized that it was not the venison he would remember, but the adventure, the laughter, and the love that had made the night truly unforgettable.
+Table 25: Example where all annotators prefer the Average Author over Sheet for an AO3 author.
+The Average Author aligns better with the story’s playful tone and king–queen dynamic, aided by LLM familiarity with Lord of the Rings fanfiction tropes.
+Sheet adds internal conflict that doesn’t fit this prompt.
+Writing Prompt
+"In a desert town under the watchful eyes of royal guards and oppressive banners, you and your team have five minutes to locate a mysterious contact in a pink house before chaos erupts—failure is not an option."
+Ground-truth Author Story
+# Chapter 1: Danger Close *La Makan, 12:54 standard time* Wind whistled through the close streets of *La Makan*, stirring grains of sand with all the effect of a man attempting to blow salt through eternity.
+It whipped around the corners of the alleys, rustling banners hanging from balconies and glassless windows.
+It looked for something, a thing it would never find…..
+Either way wouldn’t end well for them.
+Our contact gave us an area, but not an address.
+We have, he checked his watch, five minutes until the distraction starts.
+We need to find them and get them out.
+Problem is, we don’t know what they look like, only that their house is pink.
+He glanced over his shoulder at the multicolored buildings.
+Any one of the scoured pink buildings could’ve been their one.
+“Vic, you knew Xitor the best of any of us.
+I’m trusting you to get us there.
+Everyone else, let’s keep our hats on, kay?"
+Average Author
+The sun hung low in the sky, casting long shadows across the dusty streets of the desert town of Vespera.
+Time was of the essence; you had five minutes to locate your mysterious contact before chaos erupted.
+Failure was not an option.
+“Stay sharp,” you whispered to your team, your voice barely audible over the din of the market.
+"But time is short.
+We must act quickly." The weight of the moment settled over you like a heavy mantle.
+The fate of Vespera rested on the success of this meeting.
+As the figure began to speak, outlining the plan that would either save or doom the town, you couldn’t shake the feeling that this was only the beginning.
+The real challenge was yet to come.
+Sheet
+The sun beat down mercilessly, casting a harsh glare on the town’s whitewashed walls.
+The narrow alleys were alive with whispers and hurried footsteps, each person a part of the unfolding drama.
+My team and I moved through the bustling streets, the oppressive presence of royal guards a constant reminder of the stakes.
+We had five minutes—no more, no less—to locate the contact in the elusive pink house before chaos erupted.
+The memory of a previous mission flickered in my mind: the sun setting behind the dunes, the sound of distant drums, and the rush of adrenaline as we narrowly escaped a similar fate.
+That mission had gone wrong, leaving scars that still haunted my dreams.
+Failure was not an option today… Outside, the tension remained, a silent storm waiting to break.
+But for now, we had succeeded.
+The town’s frenetic energy continued, unaware of our brief triumph.
+The banners flapped, the guards watched, and the countdown to chaos continued, but we had found our mark.
+The mission was still on course, and the desert town held its breath, waiting for the next move.
+Table 26: Qualitative Example where all annotators select ‘Tie’ for a Storium author.
+While the Author Story uses an open-ended narrative with unresolved character dynamics, both the Average Author and Sheet introduce resolution and tension-driven plots, missing the author’s intended tone.
 G.3 Qualitative Examples
-We present qualitative examples from our human annotation process, covering various cases: instances where all annotators prefer Writing Sheet, all prefer Average Author, or select ‘Tie.’ Additionally, we include an example where annotators disagree, choosing different winners, highlighting the challenging and subjective nature of the annotation task.
-All annotators prefer Writing Sheet:
-Table 17 shows an example where all three annotators prefer the Writing Sheet method over the Average Author for similarity to the author story of a Reddit author.
-The annotators highlight several key aspects that make the Writing Sheet story more aligned with the ground-truth author story: (1) The Writing Sheet method incorporates humor and satire, using phrases such as “a tale of capes and cereal,” “invented breakfast hurricane,” and “World’s Greatest Hero Faces Greatest Challenge Yet: Parenting,” which mirror the lighthearted and satirical tone of the author story.
-(2) Unlike the Average Author story, which concludes with the son’s death, the Writing Sheet story reinterprets the prompt, portraying Titania as an advocate for families with special children.
+We present qualitative examples from our human annotation process, covering various cases: instances where all annotators prefer Sheet, all prefer Average Author, or select ‘Tie.’ Additionally, we include an example where annotators disagree, choosing different winners, highlighting the challenging and subjective nature of the annotation task.
+All annotators prefer Sheet:
+Table 23 shows an example where all three annotators prefer the Sheet method over the Average Author for similarity to the author story of a Reddit author.
+The annotators highlight several key aspects that make the Sheet story more aligned with the ground-truth author story: (1) The Sheet method incorporates humor and satire, using phrases such as “a tale of capes and cereal,” “invented breakfast hurricane,” and “World’s Greatest Hero Faces Greatest Challenge Yet: Parenting,” which mirror the lighthearted and satirical tone of the author story.
+(2) Unlike the Average Author story, which concludes with the son’s death, the Sheet story reinterprets the prompt, portraying Titania as an advocate for families with special children.
 This aligns with the ground-truth author story’s approach, where the son engages with coding and the narrative avoids a sorrowful ending.
-(3) In terms of language style, the Writing Sheet story employs imaginative elements such as “The Hero’s Journey: A Tale of Capes and Cereal” and “The Avalon Gazette,” which parallel the ground-truth author story’s unconventional use of computer programming references, such as “Syntax Error: Unexpected Power.”
+(3) In terms of language style, the Sheet story employs imaginative elements such as “The Hero’s Journey: A Tale of Capes and Cereal” and “The Avalon Gazette,” which parallel the ground-truth author story’s unconventional use of computer programming references, such as “Syntax Error: Unexpected Power.”
 All annotators prefer Average Author:
-Table 18 shows an example where all three annotators prefer the Average Author method over the Writing Sheet for similarity to the author story of an AO3 author.
-The primary reason cited by annotators is that the Writing Sheet method introduces a strong internal conflict for the king, portraying him as guilty or hesitant about sneaking into the kitchen for a late-night snack.
+Table 25 shows an example where all three annotators prefer the Average Author method over the Sheet for similarity to the author story of an AO3 author.
+The primary reason cited by annotators is that the Sheet method introduces a strong internal conflict for the king, portraying him as guilty or hesitant about sneaking into the kitchen for a late-night snack.
 His character is depicted as struggling with the decision, given his stature as a formidable ruler expected to lead battles.
-Additionally, the Writing Sheet creates a sharp contrast between the king and queen, with the queen depicted as joyful and mischievous, encouraging the king to abandon his guilt.
+Additionally, the Sheet creates a sharp contrast between the king and queen, with the queen depicted as joyful and mischievous, encouraging the king to abandon his guilt.
 In contrast, both the ground-truth Author Story and the Average Author story emphasize the playful banter and affectionate dynamic between the king and queen, focusing on the shared experience of sneaking into the kitchen rather than introducing introspection or internal conflict.
-This shift in focus stems from the Writing Sheet being influenced by the author’s past writing history, where internal conflicts are a recurring theme.
-However, in this case, character development—being more writing-prompt-specific than broader categories like creativity or language use—did not transfer well.
-The author’s original story prioritizes external actions and lighthearted interactions rather than deep introspection, leading to a mismatch between the Writing Sheet’s personalized adaptation and the intended tone of the new prompt.
+This shift in focus stems from the Sheet being influenced by the author’s past writing history, where internal conflicts are a recurring theme.
+However, in this case, character development being more writing-prompt-specific than broader categories like creativity or language use did not transfer well.
+The author’s original story prioritizes external actions and lighthearted interactions rather than deep introspection, leading to a mismatch between the Sheet’s personalized adaptation and the intended tone of the new prompt.
 All Annotators Choose Tie:
 Table 26 shows an example where all three annotators select ‘Tie’ for similarity to the author story of a Storium author.
-The primary reason is that both the Average Author story and the Writing Sheet story deviate significantly from the ground-truth author story in terms of plot.
+The primary reason is that both the Average Author story and the Sheet story deviate significantly from the ground-truth author story in terms of plot.
 The ground-truth story adopts an open-ended approach to both character development and plot, where character relationships remain undefined, and the narrative focuses on anticipation rather than action.
 In contrast, the Average Author story follows a conventional narrative arc, featuring fast-paced dialogue and character interactions that build tension and lead to resolution.
-Similarly, the Writing Sheet story also resolves the narrative but differs from the Average Author method by emphasizing past missions and their consequences, aligning with the author’s tendency to reference character backstories.
+Similarly, the Sheet story also resolves the narrative but differs from the Average Author method by emphasizing past missions and their consequences, aligning with the author’s tendency to reference character backstories.
 However, both methods fail to capture the open-ended nature of the original story.
@@
 Not sad but joyous, a honeymoon.”
-The annotator who selected Writing Sheet as the winner states, ‘‘Although Average Author and the Author Story depict relationships between close family members, Writing Sheet is more similar to the Author Story in Plot, Creativity, and Development by exploring the meaning and emotional responses to the term hospice.’’ This annotator highlights the presence of subtext in the Writing Sheet story and its attempts to mimic the author’s cyclic style—“Each door held a story, a life in its final chapter… she would ask the same question…” and “Clara knew that work was not about erasing the pain or the grief, but about finding moments of grace within it…”.
+The annotator who selected Sheet as the winner states, ‘‘Although Average Author and the Author Story depict relationships between close family members, Sheet is more similar to the Author Story in Plot, Creativity, and Development by exploring the meaning and emotional responses to the term hospice.’’ This annotator highlights the presence of subtext in the Sheet story and its attempts to mimic the author’s cyclic style—“Each door held a story, a life in its final chapter… she would ask the same question…” and “Clara knew that work was not about erasing the pain or the grief, but about finding moments of grace within it…”.
 The annotator who selected Average Author as the winner states, ‘‘Average Author aligns more closely with the Author’s Story in terms of Plot and Language Use.
 This is because both stories center on someone taking care of a family member in their own home, and growing from a lack of peace into a sense of peace, which is expressed through the choice of language.
-Writing Sheet also indicates a development from lack of peace to acceptance, but the familial tie is what pushes Average Author over the top.’’
+Sheet also indicates a development from lack of peace to acceptance, but the familial tie is what pushes Average Author over the top.’’
 The annotator who selected ‘Tie’ states, ‘‘I think it will be hard for one of these to resemble the Author Story since it was so unique.

### v3

@@
 Stage 1 constructs an Author Writing Sheet with Claim ($C$) and Evidence ($E$) pairs capturing the author’s story-writing characteristics across narrative categories.
-It is derived from the author’s history of writing prompts ($wp$), author-written stories ($s_{a}$), and LLM-generated Average Stories ($s_{b}$) representing a typical author’s response to the same prompt.
+It is derived from the author’s history of writing prompts (wpwp), author-written stories ($s_{a}$), and LLM-generated Average Stories ($s_{b}$) representing a typical author’s response to the same prompt.
 Stage 2 uses the Author Writing Sheet to role-play the author, incorporating tailored story rules and a persona description for personalized generation.
 Large Language Models (LLMs) are increasingly used in interactive systems to support human users in co-writing tasks (Fang et al., 2024; Yuan et al., 2022; Chung et al., 2022).
-A key challenge in these settings to generate personalized outputs that align with individual user preferences (Yeh et al., 2024), which improves engagement, agency, and writing quality (Zhao et al., 2025).
-Personalization is also key in educational applications to deliver effective feedback to students (Anderson et al., 1985), and may support users facing writer’s block (Yuan et al., 2022) or help second language learners in learning a new language (Baffour et al., 2023; Su et al., 2023).
+A key challenge in these settings is to generate personalized outputs that align with individual user preferences (Yeh et al., 2024), which improves engagement, agency, and writing quality (Zhao et al., 2025).
+Personalization is also key in educational applications to deliver effective feedback to students (Anderson et al., 1985; Ashok Kumar and Lan, 2024), and may support users facing writer’s block (Yuan et al., 2022) or help second language learners in learning a new language (Baffour et al., 2023; Su et al., 2023).
 Despite these potentials, personalization in story writing has received limited attention.
-Story-writing personalization faces several challenges, including a lack of existing resources for evaluating personalization, the inherent subjectivity in user preferences, and methods to evaluate the extent of personalization.
-To address these gaps, we construct Mythos, a dataset of 3.6k human-written stories ($\approx$1500 tokens each) from 112 authors (16 stories each) across five distinct writing settings: Reddit, AO3, Storium, Narrative Magazine, and New Yorker.
+Story-writing personalization faces several challenges, including a lack of existing resources for evaluating personalization and the inherent subjectivity in user preferences.
+To address these gaps, we construct Mythos, a dataset of 3.6k human-written stories ($\approx$1500 tokens each) from 112 authors (16 stories each [^1] The reported average (16) is the mean of source-level stories-per-author counts (Table 8), while the global average across all authors is 32 because Reddit has more stories per author than New Yorker and Narrative Magazine.) across five distinct writing settings: Reddit, AO3, Storium, Narrative Magazine, and New Yorker.
 We split each author’s stories chronologically into two sets: an earlier Profiling set used to infer their writing characteristics, and a later Generation set used to evaluate how well LLMs can emulate their distinctive writing behavior (Salemi et al., 2024).
@@
 Details on our dataset can be found in Appendix A.
+Stories per Author Across Sources:
+As summarized in Table 8, Reddit authors contribute substantially more stories on average (54 per author), while expert sources such as N.Mag and N.York have only 3–4 stories per author due to limited public availability, reflecting natural differences in data accessibility between amateur and expert sources.
 Dataset Name Size Sources Prompt Length Story Length Author IDs
@@
 This yields an intermediate Author Writing Sheet $A_{t}^{\prime}$ (Shashidhar et al., 2024; Krishna et al., 2020).
-To construct the final Author Writing Sheet $A_{|P|}$, we iteratively merge each intermediate sheet $A_{t}^{\prime}$ with the aggregated sheet $A_{t-1}$, following a moving update approach (Chang et al., 2023).
+To construct the final Author Writing Sheet $A_{|P|}$, we iteratively merge each intermediate sheet $A_{t}^{\prime}$ with the aggregated sheet $A_{t-1}$, following a moving update approach (Chang et al., 2024).
 To reduce redundancy and manage output length, we prompt the LLM with a prompt (LLMcombine) to group equivalent Claims, select the best Evidence, and retain ungrouped Claims.
-We further limit each narrative category to 10 Claim-Evidence pairs prioritizing grouped Claims to fit within the LLM’s context window in subsequent iterations.
+We further limit each narrative category to 10 Claim-Evidence pairs prioritizing grouped Claims to fit within the LLM’s context window in subsequent iterations resulting in a total of 40 Claim-Evidence pairs per author.
 Each Evidence is tagged with a timestamp indicating its source story.
-Similar to Knowledge Tracing in education (Liu et al., 2022; Scarlatos and Lan, 2024), our method processes each story sequentially.
+Similar to Knowledge Tracing in education (Liu et al., 2022; Kumar et al., 2023; Scarlatos et al., 2025), our method processes each story sequentially.
 Doing so enables efficient updates to the Author Writing Sheet without having to re-process prior stories every time a new story is analyzed (Yeh et al., 2024).
@@
 This result is likely due to personas for expert authors failing to capture the nuanced and subtle characteristics of their writing styles, which are reflected in their Author Writing Summary.
-See Appendix E.1.2 and Tables 11, 12 for examples.
+See Appendix E.1.2 and Tables 13, 14 for examples.
 6.2 Similarity to Author Story
@@
 See Appendix E.2.4 and Figure 22 for plots from other data sources.
+Method k = 1 k = 3 k = 5
+Sheet 70 67 66
+RAG 38 43 36
+Summ 50 56 49
+Table 5: Effect of the number of few-shot demonstrations ($k$) for different methods, evaluated on 100 Reddit stories from 20 authors for Similarity to Author Story.
+The proposed Sheet method outperforms the best of other methods.
+Best results for each $k$ are bolded.
+Sheet performs best across different numbers of few-shot demonstrations:
+We analyze the impact of the number of few-shot demonstrations ($k$) on model performance using 100 Reddit stories from 20 authors for different methods.
+As shown in Table 5, the Sheet method achieves the highest win rate at $k=1$ (70%), with a slight decline as $k$ increases, consistent with observations by (Richardson et al., 2023).
+In contrast, both RAG and Summ peak at $k=3$ but drop at $k=5$, suggesting sensitivity to prompt length and context dilution.
+Even at their best settings, these methods lag behind Sheet by 27% and 14%, respectively, demonstrating the superiority of our proposed method.
+Method Plot Creat. Dev. Lang. Overall
+Sheet (full) 48 74 45 60 70
+no Plot 37 65 45 61 49
+no Creat. 46 63 51 52 53
+no Dev. 37 57 36 46 39
+no Lang. 38 61 45 46 44
+Table 6: Ablation study on the four narrative categories in the Author Writing Sheet, conducted on 100 stories from 20 Reddit authors for Similarity to Author Story.
+Each cell reports the win rate (%) of the corresponding method (row) for each narrative category (column).
+Creat., Dev., and Lang.
+denote Creativity, Development, and Language Use, respectively.
+Overall indicates the combined win rate across all categories.
+Method names such as “No Plot” indicate variants of the Sheet method with the corresponding category excluded.
+Best numbers for each category are bolded.
+All four narrative categories contribute to Sheet’s performance:
+To examine the contribution of each narrative category in the Sheet method for similarity to the author story, we conducted an ablation study on a subset of 100 stories from 20 Reddit authors.
+In this setting, we removed one category at a time from both the story rules and persona descriptions in our Sheet method while keeping the others fixed.
+As shown in Table 6, we see that, the full model that uses all four categories achieves the highest overall win rate of 70%, outperforming all ablated variants.
+We also see that, excluding any single category lowers performance on the corresponding narrative dimension, confirming that all four aspects, Plot, Creativity, Development, and Language Use jointly contribute to effective personalization.
+We further note that removing Development yields the largest drop (31% gap) compared to the full model, indicating that while Development is harder to personalize, it is also essential, as its exclusion results in the lowest overall win rate and highlights its importance to effective personalization.
 6.3 Traditional Metrics
@@
 For AO3, LLMs’ familiarity with common fanfiction tropes [^5] https://archiveofourown.org/admin_posts/25888 strengthens the Average Author baseline (see Table 27).
-For Storium, the small profile size and weak stylistic differentiation among authors beyond open-ended endings (Xu et al., 2024; Tian et al., 2024) limit the effectiveness of personalization methods (see Table 26).
+For Storium, the small profile size and weak stylistic differentiation among authors beyond open-ended endings (Xu et al., 2025; Tian et al., 2024) limit the effectiveness of personalization methods (see Table 28).
 These findings suggest that personalization is most effective when authors explore diverse topics or display distinct traits, such as complex narratives, creativity, or advanced language use.
@@
 Recent works have introduced benchmark datasets for personalizing LLM outputs in tasks like email and news writing, focusing on shorter outputs (e.g., 300 tokens for product reviews (Kumar et al., 2024), 850 for news writing (Shashidhar et al., 2024)).
-These methods infer user traits from history for task-specific personalization (Sun et al., 2024, 2025; Pal et al., 2024; Li et al., 2023; Salemi et al., 2025).
+These methods infer user traits from history for task-specific personalization (Sun et al., 2024, 2025; Pal et al., 2025; Li et al., 2023; Salemi et al., 2025).
 In contrast, we address the more subjective task of long-form story writing, with author stories averaging 1500 tokens.
@@
 To further mitigate unforeseen harms, the dataset will be released under an **Educational or Academic Research, Non-Commercial Use** license, following (Akoury et al., 2020).
+Acknowledgments
+The authors are partially supported by the NSF under grant IIS-2202506.
+We thank our Upwork annotators for their dedicated efforts in the annotation process.
+We are also grateful to Alexander Scarlatos and Jaewook Lee for their valuable discussions and feedback.
+Finally, we thank members of the UMass ML4Ed and UMass NLP labs for their constructive insights and support on this project.
 Appendix A Dataset
-In this section, we describe the details of the construction of our dataset which includes five distinct story-writing sources, Reddit [^8] https://www.reddit.com/r/WritingPrompts/, AO3999https://archiveofourown.org/, Storium [^10] https://storium.cs.umass.edu/, Narrative Magazine [^11] https://www.narrativemagazine.com/, and New Yorker [^12] https://www.newyorker.com/.
+In this section, we describe the details of the construction of our dataset which includes five distinct story-writing sources, Reddit [^9] https://www.reddit.com/r/WritingPrompts/, AO3101010https://archiveofourown.org/, Storium [^11] https://storium.cs.umass.edu/, Narrative Magazine [^12] https://www.narrativemagazine.com/, and New Yorker [^13] https://www.newyorker.com/.
 Selection Constraints:
@@
 Figure 4: Prompt for generating writing prompts for the stories in our dataset.
-Dataset Total
-Stories
-Profile Size Gen Size Total
-Authors
-Stories/
-Author
-Token Length
+Source Total Profile Gen. Authors St./Auth. Tokens
 Reddit 3262 2255 1007 60 54 1210
@@
 Overall 3682 2525 1157 112 16 1517
-Table 6: Dataset statistics, including the total number of stories, profiling (Profile) and generation (Gen) set distribution, number of authors, average stories per author, and token length in words.
+Table 8: Dataset statistics for the five sources in Mythos, reporting total stories (Total), profiling split (Profile), generation split (Gen.), number of authors (Authors), average stories per author (St./Auth.), and average story length in tokens (Tokens).
 Appendix B Author Writing Sheet
@@
 These standards provide a structured framework for evaluating narrative elements such as plot, creativity, development, and language use within the Author Writing Sheet, ensuring a systematic and interpretable representation of an author’s unique storytelling style.
-1 Input: profiling set $P=\{(wp_{t},s_{at})\mid t=1,\ldots,|P|\}$,
+1Input: profiling set $P=\{(wp_{t},s_{at})\mid t=1,\ldots,|P|\}$,
 where $wp_{t}$ is the writing prompt and $s_{at}$ is the author’s story at time-step $t$.
@@
  // average story
-[…]
+6 $s_{bt}\leftarrow\text{LLM\textsubscript{avg}}(wp_{t})$
  // Intermediate Author Writing Sheet
-[…]
+7 $A_{t}^{\prime}\leftarrow\text{LLM\textsubscript{sheet}}(wp_{t},s_{bt},s_{at})$
  // Author Writing Sheet ($A_{t}$)
-[…]
+8 $A_{t}\leftarrow\text{LLM\textsubscript{combine}}(A_{t}^{\prime},A_{t-1})$
  // Steps within combination:
@@
 We calculate the average cost per sample (a writing prompt and its corresponding story) as 0.06\$ for the Author Writing Sheet and 0.02\$ for the Author Writing Summary considering 2.50\$ per 1M tokens for input, and 10.00\$ per 1M tokens for output [^17] https://openai.com/api/pricing/.
-In an interactive writing setting, where an author adds a new submission at each time step, the total cost of generating the Author Writing Sheet is 0.06\$ * $|P|$, while the total cost of generating the Author Writing Summary is 0.0015$|P|^{2}$ + 0.019$|P|$.
-The cost of generating the Author Writing Summary scales quadratically with $|P|$, as the entire author history must be reprocessed at each step, leading to cumulative costs as the profiling set grows.
-In contrast, the cost of generating the Author Writing Sheet grows linearly, as each new submission incurs a fixed cost of 0.06\$ by running the for loop in Algorithm 1 once.
+[…]
 Figure 12 shows the cost analysis of generating the Author Writing Sheet and Summary using GPT-4o as a function of the profiling set size $|P|$ in an interactive writing setting.
 Up to $|P|=30$, the cost of generating both remains approximately the same at 1.16\$.
-However, as $|P|$ increases, the cost of generating the Author Writing Summary grows at a rate of 0.025$|P|$ + 0.30 compared to the Author Writing Sheet which continues to scale linearly.
+However, as $|P|$ increases, the cost of generating the Author Writing Summary grows at a rate of 0.025|P||P| + 0.30 compared to the Author Writing Sheet which continues to scale linearly.
 Specifically, for $|P|=100$, the cost of generating the Author Writing Summary is 17\$, compared to 6\$ for the Author Writing Sheet.
@@
 Larger models improve instruction-following and personalization:
-Tables 17 and 18 show Faithfulness to Writing History for Llama 3.1 8B and 70B, while Tables 19 and 20 report similarity to the author story.
+Tables 19 and 20 show Faithfulness to Writing History for Llama 3.1 8B and 70B, while Tables 21 and 22 report similarity to the author story.
 Across both criteria, Oracle achieves higher win-rates with Llama 3.1 70B, indicating that larger models enhance instruction-following and personalization (Chung et al., 2024).